A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7550–7599

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.01

Title word cross-reference

1 [387, 565, 288]. 2 [264, 426]. 3 [411, 348, 310, 295, 522, 426, 311, 454, 307, 427]. 3d + t [341]. 4 [303]. 5 [373]. st [231]. _{1} [112]. _{2} [112]. _{□} [470]. M [221]. h [26]. k [551, 228]. L(2, 1) [29]. L_{2} [366]. M, N [223]. n [120].

-Adhesive [221, 223]. -Center [387]. -cubics [526]. -Cycle [120]. -D [307]. -Gons [373]. -Highway [387]. -Holes [373]. -Hop [565]. -Median [288]. -Placement [551]. -Quasi [26]. -Span [29].

10th [637]. 11th [639, 623, 618]. 12th [626, 627, 628]. 13th [620]. 14th [638]. 15th [634, 624, 617]. 16th [614].

2.0 [237]. 21st [238, 622, 622]. 22nd [612, 613].

3 [256]. 38th [611]. 3D [406, 415, 463, 593, 407, 434]. 3G [501].

4 [518]. 4D [300].

60th [629]. 6th [621, 610].

7th [622].

Aberrations [344]. above [34]. Abstract [23]. Abstraction [213]. Abstractions [535]. AC [581]. AC-CS [581]. Academic [130]. Accelerator [50]. Access [261, 255, 262]. Account [15]. Accuracy [425]. Accurate [93]. Achieving [123]. Acquainted [182]. Acquisition [519, 528]. across [506, 364, 317, 507]. Action [173, 174, 430, 318, 356, 49, 358]. Activation [132]. Active [335, 286]. Activities [320, 170]. Activity [357, 464, 543]. Actuality [242]. Actually [240]. Ad [258]. Ad-hoc [258]. Adaboost [598]. Adaptable [84]. Adaptation [149, 499]. Adaptive [245, 523, 250, 170, 62, 420]. Adaptivity [63]. Adhesive [221, 223]. Adjusted [525]. Ado [493]. Advanced [270]. Advances [269, 618]. Advertisement [398]. Aerial [519]. Affect [443]. Affective [359, 177]. after [304]. against [227]. Age [595]. Age-Related [595]. Agent [394, 559, 396, 401, 395, 630, 630]. Agent-Based [396, 401, 395]. Aggregation [491]. Aging [442, 443]. Aid [478]. Aided [422]. AIMSA [617]. Airports [104]. Alcalá [629]. Algebra [192, 191, 190, 207]. Algebraic [200, 620]. Algebras [204, 199]. Algorithm [511, 587, 105, 438, 287, 581, 339, 520, 579, 264, 263, 586, 552, 132, 288, 576, 65, 582, 589]. Algorithmic [590]. Algorithmically [518]. Algorithms [560, 33, 289]. Alignment [417, 370, 334, 369]. Alignment-Based [369]. Allocation [397, 101, 498]. Alloy [521]. Alternating [17]. Alternative [473]. Alzheimer [296]. AMASE [170]. Ambient [312]. American [120]. among [374, 67]. Amorphous [47]. Analog [47]. Analysis [54, 8, 82, 430, 500, 236, 625, 238, 296, 502, 540, 440, 58, 243, 394, 416, 637, 181, 453, 640, 136, 486, 589, 359, 362, 408, 444, 409, 94, 352, 178, 595, 499]. Analytical [413]. Analyzing [299, 136]. Anatomical [593]. Anatomy [301, 600]. Anchored [417]. Aneurysms [515, 598]. Angle [391]. Annotation [348, 156, 337]. Announcement [566, 554, 558, 559, 553, 567]. Anomaly [586, 588, 573]. Anonymization [116]. Anonymous [554]. Anthropometric [455]. Anticipatory [400]. Appeal [493]. Application [346, 278, 513, 173, 520, 214, 485, 603, 604, 302, 515, 356, 106, 103]. Applications [226, 503, 133, 617, 507]. Applied [165, 583]. Applying [99]. Approach [174, 151, 397, 577, 422, 323, 61, 96, 247, 454, 583, 584, 495, 588, 262, 468, 369, 69, 361]. Approaches [152]. Approximate [332]. Approximation [288]. Approximations [512, 212]. Apps [158]. Arbitrary [553]. Architectural [532]. Architecture [54, 42, 39, 477, 518, 503, 498]. Architectures [492, 505, 494]. Archiving [125]. Area [391, 26, 132].

Ariadne [8]. Arithmetic [523]. Arrangements [246]. Arrays [484]. Arrows [78]. Art [360]. Articles [136]. Articulated [409, 315]. Artificial [580, 36, 583, 586, 572, 612, 613, 65, 639, 617, 612, 613]. Aspects [481]. Aspectual [88]. Assessment [88, 185, 436]. Assignment [568]. Associate [53]. Association [364, 163, 486]. Associative [54, 581]. Assumptions [557, 471]. Asynchronous [556, 152]. Atlas [297, 593]. Atlas-Based [297]. Atlases [601]. Atoms [1]. Attachment [132]. Attacks [256]. Attention [180]. Attribute [320, 89]. Attribute-Based [320]. Attributed [284, 217]. Attributes [452, 448, 451]. Attribution [221]. Auction [397, 398]. August [639]. Australia [630]. Austria [634]. Authentication [507]. Authoring [250]. Autism [577]. Automata [8, 534, 226, 537, 6, 541, 4, 543, 2, 546]. Automated [192, 474, 598]. Automatic [594, 225, 517, 488, 144, 410, 185, 604, 596]. Automatically [119]. Automaton [535]. Autonomous [283]. AVC [518]. AVC/H.264 [518]. Aware [472, 567, 606]. Awareness [241]. Axioms [202]. Axis [510].

B [579]. B-Cell [579]. Back [13, 90, 107, 186, 308, 402, 548, 608]. Background [333]. Backstage [242]. Backward [513]. Bag [174]. Bag-of-Key-Poses [174]. Bait [580]. Ballast [93]. Banking [281]. Banks [521, 527]. Barrel [528]. Base [55]. Based [346, 83, 233, 336, 139, 141, 195, 48, 104, 174, 582, 165, 630, 87, 287, 261, 258, 490, 265, 297, 292, 540, 590, 111, 203, 517, 143, 394, 256, 255, 303, 565, 521, 227, 542, 419, 61, 350, 527, 317, 306, 209, 161, 115, 251, 586, 95, 71, 286, 117, 362, 396, 131, 572, 618, 155, 268, 491, 408, 401, 213, 545, 320, 293, 395, 228, 147, 398, 369, 573, 113, 576, 355, 358, 607, 103, 163, 598, 499]. Baseline [421]. Batch [514]. Bayesian [175, 453, 335, 40, 369]. Beginner [575]. Behavior [225, 482, 172, 619]. Behaviors [181, 183]. Behaviour [359, 184]. Benchmark [441]. Benchmarking [104, 354]. Benefit [479]. Bertinoro [635]. Best [368]. Beta [583]. Better [484]. between [218, 6, 282, 578]. Bias [321]. Bicolour [392]. Bifurcating [55]. Binary [41, 326]. Biometrics [455]. Biosignal [270]. Bipartite [196]. Birthday [629]. Bisections [34]. Black [79]. Blind [354, 344]. Blinking [527]. BLIP [569]. Bloom [569]. Bodies [310]. Body [179, 455, 178]. Boltzmann [68, 46, 72]. Bookmarking [151]. Bordeaux [610]. Borrowed [217]. Bottom [328, 468]. Bottom-Up [328, 468]. Boundary [604]. Bounded [26, 563, 544]. Bounding [450]. Bounds [539, 30, 34]. Box [79]. Brain [299, 305, 303, 302, 300]. Brazil [614]. Brazilian [614]. Breed [322]. Bremen [621]. Bridging [218]. Brief [566, 554, 558, 559, 553, 567]. Browsers [506]. browsing [506]. Bugs [79]. Bulgaria [617]. Burst [293]. Business [139, 485].

C [165, 114, 86]. C-FOAM [165]. Calculation [97]. Calibration [524]. Cambridge [620]. Camera [368, 175, 460, 354, 524, 435, 350, 313, 454]. Canada [638]. Canals [597]. Cancelling [312]. Capability [148]. Capacitated [98]. Capacities [101]. Capacity [563]. Capture [520].

Cardiac [297]. Cardiomyocyte [306]. Cargo [92]. CARS [287]. Carving [599]. Case [562, 105, 494]. Cases [28]. Cat [71]. Categorical [115]. Categories [202, 200, 193, 216]. Categorization [582, 352]. Catering [106]. Causal [67]. CCA [359]. Ceiling [175]. Cell [582, 579, 341, 589, 307]. Cells [392, 580]. Cellular [501]. Census [122]. Center [387]. Centers [75]. Centric [502, 429]. Century [238, 231, 622]. Cerebral [598]. Certifiable [193]. Certificate [255]. Certificate-Based [255]. Chain [341]. Chaining [42]. Chains [96]. Chair [616]. Challenging [419]. Changes [299, 301, 480, 595]. Channel [533]. Characterization [306, 420]. Characterized [21]. Characterizing [298]. Cheating [154]. Checking [8, 226, 539, 5, 10, 225, 496, 531, 567, 480]. China [615]. Choice [206]. Chords [526]. Chromatic [379]. Circuits [47]. CISIM [623]. Cities [254]. Claims [471]. Clarke [105]. Classes [21]. Classification [406, 509, 581, 325, 259, 323, 322, 356, 451]. Classroom [237, 154]. Claw [19]. Claw-Free [19]. CLES [245]. Clinical [596]. Clips [350]. Clique [19, 20, 23]. Clocks [541]. Clonal [587, 585]. Closed [17, 279]. Closed-form [279]. Closures [195]. Cloud [253, 490, 501, 502, 561, 161, 495, 497, 635, 498, 505, 493, 494]. Cloud-Based [161]. Clouds [311, 427]. CloudSim [502]. Cloudy [345]. Clustering [287, 75, 565, 277, 115, 286, 117, 576, 134]. Clustering-Based [115, 576]. Co [316, 506, 284, 398, 289]. Co-browsing [506]. Co-dependent [398]. Co-detection [316]. Co-evolution [284]. Co-regularized [289]. Code [479, 488, 114, 80]. Coding [533, 456, 523]. Cognitive [142]. Cohesive [284]. Collaborations [493]. Collaborative [406, 152]. Collaborator [260]. Collapsible [2]. Color [346, 514, 430]. Color-Depth [430]. Colored [382]. Coloring [190]. Combinatorial [375, 397, 183, 407]. Combined [66, 604, 405]. Combining [428, 318, 435]. Common [579]. Commonality [323]. Communication [52, 563, 396, 550]. Communication-Efficient [550]. Community [120, 236, 246]. Compact [388, 451, 326]. Comparability [123]. Comparative [75, 486, 573]. Comparing [112]. Comparison [297, 63]. Comparisons [59]. Competence [238]. Competitive [16]. Complementation [30]. Complete [437]. Completeness [211, 198, 199]. Completeness-Driven [211]. Complex [137, 121, 207, 480, 57]. Complex-Valued [57]. Complexity [30, 544, 302]. Components [266, 78, 443]. Composed [20]. Composing [170]. Composite [320]. Composition [490, 5, 491, 499]. Compositionality [224]. Compound [143]. Comprehension [479]. Compressed [70]. Compression [523]. Computation [569, 447]. Computational [110, 629, 69, 615, 629]. Computations [189]. Compute [51]. Computer [169, 626, 627, 628, 611, 620, 636, 623, 626, 627, 628, 631, 632, 633]. Computing [195, 630, 196, 501, 579, 179, 495, 635, 498]. Concept [221, 262]. Concepts [611, 102, 589, 206]. Concurrency [215]. Concurrent [228]. Conditional [364, 598, 499]. Conditions [214]. Conductance [48]. Conductance-Based [48]. Conference [636, 639, 623, 616, 621, 626, 627, 628, 634, 624, 615, 637, 620, 635, 618, 617, 622, 612, 613]. Confidence [539, 440, 437, 286]. Confidence-Based [286]. Confidential [125]. Confidentialising [126]. Configurations [313]. Confluence [220]. Conformance [496]. Conjunction [625]. Connecting [392]. Consensus [287]. Consequences [190]. Conservation [395]. Considering [401]. Consistency [146, 480]. Consistent [111, 300]. Constant [83]. Constrained [456, 356]. Constraint [146]. Constraints [532, 348, 438, 486, 271]. Constructing [16]. Construction [593, 74, 329, 103, 163]. Constructivism [157]. Consumer [435]. Container [95, 93, 100, 94]. Containers [97]. Content [164]. Contest [256]. Context [472, 479, 504]. Context-Aware [472]. Contexts [200, 217]. Contextual [181]. Contingency [128]. Continuous [535, 380, 370, 367]. Contraction [365]. Contracts [227]. Control [183, 95, 71, 544, 262, 257]. Controlled [481]. Controllers [144]. Controlling [261]. Convex [390, 379, 380, 327]. Convexifying [381]. ConvNets [64]. Coordinated [400]. Coping [471]. Coq [194]. Core [48, 537]. Corpus [162, 131]. Corpus-Based [131]. Correct [249]. Correction [344]. Correlation [42]. Correspondence [601, 337]. Correspondences [436, 414]. Corresponding [346]. Cortical [603, 298, 300, 596]. Counter [6, 4]. Counterexample [538]. Counterexample-Guided [538]. Counterpart [212]. Counting [465]. Coupled [210]. Courses [156]. Cover [20]. Coverage [227]. Covering [484]. Creating [159, 441]. CreaTools [515]. Crises [281]. Cross [478]. Cross-Language [478]. Crossing [391]. Crossover [149]. Crowd [462]. CS [581]. CSA [587]. CSA/IE [587]. CSCL [239]. CT [596]. CTA [112, 110]. cubics [526]. Cues [318, 435]. Cultural [137]. Culture [157]. Cumulative [526]. Curation [125]. Current [238, 218]. Curves [418]. Customer [134]. Customisable [505]. Cut [347]. Cycle [120, 33, 103]. Cycle-Based [103]. Cylinders [27].

D [426, 411, 348, 310, 295, 303, 522, 264, 426, 311, 454, 307, 427]. D/ [426]. Daily [430]. Damage [321]. Dashboard [251]. Data [299, 43, 430, 145, 516, 616, 121, 564, 625, 433, 259, 590, 440, 324, 520, 303, 176, 115, 589, 126, 274, 545, 276, 320, 283, 573, 113]. Data-Variable [545]. Database [500, 354, 269]. Databases [616, 329, 337, 127]. Dataset [321]. Datasets [121]. Day [345]. DCA [582]. Decentralized [138]. Decidability [220]. Deciding [194]. Decision [204, 102, 3]. Declarative [274]. Decomposition [463, 19, 267]. Deconvolution [354]. Decreasing [389]. Dedekind [202]. Dedicated [629]. Defining [475]. Definition [592, 75]. Deformable [417, 593, 449, 450, 410]. Deformation [599]. Deformations [600]. Deforming [312]. Degrading [558]. Degree [30, 31]. Delay [38]. Delays [534]. Demand [401]. Demons [298]. Demonstrated [183]. Demonstrations [631, 632, 633]. Dendritic [582, 589]. Denoising [346]. Dense [307, 337, 436]. Density [461]. Dependent [53, 398]. Deployment [502, 515]. Depth [363, 336, 430, 433, 324, 429, 432]. Depth-Limited [363].

Derivational [131]. Derivative [58]. Descent [58, 57]. Description [143]. Descriptive [281]. Descriptor [605, 327, 358]. Descriptors [428, 296, 326]. Design [530, 479, 245, 209, 248, 480]. Design-Space [530]. Designing [254, 237, 124]. Destination [462]. Detect [324, 271]. Detect-Track [271]. Detecting [461]. Detection [442, 447, 258, 481, 580, 130, 175, 450, 454, 464, 263, 586, 584, 588, 545, 457, 293, 434, 573, 355, 427, 316]. Detectors [465]. Determination [313]. Determining [29]. determinism [197]. Deterministic [514]. Develop [196, 515]. Developable [431]. Developed [102]. Developers [479]. Developing [182, 247]. Development [211, 181, 298, 478, 262, 470, 302, 167]. Developments [275]. Device [507]. Devices [506, 240]. Diagram [392]. Diagrams [204, 40]. Dictionary [323]. Differences [177]. Different [506]. Differential [272, 513, 583]. Differentially [569, 128]. Differentially-Private [569, 128]. Diffuse [126]. Diffusion [416, 293]. Digital [48, 261, 238, 250]. Digraphs [553]. Dimension [28]. Dimensional [511, 587, 517]. Dioid [203]. Directional [521]. Discerning [242]. Discovering [448, 131, 164]. Discovery [285, 624]. Discrete [363, 510, 341]. Discrimination [319]. Discriminative [422, 452, 335, 326]. Disjoint [543]. Disjointness [379]. Disjunction [554]. Disks [378]. Displacement [317]. Disruption [96]. Distance [267, 277, 282]. Distances [112]. Distortion [528]. Distributed [564, 560, 401, 638]. Distribution [401, 94]. Diversities [419]. Diversity [145]. Divine [249]. Do [479, 231]. Documentation [479]. Does [231, 443]. Dog [322]. Domain [85, 205, 200, 166, 470]. Dominance [8]. Dominant [606]. Dominate [109]. Dotted [22]. DPO [222]. Drawing [410]. Drawings [391, 26]. Dream [178]. Dreams [178]. Driven [530, 535, 211, 634, 476, 495, 353]. DS [624]. DTW [268]. Duals [27]. Duplicating [563]. Dynamic [472, 489, 284, 563, 265, 39, 359, 413, 315]. Dynamically [535]. Dynamically-Driven [535]. Dynamics [38, 269].

E-Collaborative [152]. E-Learning [244, 156]. Ear [597]. Earnings [123]. Easy [28]. EC [622]. EC-TEL [622]. ECCV [626, 627, 628, 631, 632, 633]. Echo [453]. Eclipse [473]. Ecological [588]. Ecosystems [492, 210]. Edge [559, 377]. Edges [219, 385, 377]. Education [167]. Educational [250, 164]. Effect [558]. Effective [325, 605]. Effects [301, 246]. efficacy [239]. Efficiency [49]. Efficient [363, 226, 174, 10, 422, 285, 550, 326, 315]. Efficiently [324]. Effort [481]. EGC [629]. EIA [588]. Eigen [57]. Eigenvalues [148]. Elastic [298]. Electric [397, 399]. Electricity [61]. Electrooculography [527]. Elements [415]. Embedded [530]. Embedding [267]. Emergence [137]. Emerging [74]. EMF [476]. Emotion [176]. Emotional [136]. Empirical [128]. Employee [121]. Employer [121]. Employing [395]. Emulating [52]. eMUSE [155]. Enabling [506]. Encoder [518]. Encryption [255]. End [563]. End-to-End [563]. Energetic [49]. Energy [432, 395, 353]. Energy-Driven [353]. Engineering [634, 247]. English [169, 134]. Enhanced [105, 248, 622].

Enhancements [502]. Enhancing [156, 241]. Enlargement [9]. Ensemble [417]. Ensuring [138]. Enterprises [124]. Entities [585]. Envelope [104]. Environment [247, 429]. Environments [137, 489, 87, 159, 160]. Equal [541]. Equations [513]. Equivalence [194]. Equivalent [200]. Err [249]. Erroneous [249]. ESOCC [635]. Estimate [103]. Estimating [301, 67]. Estimation [461, 520, 525, 526, 527, 311, 439, 405]. Estimators [339]. Euclidean [267]. Euler [512]. European [626, 627, 628, 635, 622, 104, 236]. Evaluating [562, 237, 481, 179, 471]. Evaluation [234, 110, 128, 165, 256, 245, 399, 419, 524, 439]. Evaluator [339]. Even [33]. Event [139, 135]. Events [292, 461]. Evolution [472, 296, 604, 210, 284]. Evolutionary [511, 210, 576]. Evolutionary-Neural [511]. Evolving [88, 559]. EWCVT [348]. Exact [173]. Examples [249]. Exceptions [496]. Excessive [122]. Exchange [587, 74]. Exchangers [583]. Executable [560, 197]. Exhaustive [363]. Expansion [132, 585]. Experiential [234]. Experiment [481]. Experiments [130]. Explain [249]. Exploit [151]. Exploiting [233, 212, 454]. Exploration [530, 556, 522, 62]. Explorative [516]. Exploratory [152]. Exploring [340, 276]. Exponential [25, 525]. Expression [81, 84]. Expressions [194]. Extended [267, 23, 97]. Extracted [353]. Extraction [139, 135, 291, 528]. Extreme [65]. Extrinsic [524]. Eye [358].

Facade [421]. FacadeMetamodel [469]. Face [417, 364, 571, 317, 271, 334]. Facebook [240]. Faces [345, 442, 411]. Facets [248]. Facial [419, 443, 444, 352, 407]. Facility [384, 98]. FactForge [145]. Factorization [361, 44]. Fall [454]. Falling [71]. Farms [102]. Fast [290, 447, 64, 362, 343, 288]. Fault [16]. Fault-Tolerance [16]. Faults [553]. Feature [582, 88, 190, 602, 483, 414]. Feature-Oriented [190]. Feature-Preserving [602]. Features [522]. Federated [160]. Feedback [562, 183]. Feedforward [66]. Femur [596]. Ferran [629]. Few [410]. Field [39, 317, 598]. Fields [364]. FIFO [563]. Fight [154]. Files [119]. Film [47]. Filter [521, 527]. Filtered [130, 55]. Filtering [418, 412, 131]. Filters [453, 315, 569]. Finalists [256]. Finder [175]. Finding [280, 148]. Finite [144]. Finite-State [144]. First [635]. Fish [65]. Fitting [419, 432, 365]. Flattening [380]. Fleet [92]. Flexible [111]. Flight [106]. Flippable [385]. Flipper [363]. Flips [375]. Florence [626, 627, 628, 631, 632, 633]. Flow [465, 437, 439, 462, 138, 441]. Flows [92]. FOAM [165]. Focus [219, 180]. Folksonomies [141]. Forbidden [21]. Forecast [516]. Forecasting [278, 61]. Forgetful [546]. Forgetting [558]. form [279]. Formal [151, 218, 561, 494, 637]. Formalising [253]. Formalism [261]. Formally [475]. FORMATS [637]. Formulas [279]. Formulations [346]. Forward [158]. Foundations [218, 190]. Four [517]. Four-Dimensional [517]. FPGA [48, 50]. Fractals [517]. Fracture [594]. Fraenkel [1]. Fragment [287]. Fragment-Based [287]. Fragmentation [474]. Frame [368, 318]. Frame-Rate [368]. Frames [458, 364]. Framework [299, 492, 297, 292, 473, 426, 209, 437, 95, 432, 371, 228, 515, 170]. Frameworks [238]. France [625, 610, 624, 640]. Free [19, 71]. Free-Falling [71]. Freeform [592]. Frequencies [546]. Frequency [99, 49]. Frequent [325]. Frog [264, 181]. Front [14, 35, 487, 508, 529, 549, 570, 591, 609, 56, 77, 91, 108, 129, 150, 171, 187, 208, 229, 230, 232, 252, 273, 294, 309, 330, 351, 372, 393, 403, 424, 445, 466]. Full [455, 100]. Full-Body [455]. Fun [181]. Function [58]. Functional [195, 143]. Funds [74]. Fusion [564, 175, 425]. Fuzzy [45].

Gait [176]. Game [248, 167]. Games [169, 4, 168, 250]. Gap [218]. Gas [514]. Gated [72]. Gates [47]. Gaussian [332]. GDP [278]. GEMBus [490]. Gender [177]. General [540, 221, 255, 366]. Generalised [224]. Generalization [390]. Generating [121, 484]. Generation [39, 517, 488, 602, 144]. Generations [401]. Generative [43]. Generator [245]. Genes [143]. Genomic [590]. Gentle [598]. Geo [421]. Geo-localization [421]. Geodesic [333]. Geometric [510, 435, 314, 365]. Geometrical [428]. Geometry [235, 629, 69]. Geomodeling [85]. Germany [621, 622]. Get [153]. Getting [182]. GHNNs [38]. Glioma [304]. Global [146, 587, 296, 557]. Globally [307]. Globally-Optimal [307]. GM [259]. GM-SOM [259]. Goes [122]. Gons [373]. GPU [73]. Gradient [58, 521]. Grain [348]. Grammar [89]. Grammars [218, 81, 84]. Graph [233, 226, 173, 261, 219, 379, 224, 218, 215, 193, 522, 341, 329, 214, 21, 209, 216, 552, 383, 605, 213, 228, 621, 611]. Graph-Based [233, 261, 209]. Graph-Theoretic [611]. Graphical [363]. Graphics [636]. Graphs [391, 27, 19, 20, 388, 284, 26, 28, 23, 22, 32, 140, 31, 217, 16]. Gray [521]. Grid [388, 556]. Ground [438]. Group [340, 409]. Group-Valued [409]. Groups [67]. Guarantee [565]. Guard [9, 377]. Guards [377]. Guide [575, 181]. Guided [538]. Guidelines [486].

Hallucination [369]. Handed [198]. Hanoi [25]. Hard [28]. Hardware [48]. Hardware-Core [48]. Harmful [122]. Harnessing [68]. Haskell [196]. HBU [619]. HCAC [286]. Healing [574]. Heat [406, 583]. Heaviest [389]. Heavy [500]. Hebbian [73, 47]. Held [625]. Henares [629]. Heuristic [98, 106]. Heuristics [545, 547]. HEVC [523]. Hierarchical [137, 602, 286, 164]. Hierarchically [39]. High [587, 592, 590, 368, 341, 560, 425]. High-Level [560]. Higher [192, 267, 352]. Higher-Order [192, 267, 352]. Highly [140]. Highway [387]. Hippocampus [296]. History [375]. HMM [292]. hoc [258]. Holes [373]. homogeneous [1]. Hop [565]. Hopfield [37]. Hough [349]. Households [395]. HSV [283]. Human [442, 348, 173, 174, 430, 310, 339, 181, 176, 429, 249, 172, 180, 619]. Human-Centric [429]. humanoid [520]. Hurtado [629]. Hybrid [8, 539, 540, 520, 531, 117, 45]. Hydrocephalus [604]. Hypervolume [572]. Hypotheses [328].

ICANN [612, 613]. ICARIS [639]. ICCL [615]. ICCVG [636]. ICGT [621]. ICWL [618]. Idempotent [200]. Identification [99, 260, 143, 455]. IFIP [623]. IFS [511]. IHSN [114]. II [632, 11, 613]. III [627, 633]. Image [346, 336, 512, 305, 297, 625, 325, 519, 360, 329, 640, 317, 319, 495, 362, 337, 370, 436, 515, 451, 369]. Images [511, 513, 348, 592, 521, 419, 528, 410, 328, 343, 355, 300, 596]. Imitation [185]. Immune [639, 581, 580, 577, 571, 574, 583, 586, 572, 576, 578]. Immune-Inspired [581, 574]. Immunology [575]. Impact [481, 96, 269, 49]. Imperfect [433]. Implementation [226, 193, 578]. Implicit [298]. Important [449]. Improve [151, 502]. Improved [312, 438, 523]. Improvements [499]. Improving [141]. In- [194]. In-Flight [106]. Incentive [399]. Incident [219]. Including [291]. Inclusion [114]. Inconsistency [481]. Increasing [389]. Incremental [205, 480]. Independent [32]. Index [179]. Individual [295, 246]. Indoor [429, 427]. Induced [21, 136]. Induction [542, 147]. Industrial [623, 494]. Inference [128, 119]. Inferred [145]. Infinite [475, 212]. Infinite-State [212]. Influence [239, 75, 40]. Inform [181]. Information [562, 233, 139, 135, 587, 623, 205, 293, 138]. Infrastructure [52, 394, 247]. Initial [234]. Initialization [514]. Innsbruck [634]. Input [12]. Inputs [125]. Insertion [106]. Insights [441]. Inspired [581, 574]. Instance [349]. Instead [122]. Integer [17]. Integrating [482]. Intelligence [139, 617]. Intelligent [522]. Intensive [530]. Inter [486]. Inter-association [486]. Interaction [185, 94, 180, 578]. Interactions [457]. Interactive [592, 249, 425, 468, 569]. Interconnection [476]. Interest [324]. Interesting [461]. Interestingness [143]. Interface [491]. Interface-Based [491]. Interfaces [235]. Interior [44]. Interior-Point [44]. Intermodal [95]. Internal [381, 184]. International [636, 122, 639, 623, 630, 616, 625, 621, 610, 634, 624, 611, 615, 637, 620, 640, 618, 617, 638, 619, 612, 613]. Internet [398]. Interpretability [279]. Interval [23, 22]. Intervals [15]. Intracardial [270]. Intracranial [515]. Introducing [580]. Introspection [86]. Intruder [259]. Intrusion [258, 584]. Invariant [406, 78, 463, 226, 225, 431]. Invariant-Typed [78]. Invariants [12, 197]. Inverse [345, 511, 183]. Investigation [113]. IPUMS [122]. IPUMS-International [122]. Iris [263]. Isometric [414]. Isomorphism [21]. Israel [611]. Issues [254, 161]. Italy [639, 623, 616, 626, 627, 628, 631, 632, 633, 635]. item [101]. Itemset [325]. Iterating [475].

Java [166]. Jerusalem [611]. Joining [349]. Joint [305, 426, 319, 334]. June [611, 629].

Kernel [43, 32, 59]. Kernelized [347]. Kernels [406]. Key [174]. Keystroke [269]. Killer [580]. Killing [79]. Kinect [455]. Kinects [434]. KITTI [440]. Knot [11]. Know [243]. Knowledge [145, 166, 557]. Known [600]. Krivine [2].

L [112]. Lab [241]. Label [60]. Labeling [305]. Lagrangian [98]. Lamellar [521]. Landmark [407]. Lanes [399]. Language [169, 151, 85, 133, 166, 478, 470, 257]. Languages [634, 614]. Lapse [307]. Large [391, 332, 52, 50, 240, 277, 329, 247, 313, 528, 337, 474, 451, 607, 420, 134]. Large-Scale [52, 50, 247, 451, 420]. Larger [480]. Laser [175]. Latent [136, 356]. Laughter [179]. Lausanne [612, 613]. Layered [207]. Lazy [363, 189]. Leap [264]. Learn [464]. Learning [278, 233, 235, 169, 165, 151, 239, 571, 58, 423, 243, 152, 324, 244, 245, 246, 70, 419, 323, 350, 73, 168, 247, 144, 161, 452, 279, 454, 64, 183, 40, 251, 71, 286, 231, 156, 159, 618, 155, 274, 622, 302, 327, 170, 451, 355, 612, 613, 47, 163, 342, 53, 160, 65]. Learning-Based [355]. Least [371]. Lectures [240]. Left [198, 81]. Left-Handed [198]. Length [525, 11]. Lessons [441]. Level [521, 560, 604, 408, 459]. Level-Set [408]. Lexical [132]. Lexicon [448]. LF [287]. LF-CARS [287]. LHS [352]. Library [261, 248]. Lie [310]. Likelihood [73, 366]. Limitations [99]. Limited [363]. Line [384, 58]. Linear [536, 26, 67, 456, 70, 3, 371]. Linear-Time [536]. Liner [92]. Lines [88, 484, 485]. Linguistic [477]. Link [163]. Linked [121, 504]. Linking [140]. Lip [268]. Literature [94]. Live [307]. Liveness [535]. LMS [158]. Load [61, 100]. Local [296, 30, 551, 352]. Localisation [460, 497]. Locality [456]. Locality-Constrained [456]. Localization [175, 322, 349, 407, 421]. Locating [384]. Location [422, 98, 304]. LOD [145]. Logic [188]. Logistic [118]. Logistics [99, 615, 102, 103]. London [637]. Long [63, 162]. Long-Term [63]. Longest [579]. Longitudinal [299, 305, 301, 625, 296, 295, 300]. Loose [287]. Loss [360, 103]. Loss-Specific [360]. Lot [101]. Lower [34]. LTI [12]. LTL [5]. LTL-Model-Checking [5]. Lyon [624].

m [568]. Machine [72, 495, 274, 356, 612, 613]. Machines [68, 46, 76, 561, 2, 80, 65]. Macros [89]. Magnitude [113]. Maintaining [381]. Make [152]. Malaria [578]. Manage [210]. Management [623, 161, 401, 494]. Manifold [310]. Manifolds [331]. Manpower [106]. Mapping [57]. Maps [222, 142, 126]. Margin [256]. Maritime [102]. Marked [442]. Markets [74]. marshalling [97]. Mashup [159]. Masking [469, 113]. Match [329]. Matched [119]. Matching [266, 421, 173, 438, 601, 317, 404, 413, 431, 420]. Matchings [196, 389]. material [602]. Math [249]. Mathematical [578]. Matrices [204, 545]. Matrix [42, 41, 343, 361, 44]. Matrix-Valued [343]. MATSim [400]. Matter [13, 90, 107, 186, 308, 402, 548, 608, 14, 35, 487, 508, 529, 549, 570, 591, 609, 56, 77, 91, 108, 129, 150, 171, 187, 208, 229, 230, 232, 252, 273, 294, 309, 330, 351, 372, 393, 403, 424, 445, 466]. Maximal [552]. Maximum [196, 23, 73, 366]. Mcta [547]. Mean [231]. Means [241]. Measure [143, 147]. Measure-Based [147]. Measurement [300]. Measures [440, 394, 454]. Mechanical [266]. Mechanism [398]. Mechanisms [478, 241]. Medial [510]. Median [288]. Mediation [504]. Medical [592, 640, 515]. Medium [124]. Meet [473]. Meeting [629]. Memory [54, 42, 422]. Mesh [593, 602, 599, 412, 604, 640]. Meshes [607]. Meshing [606, 596]. MeshMed [640]. Message [567]. Message-Passing [567]. Meta [468]. Meta-Modelling [468]. Metacognitive [234]. Metaheuristics [133]. Metamodeling [210]. Metastatic [305]. Method [260, 521, 41, 524, 453, 603, 604, 118]. Methodology [165, 248, 617]. Methods [122, 116, 620, 542, 114, 362]. Metric [418, 28]. MICCAI [625, 640]. Microdata [122, 116, 117, 113]. Microenvironments [307]. Microtomography [521]. Mid [459]. Mid-level [459]. Middle [249]. Migrating [561]. Minimal [76, 551, 338, 557]. Minimally [122]. Minimization [432, 366]. Minimum [20, 30]. Mining [275, 325, 274]. Mis [97]. Mis-overlay [97]. Misdetection [341]. Mix [41]. Mix-Matrix [41]. Mixed [301, 126]. Mixed-Effects [301]. Mixture [290]. Mixtures [332]. Mobile [501, 240, 175, 559, 555, 283]. modal [426]. Model [290, 530, 411, 87, 449, 539, 5, 10, 590, 225, 203, 520, 303, 37, 341, 227, 419, 133, 531, 567, 495, 117, 93, 396, 588, 480, 432, 293, 474, 113, 257, 365, 634]. Model-Based [87, 227, 117, 113]. Model-Driven [530, 495]. Modeling [442, 510, 500, 592, 477, 281, 577, 482, 72, 484, 637, 166, 429, 486, 274, 47, 494]. Modelling [492, 473, 140, 468]. MODELS [634, 290, 472, 363, 52, 475, 593, 301, 67, 473, 212, 221, 601, 476, 467, 360, 70, 450, 335, 464, 474, 483, 334]. Moderates [239]. Modified [265, 65]. Modular [426]. Modulated [53]. Modules [88]. Molecules [140]. Monitoring [505]. Monocular [435, 454]. Monotone [381]. Moodle [158]. Morphable [317]. Morphing [408]. Morphometry [595]. Most [571]. Mostowski [1]. Motion [463, 297, 520, 350, 306, 464, 455, 409, 405]. Motion-Based [306]. Motivational [244]. Mounted [175]. Move [464]. Movements [177, 358]. Moves [11, 365]. Movies [307]. MP [567]. MP-State [567]. MPEG [518]. MPEG-4 [518]. MR [300]. MRI [299]. Much [493]. Multi [465, 204, 25, 68, 174, 430, 630, 537, 397, 394, 101, 602, 318, 426, 404, 271, 307, 572, 483, 576, 505]. Multi-agent [630]. Multi-Agent-Based [394]. Multi-constraints [271]. Multi-core [537]. Multi-item [101]. Multi-material [602]. Multi-modal [426]. Multi-Objective [572, 576]. Multi-part [404]. Multi-peg [25]. Multi-person [318]. Multi-perspectives [483]. Multi-photon [307]. Multi-round [397]. Multi-sized [465]. Multi-tempering [68]. Multi-temporal [307]. Multi-tenant [505]. Multi-unit [397]. Multi-valued [204]. Multi-view [174, 430]. Multichannel [348]. Multidimensional [516]. Multifeature [444]. Multifocal [528]. Multilabel [367]. Multilayer [60, 57]. Multiobjective [74]. Multiple [418, 340, 124, 23, 460, 589, 434]. Multiset [43]. Mutual [340]. My [339].

Natal [614]. Natural [580, 133, 355]. Near [414]. Near-Isometric [414]. Negative [214, 584, 573]. Neonates [298]. Nested [76, 366]. Net [96]. Nets [9, 532, 540, 577]. Network [48, 50, 554, 236, 583, 40, 401, 138, 163]. Networks [54, 137, 146, 52, 258, 563, 260, 501, 36, 541, 460, 453, 59, 574, 313, 66, 545, 94, 63, 550, 69, 612, 613, 45]. Neural [52, 511, 514, 50, 73, 66, 63, 69, 612, 613, 45, 47]. Neuronal [54, 48]. Neurons [54, 51, 55, 62, 45]. Newly [102]. News [136]. Newton [66]. Nice [625, 640]. Nitride [47]. Noise [111]. Non [415, 569, 266, 1, 360, 520, 460, 181, 197, 371, 370, 436, 334, 563]. Non-dense [436]. Non-determinism [197]. non-FIFO [563]. Non-homogeneous [1]. Non-interactive [569]. Non-overlapping [460]. Non-Parametric [360, 334]. Non-Rigid [415, 266, 371, 370]. Non-verbal [181]. Nonlinear [8]. Nonmetric [367]. Nonnegative [44]. Nonseparating [32]. Norms [395]. Notebook [240]. Notes [383]. Novel [587, 95]. November [630]. Nucleotide [282]. Number [382, 379, 134]. Numerically [338].

Object [406, 316, 418, 340, 433, 423, 328, 434]. Objective [572, 576]. Objects [410, 446]. Oblivious [568, 556, 555]. Observability [544]. Observation [538]. Observations [203]. Obstacle [427]. Occasion [629]. Occlusion [312]. OCL [482]. October [625, 626, 627, 628, 634, 631, 632, 633, 624, 640, 638, 619]. Offshore [102]. Omega [199]. Omega-Regular [199]. Omitting [563]. On-Line [58]. One [540]. One-Shot [540]. Online [347, 152, 453, 66, 138, 342, 160, 289]. Online-Teaching [152]. Online/Batch [66]. Only [410]. Ontologies [140]. Ontology [139, 261, 166]. Ontology-Based [139]. Open [503, 222, 162, 377]. OpenSocial [158]. Operations [95]. Operator [583, 149]. Opportunities [99]. Optical [437, 344, 441]. Optimal [556, 313, 307, 69]. Optimally [534]. Optimisation [74, 327]. Optimization [587, 110, 76, 22, 41, 338, 531, 583, 367, 576, 65]. Optimized [518, 45]. Optimizing [85, 95]. Option [502]. Oracles [602]. Orchestration [237, 488, 241]. Order [192, 267, 67, 352]. Orderings [382]. Organization [205, 328]. Organizational [246]. Organized [39]. Organs [510]. Oriented [503, 190, 635]. Oscillations [38]. Other [122]. Our [248]. Outdoor [181]. Outlier [264]. Output [127]. Over- [212]. overlapping [460]. overlay [97]. Overview [99]. Oxide [47].

PaaS [490, 503]. PaaSSOA [503]. Pair [53]. Pair-Associate [53]. Palermo [616]. Paper [235]. Papers [130, 630, 611, 629]. Paradigm [590]. Parallelism [215]. Parallelization [68]. Parallelohedra [374]. Parameter [83, 70, 362]. Parameterization [525]. Parameterized [33]. Parameters [70, 133]. Parametric [360, 149, 334]. Pareto [69]. Pareto-Optimal [69]. Parsing [81, 84]. Part [626, 627, 628, 631, 632, 633, 450, 322, 328, 612, 613, 404]. Partial [601, 484, 544]. Partiality [197]. Particle [74, 607]. Particle-Based [607]. Particularity [323]. Partition [603]. Partitioning [552]. Parts [449, 448]. Passing [567]. Patch [336, 601]. Path [146, 163]. Path-Consistency [146].

Paths [389]. Patient [135]. Pattern [275, 267, 291, 479, 248, 263, 213]. Pattern-Based [213]. Patterns [284, 36, 37]. Payments [24]. PCA [73]. PDE [362]. PDE-Based [362]. PDEs [346]. Pedestrian [447]. peg [25]. Pelvic [510]. [426]. Batch [66]. H.264 [518]. IE [587]. People [175]. Per-frame [318]. Per-track [318]. Perceived [152]. Perception [177]. Perceptron [60, 57]. Perceptual [328]. Perfect [20]. Performance [500, 524]. Permutation [389]. Persistent [507, 414]. Persisting [474]. Person [455, 318]. Personal [247, 159]. Personalised [170]. Personalized [244]. Personalizing [141]. Persons [260]. Perspective [536, 236]. perspectives [483]. Perturbation [122, 419, 127]. PET [303]. Petri [9, 532, 540, 577, 96]. Pharmacokinetic [303]. photon [307]. Pieces [404]. Piecewise [526]. Piecewise-quadratics [526]. Piles [7]. Pipeline [515]. Pipelined [501]. Pixels [446]. Placement [551, 80]. Planar [391, 27, 26, 382, 386, 32]. Plane [438, 427]. Planning [102, 93]. Plasticity [53]. Platform [490]. Playback [354]. Playing [534, 4]. PLE [158]. Point [442, 99, 382, 202, 386, 311, 126, 389, 427, 44]. Points [324, 414]. Poland [636]. Policies [399]. Policy [394, 279, 71, 257, 505]. Policy-Based [71]. Polygon [607]. Polygonization [386]. Polygons [381, 377]. Polyhedra [390, 380]. Polynomial [331, 29, 338, 45]. Population [296, 557, 408]. Port [96]. Ports [94]. Portugal [619]. Pose [339, 317, 311]. Poses [174]. Positions [282]. Post [111, 118]. Post-tabular [111]. Poster [18]. Potentials [49]. Power [397, 401]. Powerset [201]. Practical [562, 269]. Practice [218]. Pre [97]. Pre-marshalling [97]. Precision [82]. Predicates [538]. Predicting [292, 304]. Prediction [278, 418]. Predictive [95]. Preferences [207]. Presence [553]. Preservation [225, 561]. Preserved [593]. Preserving [602]. Pressure [272]. PRIMA [630]. Print [268]. Priors [367, 333]. Privacy [122, 616, 616]. Private [569, 128]. Probabilistic [539, 10, 586, 359, 361]. Probabilities [10]. Problem [511, 387, 23, 41, 98, 288, 92, 24]. Problems [25, 587, 105, 610, 6, 22, 338, 3]. Proceedings [626, 627, 628, 631, 632, 633, 612, 613, 636, 639, 623, 616, 625, 621, 610, 634, 624, 615, 614, 637, 620, 640, 635, 618, 617, 622, 638, 619]. Processes [442, 453]. Processing [133, 640, 362, 270, 515]. Product [88, 485, 484]. Production [101, 162]. Productivity [103]. Products [477]. Professional [479]. Profitability [104]. Program [196]. Programming [87, 190, 166, 285, 413, 315, 614]. Programs [562, 195, 566, 191]. Progressive [290]. Project [245, 251, 155]. Project-Based [251, 155]. Projections [326]. Projects [236]. Promoting [399]. Propagation [562, 83, 337]. Properties [562, 37]. Proposed [125]. Protect [122]. Protection [115]. Protocol [496, 507]. Protocols [557]. Prototype [267, 519]. Provider [493]. Proving [535]. Provisioning [495]. PSD [616]. Pseudo [205]. Publications [94]. Pushdown [2]. Putting [404].

Quad [606]. Quadratic [517, 41, 285]. quadratics [526]. Qualitative [146]. Quality [593, 269, 498, 436, 147]. Quantales [201]. Quantiles [278].

Quantitative [88, 467]. Quantization [514]. Quasi [26, 541, 66]. Quasi-Equal [541]. Quasi-Newton [66]. Query [476]. Query-Driven [476].

Radial [382, 59]. Radio [99]. RAMiCS [620]. Ramp [292]. Random [534, 364, 127, 598]. Randomization [118]. Range [82, 175]. Ranged [10]. Ranking [60]. Rapid [384, 450, 470]. Rate [368]. Rates [425]. Ratio [461]. Rational [220]. Ray [521]. RBFNN [75]. RC [582]. RC-DCA [582]. Reachability [8, 17, 537, 6, 610]. Reaction [36]. Reactive [489, 467]. Real [415, 48, 173, 191, 368, 419, 354, 586, 457, 607, 427]. Real-Time [415, 48, 173, 191, 368, 607, 427]. Real-World [419, 354]. Realboosted [465]. Realistic [600]. Reasoning [192]. Recognition [428, 173, 174, 267, 357, 265, 433, 571, 347, 422, 318, 317, 263, 268, 320, 184, 358]. Recognize [176]. Recognizing [180]. Recommendations [233, 486]. Recommender [159, 160]. Reconstruction [415, 312, 513, 512, 315, 405]. Recording [354]. Records [135]. Recovery [597]. Rectangle [195]. Rectangle-Based [195]. Rectangular [205]. Recurrence [3, 304]. Recursion [81]. Recursive [2]. Red [392]. Redescriptions [280]. Reduced [285]. Reducibility [57]. Reducing [541]. Reduction [594]. Redundancy [285]. Reeb [605]. Refined [216]. Refinement [203, 531]. Reflexive [195]. Reflexive-Transitive [195]. Region [540]. Region-Based [540]. Regions [349]. Registration [305, 295, 303, 298, 371]. Regression [331, 370, 147, 407, 118]. Regular [192, 199, 194]. Regularization [46, 295, 343, 409]. Regularized [404, 289]. Regulate [251]. Reidemeister [11]. Reinforcement [279, 71]. Related [595]. Relation [204, 193]. Relational [280, 201, 206, 620]. Relations [196, 340, 291, 197, 131]. Relationship [6]. Relationships [164]. Relative [205, 460, 452]. Relaxing [471]. Release [113]. Releases [124]. Relevant [138]. Relies [122]. Removal [599, 264]. Rendering [345, 420]. Rendezvous [559]. Reo [488]. Reorientation [297]. Repeatable [324]. Replicas [551]. replicating [585]. Replication [479]. Repositioning [92]. Representation [388, 310, 456, 201, 459, 607]. Request [499]. Requirements [473, 247]. Resection [304]. Resilient [16]. Resistance [277]. Resolution [336, 411, 426]. Resource [162, 498, 494]. Resources [233]. Restoration [360]. Restricted [68, 46]. Retail [99]. Rethinking [589]. Retries [499]. Retrieval [266, 319]. Reuse [162]. Reveal [243]. Revenue [499]. Reversibility [374]. Reviews [134]. Revised [630, 611, 629]. Revisited [220]. Revisiting [536]. Rewriting [220, 216]. Riemannian [331]. Rigid [415, 266, 16, 371, 370]. Rings [559, 555]. Risks [493]. Robot [182, 175, 180, 283]. Robotic [181]. Robotics [172]. Robots [556, 555]. Robust [287, 592, 433, 601, 66, 361, 342]. Robustness [9, 532, 341]. Role [262]. Romania [618]. Roommates [24]. Rotating [378]. Rotation [463, 263]. Rough [582]. round [397]. Rounding [111]. Routing [105, 100, 400]. RP [610]. Rule [42, 143, 147]. Rule-Based [143]. Rules [480, 147]. Run [471]. Runs [545]. Runtime [472, 473, 499, 86].

Saarbrücken [622]. Safety [638]. Saliency [358, 333]. Sample [571]. Sampling [122, 353, 358]. Sand [7]. SAT [256]. SAT-Based [256]. SBLP [614]. Scaffolding [234]. Scalable [592]. Scale [406, 332, 52, 50, 423, 277, 247, 313, 314, 451, 420]. Scales [291]. Scans [305]. Scenarios [253, 245]. Scene [428, 350, 439]. Scenes [427]. Scheduling [106]. Scheme [255]. Schemes [2]. School [249]. Schrödinger [416]. Science [79, 624, 611, 620]. Scientific [125]. Scientometric [94]. Scores [450]. Script [320]. sdcMicro [114]. Search [363, 141, 279, 407, 57, 547]. Searching [571]. Seaworthy [93]. Second [625]. Security [253, 254, 492, 256, 561, 161, 638, 493, 127]. SEEDS [353]. Segment [379, 521, 350]. Segmentation [348, 347, 426, 306, 307, 349, 427, 300, 598]. Segmentations [423]. Selcted [611]. Selected [630, 629]. Selection [290, 489, 587, 582, 590, 584, 184, 573, 69]. Self [554, 563, 564, 239, 565, 551, 574, 553, 552, 555, 550, 585]. Self- [565]. Self-efficacy [239]. Self-replicating [585]. Self-stabilization [555, 550]. Self-stabilizing [554, 563, 564, 565, 551, 553, 552]. Semantic [233, 136, 131]. Semantics [458, 228]. Semi [454, 286]. Semi-supervised [454, 286]. Semiconductor [47]. Semirings [200]. Sense [319]. Sensitivity [362]. Sensor [272, 574, 425]. Separating [323]. Separation [188]. Separator [19]. September [636, 623, 616, 621, 610, 634, 615, 614, 637, 620, 635, 618, 617, 622, 612, 613]. Sequence [293, 573]. Sequences [297, 39, 282, 3, 459]. Sequential [291, 76, 276]. Serendipitous [138]. Series [625]. Serious [248]. Serpentine [386]. Server [500]. Service [489, 506, 145, 384, 503, 565, 497, 491, 106, 504, 499, 635]. Service-Oriented [635]. Services [492, 490, 496, 488, 209]. Session [18]. Sessions [241]. Set [582, 386, 32, 604, 408]. Sets [1, 382, 389, 147]. SHA [256]. SHA-3 [256]. Shake [354]. Shanghai [615]. Shape [266, 593, 301, 296, 310, 517, 601, 416, 404, 335, 605, 408, 371, 302, 334, 407, 595]. Shared [507]. Sharpness [526]. Shipping [92]. Shot [540]. Should [109]. Signal [582, 44]. Signals [463, 55, 527, 178]. Signature [265]. Significantly [478]. Silent [554]. Silicon [51, 47]. SIMD [80]. Similar [571, 130]. Similarity [569, 287, 356]. Simple [195, 279, 377]. Simplified [605]. Simplifying [64]. Simulate [515]. Simulation [562, 630, 575, 394, 599, 462, 395, 103, 400]. Simulations [48, 184]. Simulator [502]. Simulators [234]. Simultaneous [101]. Simultaneously [385]. Sinaia [618]. Single [336, 540, 371]. Singular [267]. Sites [103]. Six [248]. sized [465]. Sizing [101]. Skeleton [520]. Sketch [410]. Skills [231, 622]. Skull [601]. Skyline [420]. Slicepedia [162]. Slots [568]. Small [124]. Smart [254, 630]. Smooth [7]. Smoothness [438]. SMT [566, 542]. SMT-Based [542]. Sobolev [418]. Sobolev-Type [418]. Social [137, 236, 151, 244, 157, 168, 247, 457, 395, 206, 138]. Society [630]. Soft [476, 599]. Software [78, 530, 477, 88, 567, 478, 515]. Software-Intensive [530]. Solution [125]. Solutions [24]. Solvers [110, 566, 277, 338]. Solving [511, 105, 98, 100]. SOM [259]. Some [59, 557]. Source [390, 479]. Space [530, 511, 458, 29, 70, 279, 358]. Space-Time [458]. Space-Variant [358]. Spain [629]. Span [29, 216]. Span-Categories [216]. Spanish [629]. Sparking [246]. Spatial [291, 303, 126]. Spatial-temporal [303]. Spatio [456, 295, 444, 625]. Spatio-Temporal [444, 456, 295, 625]. Special [399]. Specific [85, 360, 464, 470]. Specification [536, 257]. Specifications [560, 197]. Spectral [277, 412, 44]. Speech [176]. Speed [82]. Spherical [601]. Spielman [277]. Spike [53]. Spike-Time [53]. Spiking [54, 50, 47]. Spill [80]. Spiral [386]. Spirometer [272]. Split [47]. Sports [357]. Spreadsheets [87]. Spyware [580]. Squares [371]. Squashing [283]. SRE [247]. SSS [638]. Stability [590]. Stability-Based [590]. stabilization [555, 550, 638]. Stabilizing [562, 566, 554, 563, 564, 565, 551, 553, 552]. Stable [412, 338, 31, 24]. Stage [98]. State [212, 360, 453, 144, 567]. State-Aware [567]. States [136]. Static [545]. Statistical [278, 128, 616, 539, 119, 597, 531, 313, 408, 595, 127]. Statistics [352]. Stent [515]. Step [158]. Stereo [438, 440, 435, 425]. Stereotype [396]. STIA [625]. Stixel [447]. Stixels [446]. Stochastic [513, 111, 215, 531, 40, 62]. Storage [54, 101]. Stowage [93]. Strategies [400]. Stream [292, 589]. Stream-Based [292]. Striatum [605]. Strict [189]. String [378]. String-Wrapped [378]. Strip [20]. Strip-Composed [20]. Structural [299]. Structure [482, 255, 183, 40]. Structured [43, 214, 453, 342]. Structures [16, 606]. Student [239, 18]. Students [240, 249]. Studies [408]. Study [105, 75, 152, 597, 249, 149, 573, 494]. Style [243]. Sub [484]. Sub-product [484]. Subcortical [600]. Subgraphs [21]. Subgroup [285]. Subimages [283]. Subject [118]. Subjective [143]. Subquadratic [391]. Subsequence [389]. Subsequences [579]. Succeeded [339]. Suitable [260]. Sum [17]. Summarization [509]. Super [336, 411, 426]. Super-Resolution [336, 411, 426]. Superalloy [348]. Superfaces [411]. SuperFloxels [459]. Superpixels [353]. Supervertices [603]. Supervised [75, 423, 356, 454, 286]. Supply [96]. Support [76, 152, 102, 478, 356]. Suppression [122]. Surface [600, 597, 298]. Surfaces [415, 312, 603, 414, 431, 596]. Surgery [304]. Surveillance [313, 457]. Survey [124, 120]. Sustainable [630, 397]. Swapping [122, 120]. Swarm [74, 148, 65]. Switzerland [612, 613]. Symbolic [226]. Symmetrical [140]. Symmetry [355]. Symposium [614, 638]. Synchronization [48, 553, 545]. Synchronizer [558]. Synchronous [152]. Synthesis [336, 538, 600, 142, 544]. Synthetic [441]. System [258, 580, 577, 102, 528, 454, 271, 586, 572, 268, 160, 86, 578]. Systemic [281]. Systems [472, 530, 535, 50, 639, 623, 539, 575, 634, 10, 221, 223, 12, 637, 542, 214, 531, 567, 617, 638, 547].

T [470]. Tables [128]. Tabular [113, 111]. Tag [141]. Tag-Based [141]. Tail [162]. Taming [197]. Tangled [376]. Tanks [93]. Taormina [639]. Tasks [501]. Taxonomic [116]. TC [623]. Teaching [152, 153, 157]. Team [357]. Technique [582, 265, 604]. Techniques [509, 40, 66]. Technological [246]. Technologies [237]. Technology [99, 523, 622]. TEL [622, 236]. Telling [178]. tempering [68]. Template [295]. Temporal [146, 135, 438, 36, 347, 444, 456, 295, 303, 307, 625]. tenant [505]. Teng [277]. Tensor [346]. Tensor-Based [346]. Tensors [267]. Term [220, 58, 63]. Terminal [95]. Terminals [94]. Terminating [557]. Test [121]. Testing [227, 114]. Tetrahedral [607]. Textual [134]. Textural [428]. Texture [72, 416, 314, 352]. Their [346, 493]. Theorem [201]. Theoretic [611]. Theoretical [58]. Theories [536]. Theory [533, 582]. Thick [510]. Thickness [300]. Thin [47]. Third [615, 619]. Thrackles [376]. Throughput [590]. TIDS [258]. Tight [34]. Tikhonov [46]. Tikhonov-Type [46]. Time [415, 9, 532, 48, 458, 173, 38, 536, 191, 625, 368, 522, 586, 307, 471, 100, 457, 607, 427, 53]. Time-Lapse [307]. Time-Series [625]. Time-Travel [522]. Time-Varying [38]. Timed [533, 534, 535, 536, 537, 6, 541, 637, 542, 543, 544, 546, 547]. TinkerLamp [237]. Tissue [599, 307]. Titanium [521]. TOF [425]. Together [404]. Togetherness [210]. Token [507]. Tolerance [562, 16]. Tomography [405]. Tool [250, 156, 154]. Topology [606]. Toronto [638]. Tower [25]. Track [271, 318]. Tracking [418, 305, 340, 368, 341, 464, 307, 604, 342]. Trade [99]. Traded [74]. Traditional [435]. Trails [17]. Training [290, 234, 68, 360, 66]. Trajectories [460]. Trajectory [526, 315]. Transfer [239]. Transformation [219, 224, 225, 221, 223, 222, 193, 214]. Transformations [621, 215, 11, 228]. Transistors [47]. Transit [384]. Transition [540]. Transitive [195, 188]. Transparent [474]. Transport [394]. Transversal [33]. Trap [580]. Travel [522]. Treewidth [26]. Triangles [552]. Triangulation [19]. Triangulations [375, 385]. Trim [282]. Triple [218]. Truly [228]. Trust [258, 161]. Trust-Based [258]. Tumors [305]. Tuning [133]. Tweets [243, 293]. Twisted [383]. Twitter [243]. Two [511, 458, 203, 133, 21, 98, 545]. Two-Dimensional [511]. Two-Stage [98]. Type [418, 46, 86]. Typed [78].

Ubiquitous [253]. UK [637, 620]. Ultra [421]. Ultra-wide [421]. UML [469]. UML2 [486]. Unbiased [295]. Uncertain [12]. Uncertainty [471]. Unconstrained [364]. Under-Approximations [212]. Understanding [238, 172, 619]. Undoing [321]. UNESCO [616]. Unfolding [390]. Unified [432]. Uniform [149, 555]. Unifying [189]. unit [397]. Units [59]. Universal [167]. Unmanned [519]. Unmixing [44]. Unordered [147]. Unstructured [607, 134]. Unsupervised [461, 452, 302]. UPPAAL [545]. Upward [27]. Usage [486, 262, 257]. Use [325, 240, 152]. Useful [121, 154]. User [502, 505]. User-Centric [502]. User-Customisable [505]. Using [406, 415, 465, 272, 42, 442, 8, 348, 105, 593, 196, 74, 364, 501, 397, 259, 461, 460, 602, 277, 453, 311, 176, 322, 197, 410, 455, 66, 286, 298, 327, 407, 62, 315, 333, 47]. Utility [149]. Utilizing [57].

Vague [178]. Vagueness [178]. Valid [119]. Value [267, 398]. Value-Based [398]. Valued [513, 343, 409, 57, 204]. Variable [545]. Variables [67, 118]. Variance [551]. Variant [358]. Variational [346]. Varna [617]. Varying [38]. Vascular [606]. Vector [513, 76, 356, 57]. Vector-Valued [513]. Vehicle [105, 519, 100, 400]. Vehicles [399]. Venice [623]. verbal [181]. Verification [566, 225, 561]. Vertebral [595]. Very [134]. Vessel [93]. Vessels [515]. via [418, 305, 5, 225, 601, 193, 285, 557, 117, 604, 408, 337, 11, 353, 431]. Video [509, 364, 456, 518, 423, 523, 350, 459, 457]. Videos [429, 455]. View [99, 174, 430]. Viewpoint [431]. Views [142]. VII [628]. Vilamoura [619]. Virtual [561, 495]. Visibility [381]. Vision [636, 626, 627, 628, 631, 632, 633]. Visual [227, 452, 180, 451]. Visualization [516]. Visualizations [244]. Vocabularies [332]. Volume [299]. Volumetric [607]. Voronoi [392]. Voting [268]. Voting-Based [268]. vs [296, 16].

Warehouse [101]. Warping [604]. Warsaw [636]. Weakly [423, 356]. Web [618, 165, 423, 488, 170]. Web-Based [618, 165]. Web-Scale [423]. Wedge [415]. Weekday [61]. Weekday-Based [61]. Weight [565]. Weight-Based [565]. Weighted [20, 28, 37, 389, 342]. Weights [484]. Well [214]. Well-Structured [214]. WG [611]. Which [25, 297, 152]. While [471, 381]. Wiberg [366]. wide [421]. Wiki [154]. Wikis [153]. Wind [102]. Window [465]. Windows [265, 100]. Wireless [258, 574, 550]. within [291]. without [585]. Wollongong [630]. Word [319]. Wordnet [132]. Workbench [135]. Workflow [470]. Workload [500]. Workshop [625, 610, 611, 640, 619]. Workshops [630, 631, 632, 633]. World [419, 354]. Worldwide [311]. Wrapped [378]. Wright [105]. Wrinkles [442]. Write [500]. Write-Heavy [500]. Wrong [122].

X [521]. X-Ray [521]. XACML [257]. XIV [629].

Zeno [545].

References

Bojanczyk:2012:FMS

[1] Mikołaj Bojańczyk and Sławomir Lasota. Fraenkel–Mostowski sets with non-homogeneous atoms. Lecture Notes in Computer Science, 7550:1– 5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

     tronic). URL http://link.springer.com/chapter/10.1007/978-3-
     642-33512-9_1/.

Salvati:2012:RSK

[2] Sylvain Salvati and Igor Walukiewicz. Recursive schemes, Krivine machines, and collapsible pushdown automata. Lecture Notes in Computer Science, 7550:6–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33512-9_2/.

Ouaknine:2012:DPL

[3] Joël Ouaknine and James Worrell. Decision problems for linear recurrence sequences. Lecture Notes in Computer Science, 7550:21–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33512-9_3/.

Kucera:2012:PGC

[4] Antonín Kučera. Playing games with counter automata. Lecture Notes in Computer Science, 7550:29–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33512-9_4/.

Felscher:2012:LMC

[5] Ingo Felscher. LTL-model-checking via model composition. Lecture Notes in Computer Science, 7550:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33512-9_5/.

Haase:2012:RBR

[6] Christoph Haase, Joël Ouaknine, and James Worrell. On the relationship between reachability problems in timed and counter automata. Lecture Notes in Computer Science, 7550:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33512-9_6/.

Brocchi:2012:SSP

[7] Stefano Brocchi and Paolo Massazza. Smooth sand piles. Lecture Notes in Computer Science, 7550:66–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33512-9_7/.

Benvenuti:2012:ADC

[8] Luca Benvenuti, Davide Bresolin, Pieter Collins, and Alberto Ferrari. Ariadne: Dominance checking of nonlinear hybrid automata using reachability analysis. Lecture Notes in Computer Science, 7550:

79–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33512-9_8/.

Akshay:2012:RTPa

[9] S. Akshay, Loïc Hélouët, Claude Jard, and Pierre-Alain Reynier. Robustness of time Petri nets under guard enlargement. Lecture Notes in Computer Science, 7550:92–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33512-9_9/.

Ghorbal:2012:EPM

[10] Khalil Ghorbal, Parasara Sridhar Duggirala, and Vineet Kahlon. Efficient probabilistic model checking of systems with ranged probabilities. Lecture Notes in Computer Science, 7550:107–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33512-9_10/.

Saleh:2012:LKT

[11] Rafiq Saleh. On the length of knot transformations via Reidemeister moves I and II. Lecture Notes in Computer Science, 7550:121–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33512-9_11/.

Hansch:2012:ILS

[12] Paul Hänsch and Stefan Kowalewski. Invariants for LTI systems with uncertain input. Lecture Notes in Computer Science, 7550:137–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33512-9_12/.

Anonymous:2012:BMa

[13] Anonymous. Back matter. Lecture Notes in Computer Science, 7550: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33512-9/1.

Anonymous:2012:FMa

[14] Anonymous. Front matter. Lecture Notes in Computer Science, 7550: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33512-9/1.

Rautenbach:2012:AI

[15] Dieter Rautenbach. Account on intervals. Lecture Notes in Computer Science, 7551:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-34611-8_1.

Peleg:2012:CRS

[16] David Peleg. Constructing resilient structures in graphs: Rigid vs. competitive fault-tolerance. Lecture Notes in Computer Science, 7551:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34611-8_2.

Bhattacharya:2012:ARI

[17] Amitava Bhattacharya. Alternating reachability and integer sum of closed alternating trails. Lecture Notes in Computer Science, 7551:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-34611-8_3.

Golumbic:2012:SPS

[18] Martin Charles Golumbic, Michal Stern, Avivit Levy, and Gila Morgenstern. Student poster session. Lecture Notes in Computer Science, 7551:4–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34611-8_4/.

Berry:2012:TCS

[19] Anne Berry and Annegret Wagler. Triangulation and clique separator decomposition of claw-free graphs. Lecture Notes in Computer Science, 7551:7–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34611-8_5/.

Bonomo:2012:MWC

[20] Flavia Bonomo, Gianpaolo Oriolo, and Claudia Snels. Minimum weighted clique cover on strip-composed perfect graphs. Lecture Notes in Computer Science, 7551:22–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34611-8_6/.

Kratsch:2012:GIG

[21] Stefan Kratsch and Pascal Schweitzer. Graph isomorphism for graph classes characterized by two forbidden induced subgraphs. Lecture Notes in Computer Science, 7551:34–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34611-8_7/.

Hermelin:2012:OPD

[22] Danny Hermelin, Julián Mestre, and Dror Rawitz. Optimization problems in dotted interval graphs. Lecture Notes in Computer Science, 7551: 46–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34611-8_8/.

Francis:2012:MCP

[23] Mathew C. Francis, Daniel Gonçalves, and Pascal Ochem. The maximum clique problem in multiple interval graphs (extended abstract). Lecture Notes in Computer Science, 7551:57–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34611-8_9/.

Biro:2012:SSR

[24] Péter Biró, Matthijs Bomhoff, Petr A. Golovach, Walter Kern, and Daniël Paulusma. Solutions for the Stable Roommates Problem with payments. Lecture Notes in Computer Science, 7551:69–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34611-8_10/.

Berend:2012:WMP

[25] Daniel Berend and Amir Sapir. Which multi-peg Tower of Hanoi problems are exponential? Lecture Notes in Computer Science, 7551:81– 90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34611-8_11/.

DiGiacomo:2012:QPD

[26] Emilio Di Giacomo, Walter Didimo, Giuseppe Liotta, and Fabrizio Montecchiani. h-quasi planar drawings of bounded treewidth graphs in linear area. Lecture Notes in Computer Science, 7551:91–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34611-8_12/.

Auer:2012:DUP

[27] Christopher Auer, Christian Bachmaier, Franz J. Brandenburg, and Andreas Gleißner. The duals of upward planar graphs on cylinders. Lecture Notes in Computer Science, 7551:103–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34611-8_13/.

Epstein:2012:WMD

[28] Leah Epstein, Asaf Levin, and Gerhard J. Woeginger. The (weighted) metric dimension of graphs: Hard and easy cases. Lecture Notes in Computer Science, 7551:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34611-8_14/.

Junosza-Szaniawski:2012:DSP

[29] Konstanty Junosza-Szaniawski, Jan Kratochvíl, Mathieu Liedloff, and Paweł Rzażewski. Determining the L(2, 1)-span in polynomial space. Lecture Notes in Computer Science, 7551:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-34611-8_15/.

Javelle:2012:MDL

[30] Jérôme Javelle, Mehdi Mhalla, and Simon Perdrix. On the minimum degree up to local complementation: Bounds and complexity. Lecture Notes in Computer Science, 7551:138–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34611-8_16/.

Muller:2012:SDG

[31] Haiko Müller. On the stable degree of graphs. Lecture Notes in Computer Science, 7551:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34611-8_17/.

Kowalik:2012:KKN

[32] Łukasz Kowalik and Marcin Mucha. A 9 k kernel for nonseparating independent set in planar graphs. Lecture Notes in Computer Science, 7551: 160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34611-8_18/.

Misra:2012:PAE

[33] Pranabendu Misra, Venkatesh Raman, M. S. Ramanujan, and Saket Saurabh. Parameterized algorithms for even cycle transversal. Lecture Notes in Computer Science, 7551:172–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34611-8_19/.

Mnich:2012:BAT

[34] Matthias Mnich and Rico Zenklusen. Bisections above tight lower bounds. Lecture Notes in Computer Science, 7551:184–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34611-8_20/.

Anonymous:2012:FMb

[35] Anonymous. Front matter. Lecture Notes in Computer Science, 7551: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34611-8/1.

Gerrard:2012:TPA

[36] Claire Gerrard, John McCall, George M. Coghill, and Christopher Macleod. Temporal patterns in artificial reaction networks. Lecture Notes in Computer Science, 7552:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33269-2_1/.

Karandashev:2012:PHM

[37] Iakov Karandashev, Boris Kryzhanovsky, and Leonid Litinskii. Properties of the Hopfield model with weighted patterns. Lecture Notes in Computer Science, 7552:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33269-2_2/.

Cherif:2012:DOG

[38] Farouk Chérif. Dynamics and oscillations of GHNNs with time-varying delay. Lecture Notes in Computer Science, 7552:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33269-2_3/.

Duran:2012:DFA

[39] Boris Durán, Yulia Sandamirskaya, and Gregor Schöner. A dynamic field architecture for the generation of hierarchically organized sequences. Lecture Notes in Computer Science, 7552:25–32, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33269-2_4/.

Matuszak:2012:STI

[40] Michal Matuszak and Jacek Miekisz. Stochastic techniques in influence diagrams for learning Bayesian network structure. Lecture Notes in Computer Science, 7552:33–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33269-2_5/.

Karandashev:2012:MMM

[41] Iakov Karandashev and Boris Kryzhanovsky. The mix-matrix method in the problem of binary quadratic optimization. Lecture Notes in Computer Science, 7552:41–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33269-2_6/.

Austin:2012:RCA

[42] James Austin, Stephen Hobson, Nathan Burles, and Simon O’Keefe. A rule chaining architecture using a correlation matrix memory. Lecture Notes in Computer Science, 7552:49–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33269-2_7/.

Bacciu:2012:GMK

[43] Davide Bacciu, Alessio Micheli, and Alessandro Sperduti. A generative multiset kernel for structured data. Lecture Notes in Computer Science, 7552:57–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33269-2_8/.

Zdunek:2012:SSU

[44] Rafal Zdunek. Spectral signal unmixing with interior-point nonnegative matrix factorization. Lecture Notes in Computer Science, 7552: 65–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33269-2_9/.

Wang:2012:HOP

[45] Dan Wang, Donghong Ji, and Wei Huang. Hybrid optimized polynomial neural networks with polynomial neurons and fuzzy polynomial neurons. Lecture Notes in Computer Science, 7552:73–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33269-2_10/.

Cho:2012:TTR

[46] KyungHyun Cho, Alexander Ilin, and Tapani Raiko. Tikhonov-type regularization for restricted Boltzmann machines. Lecture Notes in Computer Science, 7552:81–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33269-2_11/.

Wood:2012:MSA

[47] Richard Wood, Ian Bruce, and Peter Mascher. Modeling of spiking analog neural circuits with Hebbian learning, using amorphous semiconductor thin film transistors with silicon oxide nitride semiconductor split gates. Lecture Notes in Computer Science, 7552:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33269-2_12/.

Beuler:2012:RTS

[48] Marcel Beuler, Aubin Tchaptchet, Werner Bonath, Svetlana Postnova, and Hans Albert Braun. Real-time simulations of synchronization in a conductance-based neuronal network with a digital FPGA hardwarecore. Lecture Notes in Computer Science, 7552:97–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33269-2_13/.

Singh:2012:IFE

[49] Anand Singh, Pierre J. Magistretti, Bruno Weber, and Renaud Jolivet. Impact of frequency on the energetic efficiency of action potentials. Lecture Notes in Computer Science, 7552:105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33269-2_14/.

Cheung:2012:LSS

[50] Kit Cheung, Simon R. Schultz, and Wayne Luk. A large-scale spiking neural network accelerator for FPGA systems. Lecture Notes in Computer Science, 7552:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33269-2_15/.

Choudhary:2012:SNC

[51] Swadesh Choudhary, Steven Sloan, Sam Fok, Alexander Neckar, Eric Trautmann, and Peiran Gao. Silicon neurons that compute. Lecture Notes in Computer Science, 7552:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33269-2_16/.

Barrera:2012:CIE

[52] Andres Gaona Barrera and Manuel Moreno Arostegui. A communication infrastructure for emulating large-scale neural networks models. Lecture Notes in Computer Science, 7552:129–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33269-2_17/.

Yusoff:2012:PAL

[53] Nooraini Yusoff, André Grüning, and Scott Notley. Pair-associate learning with modulated spike-time dependent plasticity. Lecture Notes in Computer Science, 7552:137–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33269-2_18/.

Agnes:2012:AMN

[54] Everton J. Agnes, Rubem Erichsen Jr., and Leonardo G. Brunnet. Associative memory in neuronal networks of spiking neurons: Architecture and storage analysis. Lecture Notes in Computer Science, 7552: 145–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33269-2_19/.

Kirikawa:2012:BNF

[55] Shota Kirikawa, Takashi Ogawa, and Toshimichi Saito. Bifurcating neurons with filtered base signals. Lecture Notes in Computer Science, 7552: 153–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33269-2_20/.

Anonymous:2012:FMc

[56] Anonymous. Front matter. Lecture Notes in Computer Science, 7552: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33269-2/1.

Suzumura:2012:CVM

[57] Shinya Suzumura and Ryohei Nakano. Complex-valued multilayer perceptron search utilizing eigen vector descent and reducibility mapping. Lecture Notes in Computer Science, 7553:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33266-1_1/.

Hara:2012:TAF

[58] Kazuyuki Hara, Kentaro Katahira, Kazuo Okanoya, and Masato Okada. Theoretical analysis of function of derivative term in on-line gradient descent learning. Lecture Notes in Computer Science, 7553:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33266-1_2/.

Krurkova:2012:SCN

[59] Věra Kůrková. Some comparisons of networks with radial and kernel units. Lecture Notes in Computer Science, 7553:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_3/.

Ribeiro:2012:MPL

[60] Geraldina Ribeiro, Wouter Duivesteijn, Carlos Soares, and Arno Knobbe. Multilayer perceptron for label ranking. Lecture Notes in Computer Science, 7553:25–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33266-1_4/.

Koprinska:2012:ELF

[61] Irena Koprinska, Mashud Rana, and Vassilios G. Agelidis. Electricity load forecasting: a weekday-based approach. Lecture Notes in Computer Science, 7553:33–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33266-1_5/.

Tokic:2012:AEU

[62] Michel Tokic and Günther Palm. Adaptive exploration using stochastic neurons. Lecture Notes in Computer Science, 7553:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_6/.

Steege:2012:CLT

[63] Frank-Florian Steege and Horst-Michael Groß. Comparison of long-term adaptivity for neural networks. Lecture Notes in Computer Science, 7553: 50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33266-1_7/.

Mamalet:2012:SCF

[64] Franck Mamalet and Christophe Garcia. Simplifying ConvNets for fast learning. Lecture Notes in Computer Science, 7553:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_8/.

deOliveira:2012:MAF

[65] João Fausto Lorenzato de Oliveira and Teresa B. Ludermir. A modified artificial fish swarm algorithm for the optimization of extreme learning machines. Lecture Notes in Computer Science, 7553:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_9/.

Ninomiya:2012:RTF

[66] Hiroshi Ninomiya. Robust training of feedforward neural networks using combined online/batch quasi-Newton techniques. Lecture Notes in Computer Science, 7553:74–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33266-1_10/.

Entner:2012:ECO

[67] Doris Entner and Patrik O. Hoyer. Estimating a causal order among groups of variables in linear models. Lecture Notes in Computer Science, 7553:84–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33266-1_11/.

Brakel:2012:TRB

[68] Philemon Brakel, Sander Dieleman, and Benjamin Schrauwen. Training restricted Boltzmann machines with multi-tempering: Harnessing parallelization. Lecture Notes in Computer Science, 7553:92–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_12/.

Torres:2012:CGA

[69] Luiz C. B. Torres, Cristiano L. Castro, and Antônio P. Braga. A computational geometry approach for Pareto-optimal selection of neural networks. Lecture Notes in Computer Science, 7553:100–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_13/.

Kassahun:2012:LPL

[70] Yohannes Kassahun, Hendrik Wöhrle, Alexander Fabisch, and Marc Tabie. Learning parameters of linear models in compressed parameter space. Lecture Notes in Computer Science, 7553:108–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33266-1_14/.

Nakano:2012:CFF

[71] Daichi Nakano, Shin ichi Maeda, and Shin Ishii. Control of a free-falling cat by policy-based reinforcement learning. Lecture Notes in Computer Science, 7553:116–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33266-1_15/.

Hao:2012:GBM

[72] Tele Hao, Tapani Raiko, Alexander Ilin, and Juha Karhunen. Gated Boltzmann machine in texture modeling. Lecture Notes in Computer Science, 7553:124–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33266-1_16/.

Kromer:2012:NPM

[73] Pavel Krömer, Emilio Corchado, Václav Snášel, Jan Platoš, and Laura García-Hernández. Neural PCA and maximum likelihood Hebbian learning on the GPU. Lecture Notes in Computer Science, 7553:132–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33266-1_17/.

Diez-Fernandez:2012:CEM

[74] Marta Díez-Fernández, Sergio Alvarez Teleña, and Denise Gorse. Construction of emerging markets exchange traded funds using multiobjective particle swarm optimisation. Lecture Notes in Computer Science, 7553: 140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33266-1_18/.

Goncalves:2012:ISC

[75] André R. Gonçalves, Rosana Veroneze, Salomão Madeiro, and Carlos R. B. Azevedo. The influence of supervised clustering for RBFNN centers definition: a comparative study. Lecture Notes in Computer Science, 7553:148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33266-1_19/.

Ghio:2012:NSM

[76] Alessandro Ghio, Davide Anguita, Luca Oneto, Sandro Ridella, and Carlotta Schatten. Nested sequential minimal optimization for support vector machines. Lecture Notes in Computer Science, 7553:156– 163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33266-1_20/.

Anonymous:2012:FMd

[77] Anonymous. Front matter. Lecture Notes in Computer Science, 7553: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33266-1/1.

Barbosa:2012:SCI

[78] Luis Soares Barbosa. Software components as invariant-typed arrows. Lecture Notes in Computer Science, 7554:1–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33182-4_1/.

Aichernig:2012:SKB

[79] Bernhard Klaus Aichernig. The science of Killing bugs in a black box. Lecture Notes in Computer Science, 7554:6–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33182-4_2/.

Sampaio:2012:SCP

[80] Diogo Nunes Sampaio, Elie Gedeon, and Fernando Magno Quintão Pereira. Spill code placement for SIMD machines. Lecture Notes in Computer Science, 7554:12–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33182-4_3/.

Medeiros:2012:LRP

[81] Sérgio Medeiros, Fabio Mascarenhas, and Roberto Ierusalimschy. Left recursion in parsing expression grammars. Lecture Notes in Computer Science, 7554:27–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33182-4_4/.

Campos:2012:SPR

[82] Victor Hugo Sperle Campos and Raphael Ernani Rodrigues. Speed and precision in range analysis. Lecture Notes in Computer Science, 7554: 42–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33182-4_5/.

Alves:2012:PBC

[83] Péricles Rafael Oliveira Alves and Igor Rafael de Assis Costa. Parameter based constant propagation. Lecture Notes in Computer Science, 7554:57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33182-4_6/.

Reis:2012:APE

[84] Leonardo Vieira dos Santos Reis and Roberto da Silva Bigonha. Adaptable parsing expression grammars. Lecture Notes in Computer Science, 7554:72–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33182-4_7/.

Ferreira:2012:OGD

[85] Bruno Morais Ferreira and Fernando Magno Quintão Pereira. Optimizing a geomodeling domain specific language. Lecture Notes in Computer Science, 7554:87–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33182-4_8/.

deBayser:2012:SRT

[86] Maximilien de Bayser and Renato Cerqueira. A system for runtime type introspection in C++. Lecture Notes in Computer Science, 7554: 102–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33182-4_9/.

Cunha:2012:MBP

[87] Jácome Cunha, João Saraiva, and Joost Visser. Model-based programming environments for spreadsheets. Lecture Notes in Computer Science, 7554:117–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33182-4_10/.

Gaia:2012:QAA

[88] Felipe Nunes Gaia and Gabriel Coutinho Sousa Ferreira. A quantitative assessment of aspectual feature modules for evolving software product lines. Lecture Notes in Computer Science, 7554:134–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33182-4_11/.

Viera:2012:AGM

[89] Marcos Viera and Doaitse Swierstra. Attribute grammar macros. Lecture Notes in Computer Science, 7554:150–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33182-4_12/.

Anonymous:2012:BMb

[90] Anonymous. Back matter. Lecture Notes in Computer Science, 7554: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33182-4/1.

Anonymous:2012:FMe

[91] Anonymous. Front matter. Lecture Notes in Computer Science, 7554: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33182-4/1.

Tierney:2012:LSF

[92] Kevin Tierney and Rune Møller Jensen. The liner shipping fleet repositioning problem with cargo flows. Lecture Notes in Computer Science, 7555:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_1/.

Pacino:2012:AMS

[93] Dario Pacino, Alberto Delgado, Rune Møller Jensen, and Tom Bebbington. An accurate model for seaworthy container vessel stowage planning with ballast tanks. Lecture Notes in Computer Science, 7555: 17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_2/.

Schwarze:2012:SAC

[94] Silvia Schwarze, Stefan Voß, Guohua Zhou, and Guoli Zhou. Scientometric analysis of container terminals and ports literature and interaction with publications on distribution networks. Lecture Notes in Computer Science, 7555:33–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_3/.

Nabais:2012:NPC

[95] João Lemos Nabais, Rudy R. Negenborn, and Miguel Ayala Botto. A novel predictive control based framework for optimizing intermodal container terminal operations. Lecture Notes in Computer Science, 7555: 53–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_4/.

Lam:2012:IPD

[96] Jasmine Siu Lee Lam and Tsz Leung Yip. Impact of port disruption on supply chains: a Petri net approach. Lecture Notes in Computer Science, 7555:72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_5/.

Voss:2012:EMO

[97] Stefan Voß. Extended mis-overlay calculation for pre-marshalling containers. Lecture Notes in Computer Science, 7555:86–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33587-7_6/.

Litvinchev:2012:STS

[98] Igor Litvinchev and Edith Lucero Ozuna Espinosa. Solving the two-stage capacitated facility location problem by the Lagrangian heuristic. Lecture Notes in Computer Science, 7555:92–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33587-7_7/.

Daduna:2012:ARF

[99] Joachim R. Daduna. Applying radio frequency identification technology in retail trade from a logistics point of view — an overview over opportunities and limitations. Lecture Notes in Computer Science, 7555: 104–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_8/.

Reinhardt:2012:SVR

[100] Line Blander Reinhardt, Simon Spoorendonk, and David Pisinger. Solving vehicle routing with full container load and time Windows. Lecture Notes in Computer Science, 7555:120–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33587-7_9/.

Iris:2012:MIS

[101] Cagatay Iris and Mehmet Mutlu Yenisey. Multi-item simultaneous lot sizing and storage allocation with production and warehouse capacities. Lecture Notes in Computer Science, 7555:129–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33587-7_10/.

Lange:2012:PML

[102] Kerstin Lange, André Rinne, and Hans-Dietrich Haasis. Planning maritime logistics concepts for offshore wind farms: a newly developed decision support system. Lecture Notes in Computer Science, 7555: 142–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_11/.

Xu:2012:ACB

[103] Feng Xu, Yuanbin Song, and Hao Hu. Application of cycle-based simulation to estimate loss of logistics productivity on construction sites. Lecture Notes in Computer Science, 7555:159–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33587-7_12/.

Bubalo:2012:BEA

[104] Branko Bubalo. Benchmarking European airports based on a profitability envelope. Lecture Notes in Computer Science, 7555:171–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33587-7_13/.

Cao:2012:SVR

[105] Buyang Cao. Solving vehicle routing problems using an enhanced Clarke– Wright algorithm: a case study. Lecture Notes in Computer Science, 7555: 190–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33587-7_14/.

Sze:2012:IHM

[106] San-Nah Sze, Ada Ng Suk-Fong, and Kang-Leng Chiew. An insertion heuristic manpower scheduling for in-flight catering service application. Lecture Notes in Computer Science, 7555:206–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33587-7_15/.

Anonymous:2012:BMc

[107] Anonymous. Back matter. Lecture Notes in Computer Science, 7555: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33587-7/1.

Anonymous:2012:FMf

[108] Anonymous. Front matter. Lecture Notes in Computer Science, 7555: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33587-7/1.

deWolf:2012:PSD

[109] Peter-Paul de Wolf and Anco Hundepool. p% should dominate. Lecture Notes in Computer Science, 7556:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33627-0_1/.

Castro:2012:CEO

[110] Jordi Castro. A computational evaluation of optimization solvers for CTA. Lecture Notes in Computer Science, 7556:11–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33627-0_2/.

Giessing:2012:FRB

[111] Sarah Giessing. Flexible rounding based on consistent post-tabular stochastic noise. Lecture Notes in Computer Science, 7556:22–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33627-0_3/.

Castro:2012:CDC

[112] Jordi Castro. Comparing L_{1} and L_{2} distances for CTA. Lecture Notes in Computer Science, 7556:35–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33627-0_4/.

Trottini:2012:IMB

[113] Mario Trottini, Krish Muralidhar, and Rathindra Sarathy. An investigation of model-based microdata masking for magnitude tabular data release. Lecture Notes in Computer Science, 7556:47–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33627-0_5/.

Kowarik:2012:TIC

[114] Alexander Kowarik, Matthias Templ, Bernhard Meindl, Francois Fonteneau, and Bernd Prantner. Testing of IHSN C++ code and inclusion of new methods into sdcMicro. Lecture Notes in Computer Science, 7556: 63–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33627-0_6/.

Mares:2012:CBC

[115] Jordi Marés and Vicenç Torra. Clustering-based categorical data protection. Lecture Notes in Computer Science, 7556:78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33627-0_7/.

Domingo-Ferrer:2012:AMT

[116] Josep Domingo-Ferrer, Krish Muralidhar, and Guillem Rufian-Torrell. Anonymization methods for taxonomic microdata. Lecture Notes in Computer Science, 7556:90–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33627-0_8/.

Oganian:2012:HMM

[117] Anna Oganian and Josep Domingo-Ferrer. Hybrid microdata via modelbased clustering. Lecture Notes in Computer Science, 7556:103–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33627-0_9/.

Woo:2012:LRV

[118] Yong Ming Jeffrey Woo and Aleksandra B. Slavković. Logistic regression with variables subject to post randomization method. Lecture Notes in Computer Science, 7556:116–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33627-0_10/.

Hall:2012:VSI

[119] Rob Hall and Stephen Fienberg. Valid statistical inference on automatically matched files. Lecture Notes in Computer Science, 7556: 131–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33627-0_11/.

DePersio:2012:CSA

[120] Michael DePersio, Marlow Lemons, Kaleli A. Ramanayake, Julie Tsay, and Laura Zayatz. n-cycle swapping for the American Community Survey. Lecture Notes in Computer Science, 7556:143–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33627-0_12/.

Dorner:2012:GUT

[121] Matthias Dorner, Jörg Drechsler, and Peter Jacobebbinghaus. Generating useful test data for complex linked employer–employee datasets. Lecture Notes in Computer Science, 7556:165–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33627-0_13/.

Cleveland:2012:WEP

[122] Lara Cleveland, Robert McCaa, Steven Ruggles, and Matthew Sobek. When excessive perturbation goes wrong and why IPUMS-international relies instead on sampling, suppression, swapping, and other minimally harmful methods to protect privacy of census microdata. Lecture Notes in Computer Science, 7556:179–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33627-0_14/.

Franconi:2012:ACE

[123] Luisa Franconi and Daniela Ichim. Achieving comparability of earnings. Lecture Notes in Computer Science, 7556:188–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33627-0_15/.

Foschi:2012:DMR

[124] Flavio Foschi, Maria Cristina Casciano, Luisa Franconi, and Daniela Ichim. Designing multiple releases from the small and medium enterprises survey. Lecture Notes in Computer Science, 7556:200–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33627-0_16/.

Abowd:2012:PSA

[125] John M. Abowd, Lars Vilhuber, and William Block. A proposed solution to the archiving and curation of confidential scientific inputs. Lecture Notes in Computer Science, 7556:216–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33627-0_17/.

OKeefe:2012:CMM

[126] Christine M. O’Keefe. Confidentialising maps of mixed point and diffuse spatial data. Lecture Notes in Computer Science, 7556:226–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33627-0_18/.

Zanger:2012:SRO

[127] Daniel Z. Zanger. Security of random output perturbation for statistical databases. Lecture Notes in Computer Science, 7556:241–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33627-0_19/.

Charest:2012:EES

[128] Anne-Sophie Charest. Empirical evaluation of statistical inference from differentially-private contingency tables. Lecture Notes in Computer Science, 7556:257–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33627-0_20/.

Anonymous:2012:FMg

[129] Anonymous. Front matter. Lecture Notes in Computer Science, 7556: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33627-0/1.

HaCohen-Kerner:2012:EFD

[130] Yaakov HaCohen-Kerner and Aharon Tayeb. Experiments with filtered detection of similar academic papers. Lecture Notes in Computer Science, 7557:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33185-5_1/.

Piasecki:2012:CBS

[131] Maciej Piasecki, Radosław Ramocki, and Paweł Minda. Corpus-based semantic filtering in discovering derivational relations. Lecture Notes in Computer Science, 7557:14–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_2/.

Piasecki:2012:LAA

[132] Maciej Piasecki, Roman Kurc, Radosław Ramocki, and Bartosz Broda. Lexical activation area attachment algorithm for Wordnet expansion. Lecture Notes in Computer Science, 7557:23–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33185-5_3/.

Klyk:2012:MTM

[133] Łukasz Kłyk, Paweł B. Myszkowski, Bartosz Broda, Maciej Piasecki, and David Urbansky. Metaheuristics for tuning model parameters in two natural language processing applications. Lecture Notes in Computer Science, 7557:32–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33185-5_4/.

Zizka:2012:CVL

[134] Jan Žižka, Karel Burda, and František Dařena. Clustering a very large number of textual unstructured customers’ reviews in English. Lecture Notes in Computer Science, 7557:38–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33185-5_5/.

Boytcheva:2012:WTE

[135] Svetla Boytcheva and Galia Angelova. A workbench for temporal event information extraction from patient records. Lecture Notes in Computer Science, 7557:48–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33185-5_6/.

Lupan:2012:AES

[136] Diana Lupan, Mihai Dascălu, Ştefan Trăuşan-Matu, and Philippe Dessus. Analyzing emotional states induced by news articles with latent semantic analysis. Lecture Notes in Computer Science, 7557:59–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33185-5_7/.

Ali:2012:ECH

[137] Mostafa Z. Ali and Robert G. Reynolds. The emergence of cultural hierarchical social networks in complex environments. Lecture Notes in Computer Science, 7557:69–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_8/.

Tandukar:2012:ERS

[138] Udeep Tandukar and Julita Vassileva. Ensuring relevant and serendipitous information flow in decentralized online social network. Lecture Notes in Computer Science, 7557:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_9/.

Arendarenko:2012:OBI

[139] Ernest Arendarenko and Tuomo Kakkonen. Ontology-based information and event extraction for business intelligence. Lecture Notes in Computer Science, 7557:89–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33185-5_10/.

Kutz:2012:MHS

[140] Oliver Kutz, Janna Hastings, and Till Mossakowski. Modelling highly symmetrical molecules: Linking ontologies and graphs. Lecture Notes in Computer Science, 7557:103–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_11/.

Beldjoudi:2012:PIT

[141] Samia Beldjoudi, Hassina Seridi-Bouchelaghem, and Catherine Faron-Zucker. Personalizing and improving tag-based search in folksonomies. Lecture Notes in Computer Science, 7557:112–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33185-5_12/.

LeDorze:2012:VSC

[142] Aymeric Le Dorze, Lionel Chauvin, Laurent Garcia, David Genest, and Stéphane Loiseau. Views and synthesis of cognitive maps. Lecture Notes in Computer Science, 7557:119–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_13/.

Gruca:2012:ICS

[143] Aleksandra Gruca and Marek Sikora. Identification of the compound subjective rule interestingness measure for rule-based functional description of genes. Lecture Notes in Computer Science, 7557:125–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33185-5_14/.

Leonetti:2012:AGL

[144] Matteo Leonetti, Luca Iocchi, and Fabio Patrizi. Automatic generation and learning of finite-state controllers. Lecture Notes in Computer Science, 7557:135–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33185-5_15/.

Damova:2012:FDS

[145] Mariana Damova, Kiril Simov, Zdravko Tashev, and Atanas Kiryakov. FactForge: Data service or the diversity of inferred knowledge over LOD. Lecture Notes in Computer Science, 7557:145–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33185-5_16/.

Amaneddine:2012:PCG

[146] Nouhad Amaneddine and Jean-François Condotta. From path-consistency to global consistency in temporal qualitative constraint networks. Lecture Notes in Computer Science, 7557:152–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33185-5_17/.

Sikora:2012:RQM

[147] Marek Sikora, Adam Skowron, and Łukasz Wróbel. Rule quality measurebased induction of unordered sets of regression rules. Lecture Notes in Computer Science, 7557:162–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_18/.

Kabzinski:2012:SCF

[148] Jacek Kabziński. Swarm capability of finding eigenvalues. Lecture Notes in Computer Science, 7557:172–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33185-5_19/.

Nadi:2012:SUP

[149] Farhad Nadi and Ahamad Tajudin Khader. A study on the utility of parametric uniform crossover for adaptation of crossover operator. Lecture Notes in Computer Science, 7557:178–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33185-5_20/.

Anonymous:2012:FMh

[150] Anonymous. Front matter. Lecture Notes in Computer Science, 7557: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33185-5/1.

Dettori:2012:AES

[151] Giuliana Dettori and Simone Torsani. An approach to exploit social bookmarking to improve formal language learning. Lecture Notes in Computer Science, 7558:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33642-3_1/.

Hillen:2012:PSC

[152] Stefanie Andrea Hillen and Tero Päivärinta. Perceived support in Ecollaborative learning: An exploratory study which make use of synchronous and asynchronous online-teaching approaches. Lecture Notes in Computer Science, 7558:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33642-3_2/.

Kubincova:2012:HGA

[153] Zuzana Kubincová and Martin Homola. How to get around with Wikis in teaching. Lecture Notes in Computer Science, 7558:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33642-3_3/.

Putnik:2012:WUT

[154] Zoran Putnik, Mirjana Ivanović, Zoran Budimac, and Ladislav Samuelis. Wiki — a useful tool to fight classroom cheating? Lecture Notes in Computer Science, 7558:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33642-3_4/.

Popescu:2012:PBL

[155] Elvira Popescu. Project-based learning with eMUSE. Lecture Notes in Computer Science, 7558:41–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33642-3_5/.

Nunes:2012:ATE

[156] Bernardo Pereira Nunes, Ricardo Kawase, and Stefan Dietze. Annotation tool for enhancing E-learning courses. Lecture Notes in Computer Science, 7558:51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33642-3_6/.

Kutay:2012:TCS

[157] Cat Kutay, Deirdre Howard-Wagner, Lynette Riley, and Janet Mooney. Teaching culture as social constructivism. Lecture Notes in Computer Science, 7558:61–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33642-3_7/.

Bogdanov:2012:LPS

[158] Evgeny Bogdanov, Carsten Ullrich, Erik Isaksson, Matthias Palmer, and Denis Gillet. From LMS to PLE: a step forward through OpenSocial apps in Moodle. Lecture Notes in Computer Science, 7558:69–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33642-3_8/.

Nussbaumer:2012:MRC

[159] Alexander Nussbaumer, Marcel Berthold, Daniel Dahrendorf, and Hans-Christian Schmitz. A mashup recommender for creating personal learning environments. Lecture Notes in Computer Science, 7558:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33642-3_9/.

Zhou:2012:FRS

[160] Lei Zhou, Sandy El Helou, Laurent Moccozet, Laurent Opprecht, and Omar Benkacem. A federated recommender system for online learning environments. Lecture Notes in Computer Science, 7558:89–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33642-3_10/.

Lorenz:2012:TSI

[161] Birgy Lorenz, Kätlin Kalde, and Kaido Kikkas. Trust and security issues in cloud-based learning and management. Lecture Notes in Computer Science, 7558:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33642-3_11/.

Levacher:2012:STL

[162] Killian Levacher, Seamus Lawless, and Vincent Wade. Slicepedia: Towards Long tail resource production through open corpus reuse. Lecture Notes in Computer Science, 7558:109–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33642-3_12/.

Yang:2012:LPC

[163] Fan Yang, Frederick W. B. Li, and Rynson W. H. Lau. Learning path construction based on association link network. Lecture Notes in Computer Science, 7558:120–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33642-3_13/.

Simko:2012:DHR

[164] Marián Šimko and Mária Bieliková. Discovering hierarchical relationships in educational content. Lecture Notes in Computer Science, 7558: 132–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33642-3_14/.

Ciuciu:2012:EMC

[165] Ioana Ciuciu and Yan Tang Demey. An evaluation methodology for C-FOAM applied to Web-based learning. Lecture Notes in Computer Science, 7558:142–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33642-3_15/.

Kouneli:2012:MKD

[166] Aggeliki Kouneli, Georgia Solomou, Christos Pierrakeas, and Achilles Kameas. Modeling the knowledge domain of the Java programming language as an ontology. Lecture Notes in Computer Science, 7558: 152–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33642-3_16/.

Torrente:2012:TUG

[167] Javier Torrente, Ángel del Blanco, Ángel Serrano-Laguna, and José Ángel Vallejo-Pinto. Towards universal game development in education. Lecture Notes in Computer Science, 7558:160–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33642-3_17/.

Lavoue:2012:TSL

[168] Élise Lavoué. Towards social learning games. Lecture Notes in Computer Science, 7558:170–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33642-3_18/.

Cehan:2012:CGE

[169] Sabina-Nadina Cehan and Dana-Anca Cehan. Computer games and English language learning. Lecture Notes in Computer Science, 7558: 180–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33642-3_19/.

Staikopoulos:2012:AFC

[170] Athanasios Staikopoulos, Ian O’Keeffe, Rachael Rafter, Eddie Walsh, and Bilal Yousuf. AMASE: a framework for composing adaptive and personalised learning activities on the Web. Lecture Notes in Computer Science, 7558:190–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33642-3_20/.

Anonymous:2012:FMi

[171] Anonymous. Front matter. Lecture Notes in Computer Science, 7558: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33642-3/1.

Salah:2012:HBUa

[172] Albert Ali Salah, Javier Ruiz del Solar, Çetin Meriçli, and Pierre-Yves Oudeyer. Human behavior understanding for robotics. Lecture Notes in Computer Science, 7559:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34014-7_1/.

Celiktutan:2012:RTE

[173] Oya Çeliktutan, Christian Wolf, Bülent Sankur, and Eric Lombardi. Realtime exact graph matching with application in human action recognition. Lecture Notes in Computer Science, 7559:17–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34014-7_2/.

Chaaraoui:2012:EAM

[174] Alexandros Andre Chaaraoui, Pau Climent-Pérez, and Francisco Flórez-Revuelta. An efficient approach for multi-view human action recognition based on bag-of-key-poses. Lecture Notes in Computer Science, 7559: 29–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34014-7_3/.

Hu:2012:BFC

[175] Ninghang Hu, Gwenn Englebienne, and Ben J. A. Kröse. Bayesian fusion of ceiling mounted camera and laser range finder on a mobile robot for people detection and localization. Lecture Notes in Computer Science, 7559:41–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34014-7_4/.

Lim:2012:USD

[176] Angelica Lim and Hiroshi G. Okuno. Using speech data to recognize emotion in human gait. Lecture Notes in Computer Science, 7559: 52–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34014-7_5/.

Samadani:2012:GDP

[177] Ali-Akbar Samadani, Rob Gorbet, and Dana Kulić. Gender differences in the perception of affective movements. Lecture Notes in Computer Science, 7559:65–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34014-7_6/.

Vincze:2012:VDA

[178] Laura Vincze, Isabella Poggi, and Francesca D’Errico. Vagueness and dreams: Analysis of body signals in vague dream telling. Lecture Notes in Computer Science, 7559:77–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34014-7_7/.

Mancini:2012:CEB

[179] Maurizio Mancini, Giovanna Varni, Donald Glowinski, and Gualtiero Volpe. Computing and evaluating the body laughter index. Lecture Notes in Computer Science, 7559:90–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34014-7_8/.

Sheikhi:2012:RVF

[180] Samira Sheikhi and Jean-Marc Odobez. Recognizing the visual focus of attention for human robot interaction. Lecture Notes in Computer Science, 7559:99–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34014-7_9/.

Karreman:2012:CAH

[181] Daphne E. Karreman, Elisabeth M. A. G. van Dijk, and Vanessa Evers. Contextual analysis of human non-verbal guide behaviors to inform the development of FROG, the fun robotic outdoor guide. Lecture Notes in Computer Science, 7559:113–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34014-7_10/.

Fischer:2012:GAD

[182] Kerstin Fischer and Joe Saunders. Getting acquainted with a developing robot. Lecture Notes in Computer Science, 7559:125–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34014-7_11/.

Mangin:2012:LCS

[183] Olivier Mangin and Pierre-Yves Oudeyer. Learning the combinatorial structure of demonstrated behaviors with inverse feedback control. Lecture Notes in Computer Science, 7559:134–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34014-7_12/.

Schillaci:2012:ISB

[184] Guido Schillaci, Bruno Lara, and Verena V. Hafner. Internal simulations for behaviour selection and recognition. Lecture Notes in Computer Science, 7559:148–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34014-7_13/.

Michelet:2012:AIA

[185] Stéphane Michelet, Koby Karp, Emilie Delaherche, Catherine Achard, and Mohamed Chetouani. Automatic imitation assessment in interaction. Lecture Notes in Computer Science, 7559:161–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34014-7_14/.

Anonymous:2012:BMd

[186] Anonymous. Back matter. Lecture Notes in Computer Science, 7559: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-34014-7/1.

Anonymous:2012:FMj

[187] Anonymous. Front matter. Lecture Notes in Computer Science, 7559: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34014-7/1.

Dang:2012:TSL

[188] Han-Hing Dang and Bernhard Möller. Transitive separation logic. Lecture Notes in Computer Science, 7560:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33314-9_1/.

Guttmann:2012:ULS

[189] Walter Guttmann. Unifying lazy and strict computations. Lecture Notes in Computer Science, 7560:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33314-9_2/.

Hofner:2012:FCA

[190] Peter Höfner, Bernhard Möller, and Andreas Zelend. Foundations of coloring algebra with consequences for feature-oriented programming. Lecture Notes in Computer Science, 7560:33–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33314-9_3/.

Dongol:2012:TAR

[191] Brijesh Dongol, Ian J. Hayes, Larissa Meinicke, and Kim Solin. Towards an algebra for real-time programs. Lecture Notes in Computer Science, 7560:50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33314-9_4/.

Armstrong:2012:ARH

[192] Alasdair Armstrong and Georg Struth. Automated reasoning in higherorder regular algebra. Lecture Notes in Computer Science, 7560:66– 81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33314-9_5/.

Kahl:2012:TCI

[193] Wolfram Kahl. Towards certifiable implementation of graph transformation via relation categories. Lecture Notes in Computer Science, 7560:

82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33314-9_6/.

Moreira:2012:DRE

[194] Nelma Moreira, David Pereira, and Simão Melo de Sousa. Deciding regular expressions (in-)equivalence in coq. Lecture Notes in Computer Science, 7560:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33314-9_7/.

Berghammer:2012:SRB

[195] Rudolf Berghammer and Sebastian Fischer. Simple rectangle-based functional programs for computing reflexive-transitive closures. Lecture Notes in Computer Science, 7560:114–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33314-9_8/.

Danilenko:2012:URD

[196] Nikita Danilenko. Using relations to develop a Haskell program for computing maximum bipartite matchings. Lecture Notes in Computer Science, 7560:130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33314-9_9/.

Macedo:2012:RES

[197] Nuno Macedo, Hugo Pacheco, and Alcino Cunha. Relations as executable specifications: Taming partiality and non-determinism using invariants. Lecture Notes in Computer Science, 7560:146–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33314-9_10/.

Kozen:2012:LHC

[198] Dexter Kozen and Alexandra Silva. Left-handed completeness. Lecture Notes in Computer Science, 7560:162–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33314-9_11/.

Laurence:2012:COR

[199] Michael R. Laurence and Georg Struth. On completeness of omegaregular algebras. Lecture Notes in Computer Science, 7560:179–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33314-9_12/.

Jipsen:2012:CAC

[200] Peter Jipsen. Categories of algebraic contexts equivalent to idempotent semirings and domain semirings. Lecture Notes in Computer Science, 7560:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33314-9_13/.

Nishizawa:2012:RRT

[201] Koki Nishizawa and Hitoshi Furusawa. Relational representation theorem for powerset quantales. Lecture Notes in Computer Science, 7560: 207–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33314-9_14/.

Furusawa:2012:PAD

[202] Hitoshi Furusawa and Yasuo Kawahara. Point axioms in Dedekind categories. Lecture Notes in Computer Science, 7560:219–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33314-9_15/.

Gluck:2012:TOD

[203] Roland Glück. Two observations in dioid based model refinement. Lecture Notes in Computer Science, 7560:235–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33314-9_16/.

Atampore:2012:RAM

[204] Francis Atampore and Michael Winter. Relation algebras, matrices, and multi-valued decision diagrams. Lecture Notes in Computer Science, 7560: 248–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33314-9_17/.

Ismail:2012:IPR

[205] Sahar Ismail and Ali Jaoua. Incremental pseudo rectangular organization of information relative to a domain. Lecture Notes in Computer Science, 7560:264–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33314-9_18/.

Schmidt:2012:RCS

[206] Gunther Schmidt. Relational concepts in social choice. Lecture Notes in Computer Science, 7560:278–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33314-9_19/.

Moller:2012:ALC

[207] Bernhard Möller and Patrick Roocks. An algebra of layered complex preferences. Lecture Notes in Computer Science, 7560:294–309, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33314-9_20/.

Anonymous:2012:FMk

[208] Anonymous. Front matter. Lecture Notes in Computer Science, 7560: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33314-9/1.

Lopes:2012:GBD

[209] Antónia Lopes and José Luiz Fiadeiro. A graph-based design framework for services. Lecture Notes in Computer Science, 7562:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33654-6_1/.

Ruscio:2012:ETH

[210] Davide Di Ruscio, Ludovico Iovino, and Alfonso Pierantonio. Evolutionary togetherness: How to manage coupled evolution in metamodeling ecosystems. Lecture Notes in Computer Science, 7562:20–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33654-6_2/.

Drechsler:2012:CDD

[211] Rolf Drechsler, Melanie Diepenbeck, Daniel Große, Ulrich Kühne, and Hoang M. Le. Completeness-driven development. Lecture Notes in Computer Science, 7562:38–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33654-6_3/.

Gadducci:2012:EUA

[212] Fabio Gadducci, Alberto Lluch Lafuente, and Andrea Vandin. Exploiting over- and under-approximations for infinite-state counterpart models. Lecture Notes in Computer Science, 7562:51–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-33654-6_4/.

Rensink:2012:PBG

[213] Arend Rensink and Eduardo Zambon. Pattern-based graph abstraction. Lecture Notes in Computer Science, 7562:66–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33654-6_5/.

Konig:2012:WSG

[214] Barbara König and Jan Stückrath. Well-structured graph transformation systems with negative application conditions. Lecture Notes in Computer Science, 7562:81–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33654-6_6/.

Heckel:2012:PCS

[215] Reiko Heckel, Hartmut Ehrig, Ulrike Golas, and Frank Hermann. Parallelism and concurrency of stochastic graph transformations. Lecture Notes in Computer Science, 7562:96–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33654-6_7/.

Lowe:2012:RGR

[216] Michael Löwe. Refined graph rewriting in span-categories. Lecture Notes in Computer Science, 7562:111–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33654-6_8/.

Orejas:2012:BCA

[217] Fernando Orejas, Artur Boronat, and Nikos Mylonakis. Borrowed contexts for attributed graphs. Lecture Notes in Computer Science, 7562: 126–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33654-6_9/.

Golas:2012:TBG

[218] Ulrike Golas, Leen Lambers, Hartmut Ehrig, and Holger Giese. Toward bridging the gap between formal foundations and current practice for triple graph grammars. Lecture Notes in Computer Science, 7562: 141–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33654-6_10/.

Duval:2012:GTF

[219] Dominique Duval, Rachid Echahed, and Frédéric Prost. Graph transformation with focus on incident edges. Lecture Notes in Computer Science, 7562:156–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33654-6_11/.

Aoto:2012:RTR

[220] Takahito Aoto and Jeroen Ketema. Rational term rewriting revisited: Decidability and confluence. Lecture Notes in Computer Science, 7562: 172–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33654-6_12/.

Golas:2012:GAC

[221] Ulrike Golas. A general attribution concept for models in ⇕-adhesive transformation systems. Lecture Notes in Computer Science, 7562:187– 202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33654-6_13/.

Heckel:2012:DTO

[222] Reiko Heckel. DPO transformation with open maps. Lecture Notes in Computer Science, 7562:203–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33654-6_14/.

Habel:2012:ATS

[223] Annegret Habel and Detlef Plump. M, N -adhesive transformation systems. Lecture Notes in Computer Science, 7562:218–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33654-6_15/.

Ghamarian:2012:GCG

[224] Amir Hossein Ghamarian and Arend Rensink. Generalised compositionality in graph transformation. Lecture Notes in Computer Science, 7562: 234–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33654-6_16/.

Giese:2012:TAV

[225] Holger Giese and Leen Lambers. Towards automatic verification of behavior preservation for model transformation via invariant checking. Lecture Notes in Computer Science, 7562:249–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33654-6_17/.

Blume:2012:ESI

[226] Christoph Blume, H. J. Sander Bruggink, Dominik Engelke, and Barbara König. Efficient symbolic implementation of graph automata with applications to invariant checking. Lecture Notes in Computer Science, 7562: 264–278, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33654-6_18/.

Khan:2012:TAV

[227] Tamim Ahmed Khan, Olga Runge, and Reiko Heckel. Testing against visual contracts: Model-based coverage. Lecture Notes in Computer Science, 7562:279–293, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33654-6_19/.

Serbanuta:2012:TCS

[228] Traian Florin Şerbănuţă and Grigore Roşu. A truly concurrent semantics for the K framework based on graph transformations. Lecture Notes in Computer Science, 7562:294–310, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33654-6_20/.

Anonymous:2012:FMl

[229] Anonymous. Front matter. Lecture Notes in Computer Science, 7562: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33654-6/1.

Anonymous:2012:FMm

[230] Anonymous. Front matter. Lecture Notes in Computer Science, 7563: 1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33263-0/1/1.

Noss:2012:NLN

[231] Richard Noss. 21^{st} century learning for 21^{st} century skills: What does it mean, and how do we do it? Lecture Notes in Computer Science, 7563:3–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_1/.

Anonymous:2012:FMn

[232] Anonymous. Front matter. Lecture Notes in Computer Science, 7563: 7, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33263-0/2/1.

Anjorin:2012:ESI

[233] Mojisola Anjorin, Thomas Rodenhausen, Renato Domínguez García, and Christoph Rensing. Exploiting semantic information for graph-based recommendations of learning resources. Lecture Notes in Computer Science, 7563:9–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_2/.

Berthold:2012:IEM

[234] Marcel Berthold, Adam Moore, Christina M. Steiner, Conor Gaffney, and Declan Dagger. An initial evaluation of metacognitive scaffolding for experiential training simulators. Lecture Notes in Computer Science, 7563:23–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_3/.

Bonnard:2012:PIL

[235] Quentin Bonnard, Himanshu Verma, Frédéric Kaplan, and Pierre Dillenbourg. Paper interfaces for learning geometry. Lecture Notes in Computer Science, 7563:37–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33263-0_4/.

Derntl:2012:ETP

[236] Michael Derntl and Ralf Klamma. The European TEL projects community from a social network analysis perspective. Lecture Notes in Computer Science, 7563:51–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33263-0_5/.

Do-Lenh:2012:TDE

[237] Son Do-Lenh, Patrick Jermann, Amanda Legge, Guillaume Zufferey, and Pierre Dillenbourg. TinkerLamp 2.0: Designing and evaluating orchestration technologies for the classroom. Lecture Notes in Computer Science, 7563:65–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_6/.

Ferrari:2012:UDC

[238] Anusca Ferrari, Yves Punie, and Christine Redecker. Understanding digital competence in the 21st century: An analysis of current frameworks. Lecture Notes in Computer Science, 7563:79–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33263-0_7/.

Gegenfurtner:2012:HCM

[239] Andreas Gegenfurtner, Koen Veermans, and Marja Vauras. How CSCL moderates the influence of self-efficacy on students’ transfer of learning. Lecture Notes in Computer Science, 7563:93–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33263-0_8/.

Gehlen-Baum:2012:NFH

[240] Vera Gehlen-Baum and Armin Weinberger. Notebook or facebook? How students actually use mobile devices in large lectures. Lecture Notes in Computer Science, 7563:103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33263-0_9/.

Rojas:2012:EOL

[241] Israel Gutiérrez Rojas, Raquel M. Crespo García, and Carlos Delgado Kloos. Enhancing orchestration of lab sessions by means of awareness mechanisms. Lecture Notes in Computer Science, 7563:113–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33263-0_10/.

Hadersberger:2012:DAB

[242] Julia Hadersberger, Alexander Pohl, and François Bry. Discerning actuality in backstage. Lecture Notes in Computer Science, 7563:126– 139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_11/.

Hauff:2012:TRM

[243] Claudia Hauff, Marcel Berthold, Geert-Jan Houben, Christina M. Steiner, and Dietrich Albert. Tweets reveal more than you know: a learning style analysis on Twitter. Lecture Notes in Computer Science, 7563: 140–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_12/.

Hsiao:2012:MSV

[244] I.-Han Hsiao and Peter Brusilovsky. Motivational social visualizations for personalized E-learning. Lecture Notes in Computer Science, 7563: 153–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_13/.

Hussaan:2012:GAL

[245] Aarij Mahmood Hussaan and Karim Sehaba. Generator of adaptive learning scenarios: Design and evaluation in the project CLES. Lecture Notes in Computer Science, 7563:166–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33263-0_14/.

Kaschig:2012:TOA

[246] Andreas Kaschig, Ronald Maier, Alexander Sandow, Alan Brown, and Tobias Ley. Technological and organizational arrangements sparking effects on individual, community and organizational learning. Lecture Notes in Computer Science, 7563:180–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33263-0_15/.

Law:2012:SRE

[247] Effie Lai-Chong Law, Arunangsu Chatterjee, Dominik Renzel, and Ralf Klamma. The social requirements engineering (SRE) approach to developing a large-scale personal learning environment infrastructure. Lecture Notes in Computer Science, 7563:194–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33263-0_16/.

Marne:2012:SFS

[248] Bertrand Marne, John Wisdom, Benjamin Huynh-Kim-Bang, and Jean-Marc Labat. The six facets of serious game design: a methodology enhanced by our design pattern library. Lecture Notes in Computer Science, 7563:208–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33263-0_17/.

McLaren:2012:EHE

[249] Bruce M. McLaren, Deanne Adams, Kelley Durkin, George Goguadze, and Richard E. Mayer. To err is human, to explain and correct is divine: a study of interactive erroneous examples with middle school math students. Lecture Notes in Computer Science, 7563:222–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33263-0_18/.

Mehm:2012:ATA

[250] Florian Mehm, Johannes Konert, Stefan Göbel, and Ralf Steinmetz. An authoring tool for adaptive digital educational games. Lecture Notes in Computer Science, 7563:236–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33263-0_19/.

Michel:2012:DRP

[251] Christine Michel, Elise Lavoué, and Laurent Pietrac. A dashboard to regulate project-based learning. Lecture Notes in Computer Science, 7563: 250–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33263-0_20/.

Anonymous:2012:FMo

[252] Anonymous. Front matter. Lecture Notes in Computer Science, 7563: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33263-0/1.

Bodei:2012:FSU

[253] Chiara Bodei, Pierpaolo Degano, Gian-Luigi Ferrari, and Letterio Galletta. Formalising security in ubiquitous and cloud scenarios. Lecture Notes in Computer Science, 7564:1–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33260-9_1/.

Cho:2012:DSC

[254] Young Im Cho. Designing smart cities: Security issues. Lecture Notes in Computer Science, 7564:30–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33260-9_2/.

Hyla:2012:CBE

[255] Tomasz Hyla and Jerzy Pejaś. Certificate-based encryption scheme with general access structure. Lecture Notes in Computer Science, 7564: 41–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33260-9_3/.

Homsirikamol:2012:SME

[256] Ekawat Homsirikamol, Paweł Morawiecki, Marcin Rogawski, and Marian Srebrny. Security margin evaluation of SHA-3 contest finalists through SAT-based attacks. Lecture Notes in Computer Science, 7564: 56–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33260-9_4/.

Um-e-Ghazia:2012:UCM

[257] Um e Ghazia, Rahat Masood, Muhammad Awais Shibli, and Muhammad Bilal. Usage control model specification in XACML policy language. Lecture Notes in Computer Science, 7564:68–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33260-9_5/.

Deb:2012:TTB

[258] Novarun Deb and Nabendu Chaki. TIDS: Trust-based intrusion detection system for wireless ad-hoc networks. Lecture Notes in Computer Science, 7564:80–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33260-9_6/.

Gajdos:2012:IDC

[259] Petr Gajdoš and Pavel Moravec. Intruder data classification using GM-SOM. Lecture Notes in Computer Science, 7564:92–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33260-9_7/.

Drazdilova:2012:MIS

[260] Pavla Dráždilová, Alisa Babskova, Jan Martinovič, Kateřina Slaninová, and Štěpán Minks. Method for identification of suitable persons in collaborators’ networks. Lecture Notes in Computer Science, 7564:101– 110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33260-9_8/.

Dasgupta:2012:GBF

[261] Subhasis Dasgupta and Aditya Bagchi. A graph-based formalism for controlling access to a digital library ontology. Lecture Notes in Computer Science, 7564:111–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33260-9_9/.

Poniszewska-Maranda:2012:RAA

[262] Aneta Poniszewska-Maranda. Role approach in access control development with the usage control concept. Lecture Notes in Computer Science, 7564:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33260-9_10/.

Misztal:2012:NAR

[263] Krzysztof Misztal, Jacek Tabor, and Khalid Saeed. A new algorithm for rotation detection in Iris pattern recognition. Lecture Notes in Computer Science, 7564:135–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33260-9_11/.

Kozera:2012:ORL

[264] Ryszard Kozera and Jacek Tchórzewski. Outlier removal in 2D leap frog algorithm. Lecture Notes in Computer Science, 7564:146–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33260-9_12/.

Doroz:2012:DSR

[265] Rafal Doroz and Krzysztof Wrobel. Dynamic signature recognition based on modified Windows technique. Lecture Notes in Computer Science, 7564:158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33260-9_13/.

Albarelli:2012:RNR

[266] Andrea Albarelli, Filippo Bergamasco, and Andrea Torsello. Rigid and non-rigid shape matching for mechanical components retrieval. Lecture Notes in Computer Science, 7564:168–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33260-9_14/.

Cyganek:2012:EEE

[267] Bogusław Cyganek. Embedding of the extended Euclidean distance into pattern recognition with higher-order singular value decomposition of prototype tensors. Lecture Notes in Computer Science, 7564: 180–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33260-9_15/.

Porwik:2012:DVB

[268] Piotr Porwik and Tomasz Orczyk. DTW and voting-based lip print recognition system. Lecture Notes in Computer Science, 7564:191–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33260-9_16/.

Rybnik:2012:AKD

[269] Mariusz Rybnik, Piotr Panasiuk, Khalid Saeed, and Marcin Rogowski. Advances in the keystroke dynamics: The practical impact of database quality. Lecture Notes in Computer Science, 7564:203–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33260-9_17/.

Penhaker:2012:AIB

[270] Marek Penhaker, Petr Klimes, Jakub Pindor, and David Korpas. Advanced intracardial biosignal processing. Lecture Notes in Computer Science, 7564:215–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33260-9_18/.

Mliki:2012:MCF

[271] Hazar Mliki, Mohamed Hammami, and Hanêne Ben-Abdallah. Multiconstraints face detect-track system. Lecture Notes in Computer Science, 7564:224–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33260-9_19/.

Augustynek:2012:UDP

[272] Martin Augustynek, Ondrej Adamec, and David Micanik. Using a differential pressure sensor as spirometer. Lecture Notes in Computer Science, 7564:236–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33260-9_20/.

Anonymous:2012:FMp

[273] Anonymous. Front matter. Lecture Notes in Computer Science, 7564: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33260-9/1.

Raedt:2012:DMM

[274] Luc De Raedt. Declarative modeling for machine learning and data mining. Lecture Notes in Computer Science, 7569:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-33492-4_1.

Calders:2012:RDP

[275] Toon Calders. Recent developments in pattern mining. Lecture Notes in Computer Science, 7569:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-33492-4_2.

Ritschard:2012:ESD

[276] Gilbert Ritschard. Exploring sequential data. Lecture Notes in Computer Science, 7569:3–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33492-4_3/.

Khoa:2012:LSS

[277] Nguyen Lu Dang Khoa and Sanjay Chawla. Large scale spectral clustering using resistance distance and Spielman–Teng solvers. Lecture Notes in Computer Science, 7569:7–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33492-4_4/.

Alquier:2012:PQS

[278] Pierre Alquier and Xiaoyin Li. Prediction of quantiles by statistical learning and application to GDP forecasting. Lecture Notes in Computer Science, 7569:22–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33492-4_5/.

Maes:2012:PSS

[279] Francis Maes, Raphael Fonteneau, Louis Wehenkel, and Damien Ernst. Policy search in a space of simple closed-form formulas: Towards interpretability of reinforcement learning. Lecture Notes in Computer Science, 7569:37–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33492-4_6/.

Galbrun:2012:TFR

[280] Esther Galbrun and Angelika Kimmig. Towards finding relational redescriptions. Lecture Notes in Computer Science, 7569:52–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33492-4_7/.

Gamberger:2012:DMS

[281] Dragan Gamberger, Dražen Lučanin, and Tomislav Šmuc. Descriptive modeling of systemic banking crises. Lecture Notes in Computer Science, 7569:67–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33492-4_8/.

Makino:2012:TDB

[282] Shunsuke Makino, Takaharu Shimada, Kouichi Hirata, Kouki Yonezawa, and Kimihito Ito. A trim distance between positions in nucleotide sequences. Lecture Notes in Computer Science, 7569:81–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33492-4_9/.

Suzuki:2012:DSH

[283] Einoshin Suzuki, Emi Matsumoto, and Asuki Kouno. Data squashing for HSV subimages by an autonomous mobile robot. Lecture Notes in Computer Science, 7569:95–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33492-4_10/.

Desmier:2012:CCE

[284] Elise Desmier, Marc Plantevit, Céline Robardet, and Jean-François Boulicaut. Cohesive co-evolution patterns in dynamic attributed graphs. Lecture Notes in Computer Science, 7569:110–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33492-4_11/.

Li:2012:ERR

[285] Rui Li and Stefan Kramer. Efficient redundancy reduced subgroup discovery via quadratic programming. Lecture Notes in Computer Science, 7569:125–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33492-4_12/.

Nogueira:2012:HSS

[286] Bruno M. Nogueira, Alípio M. Jorge, and Solange O. Rezende. HCAC: Semi-supervised hierarchical clustering using confidence-based active learning. Lecture Notes in Computer Science, 7569:139–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33492-4_13/.

Dai:2012:LCL

[287] Bi-Ru Dai and Chih-Heng Chung. LF-CARS: a loose fragment-based consensus clustering algorithm with a robust similarity. Lecture Notes in Computer Science, 7569:154–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33492-4_14/.

Tabata:2012:FAA

[288] Koji Tabata, Atsuyoshi Nakamura, and Mineichi Kudo. Fast approximation algorithm for the 1-median problem. Lecture Notes in Computer Science, 7569:169–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33492-4_15/.

deRuijter:2012:OCR

[289] Tom de Ruijter, Evgeni Tsivtsivadze, and Tom Heskes. Online coregularized algorithms. Lecture Notes in Computer Science, 7569:184– 193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33492-4_16/.

Adhikari:2012:FPT

[290] Prem Raj Adhikari and Jaakko Hollmén. Fast progressive training of mixture models for model selection. Lecture Notes in Computer Science, 7569:194–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33492-4_17/.

Fabregue:2012:ISR

[291] Mickaël Fabrègue, Agnès Braud, Sandra Bringay, Florence Le Ber, and Maguelonne Teisseire. Including spatial relations and scales within sequential pattern extraction. Lecture Notes in Computer Science, 7569: 209–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33492-4_18/.

Ferreira:2012:PRE

[292] Carlos Abreu Ferreira, João Gama, Vítor Santos Costa, Vladimiro Miranda, and Audun Botterud. Predicting ramp events with a streambased HMM framework. Lecture Notes in Computer Science, 7569:

224–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33492-4_19/.

Saito:2012:BDS

[293] Kazumi Saito, Kouzou Ohara, Masahiro Kimura, and Hiroshi Motoda. Burst detection in a sequence of tweets based on information diffusion model. Lecture Notes in Computer Science, 7569:239–253, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33492-4_20/.

Anonymous:2012:FMq

[294] Anonymous. Front matter. Lecture Notes in Computer Science, 7569: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33492-4/1.

Guizard:2012:STR

[295] Nicolas Guizard, Vladimir S. Fonov, Daniel García-Lorenzo, and Bérengère Aubert-Broche. Spatio-temporal regularization for longitudinal registration to an unbiased 3D individual template. Lecture Notes in Computer Science, 7570:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33555-6_1/.

Fiot:2012:LVG

[296] Jean-Baptiste Fiot, Laurent Risser, Laurent D. Cohen, and Jurgen Fripp. Local vs global descriptors of hippocampus shape evolution for Alzheimer’s longitudinal population analysis. Lecture Notes in Computer Science, 7570:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33555-6_2/.

Duchateau:2012:WRF

[297] Nicolas Duchateau, Mathieu De Craene, Xavier Pennec, Beatriz Merino, and Marta Sitges. Which reorientation framework for the atlas-based comparison of motion from cardiac image sequences? Lecture Notes in Computer Science, 7570:25–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33555-6_3/.

Pearlman:2012:EDC

[298] Paul C. Pearlman, Ivana Išgum, Karina J. Kersbergen, and Manon J. N. L. Benders. Elastic demons: Characterizing cortical development in neonates using an implicit surface registration. Lecture Notes in Computer Science, 7570:38–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33555-6_4/.

Aubert-Broche:2012:NFA

[299] Bérengère Aubert-Broche, Vladimir S. Fonov, Daniel García-Lorenzo, and Abderazzak Mouiha. A new framework for analyzing structural volume changes of longitudinal brain MRI data. Lecture Notes in Computer Science, 7570:50–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33555-6_5/.

Wang:2012:SLB

[300] Li Wang, Feng Shi, Gang Li, and Dinggang Shen. 4D segmentation of longitudinal brain MR images with consistent cortical thickness measurement. Lecture Notes in Computer Science, 7570:63–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33555-6_6/.

Datar:2012:MES

[301] Manasi Datar, Prasanna Muralidharan, Abhishek Kumar, Sylvain Gouttard, and Joseph Piven. Mixed-effects shape models for estimating longitudinal changes in anatomy. Lecture Notes in Computer Science, 7570: 76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_7/.

Serag:2012:ULS

[302] Ahmed Serag, Ioannis S. Gousias, Antonios Makropoulos, Paul Aljabar, and Joseph V. Hajnal. Unsupervised learning of shape complexity: Application to brain development. Lecture Notes in Computer Science, 7570: 88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_8/.

Jiao:2012:STP

[303] Jieqing Jiao, Graham E. Searle, Andri C. Tziortzi, Cristian A. Salinas, and Roger N. Gunn. Spatial-temporal pharmacokinetic model based registration of 4D brain PET data. Lecture Notes in Computer Science, 7570: 100–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_9/.

Stretton:2012:PLG

[304] Erin Stretton, Emmanuel Mandonnet, Ezequiel Geremia, Bjoern H. Menze, and Hervé Delingette. Predicting the location of glioma recurrence after a resection surgery. Lecture Notes in Computer Science, 7570: 113–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_10/.

Chitphakdithai:2012:TMB

[305] Nicha Chitphakdithai, Veronica L. Chiang, and James S. Duncan. Tracking metastatic brain tumors in longitudinal scans via joint image registration and labeling. Lecture Notes in Computer Science, 7570:124– 136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_11/.

Liu:2012:MBS

[306] Xiaofeng Liu and Dirk Padfield. Motion-based segmentation for cardiomyocyte characterization. Lecture Notes in Computer Science, 7570: 137–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_12/.

Narayanaswamy:2012:MTG

[307] Arunachalam Narayanaswamy, Amine Merouane, Antonio Peixoto, Ena Ladi, and Paul Herzmark. Multi-temporal globally-optimal dense 3-D cell segmentation and tracking from multi-photon time-lapse movies of live tissue microenvironments. Lecture Notes in Computer Science, 7570: 147–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33555-6_13/.

Anonymous:2012:BMe

[308] Anonymous. Back matter. Lecture Notes in Computer Science, 7570: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33555-6/1.

Anonymous:2012:FMr

[309] Anonymous. Front matter. Lecture Notes in Computer Science, 7570: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33555-6/1.

Freifeld:2012:LBM

[310] Oren Freifeld and Michael J. Black. Lie bodies: a manifold representation of 3D human shape. Lecture Notes in Computer Science, 7572: 1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33718-5_1/.

Li:2012:WPE

[311] Yunpeng Li, Noah Snavely, Dan Huttenlocher, and Pascal Fua. Worldwide pose estimation using 3D point clouds. Lecture Notes in Computer Science, 7572:15–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33718-5_2/.

Beeler:2012:IRD

[312] Thabo Beeler, Derek Bradley, Henning Zimmer, and Markus Gross. Improved reconstruction of deforming surfaces by cancelling ambient occlusion. Lecture Notes in Computer Science, 7572:30–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33718-5_3/.

Liu:2012:SDO

[313] Junbin Liu, Clinton Fookes, Tim Wark, and Sridha Sridharan. On the statistical determination of optimal camera configurations in large scale surveillance networks. Lecture Notes in Computer Science, 7572: 44–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33718-5_4/.

Oxholm:2012:SGT

[314] Geoffrey Oxholm, Prabin Bariya, and Ko Nishino. The scale of geometric texture. Lecture Notes in Computer Science, 7572:58–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33718-5_5/.

Valmadre:2012:EAT

[315] Jack Valmadre, Yingying Zhu, Sridha Sridharan, and Simon Lucey. Efficient articulated trajectory reconstruction using dynamic programming and filters. Lecture Notes in Computer Science, 7572:72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33718-5_6/.

Bao:2012:OCD

[316] Sid Yingze Bao, Yu Xiang, and Silvio Savarese. Object co-detection. Lecture Notes in Computer Science, 7572:86–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33718-5_7/.

Li:2012:MDF

[317] Shaoxin Li, Xin Liu, Xiujuan Chai, Haihong Zhang, Shihong Lao, and Shiguang Shan. Morphable displacement field based image matching for face recognition across pose. Lecture Notes in Computer Science, 7572: 102–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33718-5_8/.

Khamis:2012:CPF

[318] Sameh Khamis, Vlad I. Morariu, and Larry S. Davis. Combining per-frame and per-track cues for multi-person action recognition. Lecture Notes in Computer Science, 7572:116–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33718-5_9/.

Lucchi:2012:JIW

[319] Aurelien Lucchi and Jason Weston. Joint image and word sense discrimination for image retrieval. Lecture Notes in Computer Science, 7572: 130–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33718-5_10/.

Rohrbach:2012:SDA

[320] Marcus Rohrbach, Michaela Regneri, Mykhaylo Andriluka, Sikandar Amin, and Manfred Pinkal. Script data for attribute-based recognition of composite activities. Lecture Notes in Computer Science, 7572: 144–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33718-5_11/.

Khosla:2012:UDD

[321] Aditya Khosla, Tinghui Zhou, Tomasz Malisiewicz, Alexei A. Efros, and Antonio Torralba. Undoing the damage of dataset bias. Lecture Notes in Computer Science, 7572:158–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33718-5_12/.

Liu:2012:DBC

[322] Jiongxin Liu, Angjoo Kanazawa, David Jacobs, and Peter Belhumeur. Dog breed classification using part localization. Lecture Notes in Computer Science, 7572:172–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33718-5_13/.

Kong:2012:DLA

[323] Shu Kong and Donghui Wang. A dictionary learning approach for classification: Separating the particularity and the commonality. Lecture Notes in Computer Science, 7572:186–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33718-5_14/.

Holzer:2012:LED

[324] Stefan Holzer, Jamie Shotton, and Pushmeet Kohli. Learning to efficiently detect repeatable interest points in depth data. Lecture Notes in Computer Science, 7572:200–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33718-5_15/.

Fernando:2012:EUF

[325] Basura Fernando, Elisa Fromont, and Tinne Tuytelaars. Effective use of frequent itemset mining for image classification. Lecture Notes in Computer Science, 7572:214–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33718-5_16/.

Trzcinski:2012:EDP

[326] Tomasz Trzcinski and Vincent Lepetit. Efficient discriminative projections for compact binary descriptors. Lecture Notes in Computer Science, 7572: 228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33718-5_17/.

Simonyan:2012:DLU

[327] Karen Simonyan, Andrea Vedaldi, and Andrew Zisserman. Descriptor learning using convex optimisation. Lecture Notes in Computer Science, 7572:243–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33718-5_18/.

Narayanan:2012:BPO

[328] Maruthi Narayanan and Benjamin Kimia. Bottom-up perceptual organization of images into object part hypotheses. Lecture Notes in Computer Science, 7572:257–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33718-5_19/.

Kim:2012:MGC

[329] Kwang In Kim, James Tompkin, Martin Theobald, Jan Kautz, and Christian Theobalt. Match graph construction for large image databases. Lecture Notes in Computer Science, 7572:272–285, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33718-5_20/.

Anonymous:2012:FMs

[330] Anonymous. Front matter. Lecture Notes in Computer Science, 7572: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33718-5/1.

Hinkle:2012:PRR

[331] Jacob Hinkle, Prasanna Muralidharan, P. Thomas Fletcher, and Sarang Joshi. Polynomial regression on Riemannian manifolds. Lecture Notes in Computer Science, 7574:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33712-3_1/.

Avrithis:2012:AGM

[332] Yannis Avrithis and Yannis Kalantidis. Approximate Gaussian mixtures for large scale vocabularies. Lecture Notes in Computer Science, 7574: 15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33712-3_2/.

Wei:2012:GSU

[333] Yichen Wei, Fang Wen, Wangjiang Zhu, and Jian Sun. Geodesic saliency using background priors. Lecture Notes in Computer Science, 7574: 29–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33712-3_3/.

Smith:2012:JFA

[334] Brandon M. Smith and Li Zhang. Joint face alignment with nonparametric shape models. Lecture Notes in Computer Science, 7574: 43–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33712-3_4/.

Martins:2012:DBA

[335] Pedro Martins, Rui Caseiro, João F. Henriques, and Jorge Batista. Discriminative Bayesian active shape models. Lecture Notes in Computer Science, 7574:57–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33712-3_5/.

Aodha:2012:PBS

[336] Oisin Mac Aodha, Neill D. F. Campbell, Arun Nair, and Gabriel J. Brostow. Patch based synthesis for single depth image super-resolution. Lecture Notes in Computer Science, 7574:71–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33712-3_6/.

Rubinstein:2012:APL

[337] Michael Rubinstein, Ce Liu, and William T. Freeman. Annotation propagation in large image databases via dense image correspondence. Lecture Notes in Computer Science, 7574:85–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33712-3_7/.

Kuang:2012:NSO

[338] Yubin Kuang and Kalle Åström. Numerically stable optimization of polynomial solvers for minimal problems. Lecture Notes in Computer Science, 7574:100–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33712-3_8/.

Jammalamadaka:2012:MAS

[339] Nataraj Jammalamadaka, Andrew Zisserman, Marcin Eichner, and Vittorio Ferrari. Has my algorithm succeeded? An evaluator for human pose estimators. Lecture Notes in Computer Science, 7574:114– 128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33712-3_9/.

Duan:2012:GTE

[340] Genquan Duan, Haizhou Ai, Song Cao, and Shihong Lao. Group tracking: Exploring mutual relations for multiple object tracking. Lecture Notes in Computer Science, 7574:129–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33712-3_10/.

Kausler:2012:DCG

[341] Bernhard X. Kausler, Martin Schiegg, Bjoern Andres, Martin Lindner, and Ullrich Koethe. A discrete chain graph model for 3d + t cell tracking with high misdetection robustness. Lecture Notes in Computer Science, 7574:144–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33712-3_11/.

Yao:2012:RTW

[342] Rui Yao, Qinfeng Shi, Chunhua Shen, Yanning Zhang, and Anton van den Hengel. Robust tracking with weighted online structured learning. Lecture Notes in Computer Science, 7574:158–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33712-3_12/.

Rosman:2012:FRM

[343] Guy Rosman, Yu Wang, Xue-Cheng Tai, Ron Kimmel, and Alfred M. Bruckstein. Fast regularization of matrix-valued images. Lecture Notes in Computer Science, 7574:173–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33712-3_13/.

Schuler:2012:BCO

[344] Christian J. Schuler, Michael Hirsch, Stefan Harmeling, and Bernhard Schölkopf. Blind correction of optical aberrations. Lecture Notes in Computer Science, 7574:187–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33712-3_14/.

Aldrian:2012:IRF

[345] Oswald Aldrian and William A. P. Smith. Inverse rendering of faces on a cloudy day. Lecture Notes in Computer Science, 7574:201–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33712-3_15/.

Aastrom:2012:TBP

[346] Freddie Åström, George Baravdish, and Michael Felsberg. On tensorbased PDEs and their corresponding variational formulations with application to color image denoising. Lecture Notes in Computer Science, 7574: 215–228, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33712-3_16/.

Gong:2012:KTC

[347] Dian Gong, Gérard Medioni, Sikai Zhu, and Xuemei Zhao. Kernelized temporal cut for online temporal segmentation and recognition. Lecture Notes in Computer Science, 7574:229–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33712-3_17/.

Cao:2012:GSS

[348] Yu Cao, Lili Ju, and Song Wang. Grain segmentation of 3D superalloy images using multichannel EWCVT under human annotation constraints. Lecture Notes in Computer Science, 7574:244–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33712-3_18/.

Riemenschneider:2012:HRJ

[349] Hayko Riemenschneider, Sabine Sternig, Michael Donoser, Peter M. Roth, and Horst Bischof. Hough regions for joining instance localization and segmentation. Lecture Notes in Computer Science, 7574: 258–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33712-3_19/.

Kowdle:2012:LSV

[350] Adarsh Kowdle and Tsuhan Chen. Learning to segment a video to clips based on scene and camera motion. Lecture Notes in Computer Science, 7574:272–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33712-3_20/.

Anonymous:2012:FMt

[351] Anonymous. Front matter. Lecture Notes in Computer Science, 7574: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33712-3/1.

Sharma:2012:LHO

[352] Gaurav Sharma, Sibt ul Hussain, and Frédéric Jurie. Local higher-order statistics (LHS) for texture categorization and facial analysis. Lecture Notes in Computer Science, 7578:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33786-4_1/.

VandenBergh:2012:SSE

[353] Michael Van den Bergh, Xavier Boix, Gemma Roig, Benjamin de Capitani, and Luc Van Gool. SEEDS: Superpixels extracted via energy-driven sampling. Lecture Notes in Computer Science, 7578:13–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33786-4_2/.

Kohler:2012:RPC

[354] Rolf Köhler, Michael Hirsch, Betty Mohler, Bernhard Schölkopf, and Stefan Harmeling. Recording and playback of camera shake: Benchmarking blind deconvolution with a real-world database. Lecture Notes in Computer Science, 7578:27–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33786-4_3/.

Tsogkas:2012:LBS

[355] Stavros Tsogkas and Iasonas Kokkinos. Learning-based symmetry detection in natural images. Lecture Notes in Computer Science, 7578: 41–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33786-4_4/.

Shapovalova:2012:SCL

[356] Nataliya Shapovalova, Arash Vahdat, Kevin Cannons, Tian Lan, and Greg Mori. Similarity constrained latent support vector machine: An application to weakly supervised action classification. Lecture Notes in Computer Science, 7578:55–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33786-4_5/.

Direkoglu:2012:TAR

[357] Cem Direkoǧlu and Noel E. O’Connor. Team activity recognition in sports. Lecture Notes in Computer Science, 7578:69–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33786-4_6/.

Vig:2012:SVD

[358] Eleonora Vig, Michael Dorr, and David Cox. Space-variant descriptor sampling for action recognition based on saliency and eye movements. Lecture Notes in Computer Science, 7578:84–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33786-4_7/.

Nicolaou:2012:DPC

[359] Mihalis A. Nicolaou, Vladimir Pavlovic, and Maja Pantic. Dynamic probabilistic CCA for analysis of affective behaviour. Lecture Notes in Computer Science, 7578:98–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33786-4_8/.

Jancsary:2012:LST

[360] Jeremy Jancsary, Sebastian Nowozin, and Carsten Rother. Loss-specific training of non-parametric image restoration models: a new state of the art. Lecture Notes in Computer Science, 7578:112–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33786-4_9/.

Wang:2012:PAR

[361] Naiyan Wang, Tiansheng Yao, Jingdong Wang, and Dit-Yan Yeung. A probabilistic approach to robust matrix factorization. Lecture Notes in Computer Science, 7578:126–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33786-4_10/.

Patz:2012:FPS

[362] Torben Pätz and Tobias Preusser. Fast parameter sensitivity analysis of PDE-based image processing methods. Lecture Notes in Computer Science, 7578:140–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33786-4_11/.

Andres:2012:LFE

[363] Bjoern Andres, Jörg H. Kappes, Thorsten Beier, Ullrich Köthe, and Fred A. Hamprecht. The lazy flipper: Efficient depth-limited exhaustive search in discrete graphical models. Lecture Notes in Computer Science, 7578:154–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33786-4_12/.

Du:2012:FAA

[364] Ming Du and Rama Chellappa. Face association across unconstrained video frames using conditional random fields. Lecture Notes in Computer Science, 7578:167–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33786-4_13/.

Woodford:2012:CMG

[365] Oliver J. Woodford, Minh-Tri Pham, Atsuto Maki, Riccardo Gherardi, and Frank Perbet. Contraction moves for geometric model fitting. Lecture Notes in Computer Science, 7578:181–194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33786-4_14/.

Strelow:2012:GNW

[366] Dennis Strelow. General and nested Wiberg minimization: L_{2} and maximum likelihood. Lecture Notes in Computer Science, 7578:195– 207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33786-4_15/.

Strekalovskiy:2012:NPC

[367] Evgeny Strekalovskiy, Claudia Nieuwenhuis, and Daniel Cremers. Nonmetric priors for continuous multilabel optimization. Lecture Notes in Computer Science, 7578:208–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33786-4_16/.

Handa:2012:RTC

[368] Ankur Handa, Richard A. Newcombe, Adrien Angeli, and Andrew J. Davison. Real-time camera tracking: When is high frame-rate best? Lecture Notes in Computer Science, 7578:222–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33786-4_17/.

Tappen:2012:BAA

[369] Marshall F. Tappen and Ce Liu. A Bayesian approach to alignmentbased image hallucination. Lecture Notes in Computer Science, 7578: 236–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33786-4_18/.

Sanchez-Lozano:2012:CRN

[370] Enrique Sánchez-Lozano, Fernando De la Torre, and Daniel González-Jiménez. Continuous regression for non-rigid image alignment. Lecture Notes in Computer Science, 7578:250–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33786-4_19/.

Rouhani:2012:NRS

[371] Mohammad Rouhani and Angel D. Sappa. Non-rigid shape registration: a single linear least squares framework. Lecture Notes in Computer Science, 7578:264–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33786-4_20/.

Anonymous:2012:FMu

[372] Anonymous. Front matter. Lecture Notes in Computer Science, 7578: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33786-4/1.

Aichholzer:2012:GH

[373] Oswin Aichholzer, Thomas Hackl, and Birgit Vogtenhuber. On 5-gons and 5-holes. Lecture Notes in Computer Science, 7579:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34191-5_1/.

Akiyama:2012:RAP

[374] Jin Akiyama, Ikuro Sato, and Hyunwoo Seong. On reversibility among parallelohedra. Lecture Notes in Computer Science, 7579:14– 28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34191-5_2/.

Bose:2012:HFC

[375] Prosenjit Bose and Sander Verdonschot. A history of flips in combinatorial triangulations. Lecture Notes in Computer Science, 7579: 29–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34191-5_3/.

Pach:2012:TT

[376] János Pach, Radoš Radoičić, and Géza Tóth. Tangled thrackles. Lecture Notes in Computer Science, 7579:45–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-34191-5_4/.

Toth:2012:OGE

[377] Csaba D. Tóth, Godfried T. Toussaint, and Andrew Winslow. Open guard edges and edge guards in simple polygons. Lecture Notes in Computer Science, 7579:54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34191-5_5/.

ORourke:2012:SWR

[378] Joseph O’Rourke. String-wrapped rotating disks. Lecture Notes in Computer Science, 7579:65–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34191-5_6/.

Fabila-Monroy:2012:CNC

[379] Ruy Fabila-Monroy and David R. Wood. The chromatic number of the convex segment disjointness graph. Lecture Notes in Computer Science, 7579:79–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34191-5_7/.

Itoh:2012:CFC

[380] Jin ichi Itoh, Chie Nara, and Costin Vîlcu. Continuous flattening of convex polyhedra. Lecture Notes in Computer Science, 7579:85–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-34191-5_8/.

Aichholzer:2012:CMP

[381] Oswin Aichholzer, Mario Cetina, Ruy Fabila-Monroy, and Jesús Leaños. Convexifying monotone polygons while maintaining internal visibility. Lecture Notes in Computer Science, 7579:98–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.com/chapter/10.1007/978-3-642-34191-5_9/.

Diaz-Banez:2012:NRO

[382] José M. Díaz-Báñez, Ruy Fabila-Monroy, and Pablo Pérez-Lantero. On the number of radial orderings of colored planar point sets. Lecture Notes in Computer Science, 7579:109–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34191-5_10/.

Omana-Pulido:2012:NTG

[383] Elsa Omaña-Pulido and Eduardo Rivera-Campo. Notes on the twisted graph. Lecture Notes in Computer Science, 7579:119–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-34191-5_11/.

Diaz-Banez:2012:LSF

[384] José Miguel Díaz-Báñez, Matias Korman, and Pablo Pérez-Lantero. Locating a service facility and a rapid transit line. Lecture Notes in Computer Science, 7579:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34191-5_12/.

Souvaine:2012:SFE

[385] Diane L. Souvaine, Csaba D. Tóth, and Andrew Winslow. Simultaneously flippable edges in triangulations. Lecture Notes in Computer Science, 7579:138–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34191-5_13/.

Iwerks:2012:SSP

[386] Justin Iwerks and Joseph S. B. Mitchell. Spiral serpentine polygonization of a planar point set. Lecture Notes in Computer Science, 7579: 146–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34191-5_14/.

Diaz-Banez:2012:CHP

[387] José Miguel Díaz-Báñez, Matias Korman, and Pablo Pérez-Lantero. The 1-center and 1-highway problem. Lecture Notes in Computer Science, 7579:155–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-34191-5_15/.

Caceres:2012:CGR

[388] José Cáceres, Carmen Cortés, and Clara Isabel Grima. Compact grid representation of graphs. Lecture Notes in Computer Science, 7579: 166–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34191-5_16/.

Sakai:2012:HID

[389] Toshinori Sakai and Jorge Urrutia. On the heaviest increasing or decreasing subsequence of a permutation, and paths and matchings on weighted point sets. Lecture Notes in Computer Science, 7579:175– 184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34191-5_17/.

Demaine:2012:GSU

[390] Erik D. Demaine and Anna Lubiw. A generalization of the source unfolding of convex polyhedra. Lecture Notes in Computer Science, 7579: 185–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-34191-5_18/.

Angelini:2012:LAC

[391] Patrizio Angelini, Giuseppe Di Battista, and Walter Didimo. Large angle crossing drawings of planar graphs in subquadratic area. Lecture Notes in Computer Science, 7579:200–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-34191-5_19/.

Abellanas:2012:CRC

[392] Manuel Abellanas, Antonio L. Bajuelos, and Santiago Canales. Connecting red cells in a bicolour Voronoi diagram. Lecture Notes in Computer Science, 7579:210–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-34191-5_20/.

Anonymous:2012:FMv

[393] Anonymous. Front matter. Lecture Notes in Computer Science, 7579: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-34191-5/1.

Holmgren:2012:MAB

[394] Johan Holmgren, Linda Ramstedt, Paul Davidsson, and Jan A. Persson. Multi-agent-based simulation for analysis of transport policy and infrastructure measures. Lecture Notes in Computer Science, 7580:1– 15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-35612-4_1/.

Savarimuthu:2012:ABS

[395] Bastin Tony Roy Savarimuthu, Maryam Purvis, and Harko Verhagen. An agent-based simulation of employing social norms in energy conservation in households. Lecture Notes in Computer Science, 7580: 16–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35612-4_2/.

Pfau:2012:ABM

[396] Jens Pfau, Michael Kirley, and Yoshihisa Kashima. An agent-based model of stereotype communication. Lecture Notes in Computer Science, 7580: 32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35612-4_3/.

Fukuta:2012:ASE

[397] Naoki Fukuta and Takayuki Ito. An approach to sustainable electric power allocation using a multi-round multi-unit combinatorial auction. Lecture Notes in Computer Science, 7580:48–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35612-4_4/.

Takahashi:2012:CDV

[398] Satoshi Takahashi, Tokuro Matsuo, Takayuki Ito, and Roger Y. Lee. A co-dependent value-based mechanism for the Internet advertisement auction. Lecture Notes in Computer Science, 7580:64–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35612-4_5/.

Kanamori:2012:ESL

[399] Ryo Kanamori, Takayuki Morikawa, and Takayuki Ito. Evaluation of special lanes as incentive policies for promoting electric vehicles. Lecture Notes in Computer Science, 7580:78–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-35612-4_6/.

delaHoz:2012:SCA

[400] Enrique de la Hoz, Ivan Marsa-Maestre, and Miguel A. Lopez-Carmona. Simulation of coordinated anticipatory vehicle routing strategies on MAT-Sim. Lecture Notes in Computer Science, 7580:90–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-35612-4_7/.

Ren:2012:ABD

[401] Fenghui Ren, Minjie Zhang, and Danny Soetanto. Agent-based demand management in a power distribution network by considering distributed generations. Lecture Notes in Computer Science, 7580:109– 124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-35612-4_8/.

Anonymous:2012:BMf

[402] Anonymous. Back matter. Lecture Notes in Computer Science, 7580: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-35612-4/1.

Anonymous:2012:FMw

[403] Anonymous. Front matter. Lecture Notes in Computer Science, 7580: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-35612-4/1.

Litany:2012:PPT

[404] Or Litany, Alexander M. Bronstein, and Michael M. Bronstein. Putting the pieces together: Regularized multi-part shape matching. Lecture Notes in Computer Science, 7583:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33863-2_1/.

VanEyndhoven:2012:CME

[405] Geert Van Eyndhoven, Jan Sijbers, and Joost Batenburg. Combined motion estimation and reconstruction in tomography. Lecture Notes in Computer Science, 7583:12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33863-2_2/.

Abdelrahman:2012:OCU

[406] Mostafa Abdelrahman, Moumen El-Melegy, and Aly Farag. 3D object classification using scale invariant heat kernels with collaborative classification. Lecture Notes in Computer Science, 7583:22–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33863-2_3/.

Sukno:2012:FLL

[407] Federico M. Sukno, John L. Waddington, and Paul F. Whelan. 3D facial landmark localization using combinatorial search and shape regression. Lecture Notes in Computer Science, 7583:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33863-2_4/.

Raviv:2012:SSA

[408] Tammy Riklin Raviv, Yi Gao, James J. Levitt, and Sylvain Bouix. Statistical shape analysis for population studies via level-set based shape morphing. Lecture Notes in Computer Science, 7583:42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33863-2_5/.

Rosman:2012:GVR

[409] Guy Rosman, Alex M. Bronstein, Michael M. Bronstein, Xue-Cheng Tai, and Ron Kimmel. Group-valued regularization for analysis of articulated motion. Lecture Notes in Computer Science, 7583:52–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33863-2_6/.

Marvaniya:2012:DAS

[410] Smit Marvaniya, Sreyasee Bhattacharjee, Venkatesh Manickavasagam, and Anurag Mittal. Drawing an automatic sketch of deformable objects using only a few images. Lecture Notes in Computer Science, 7583: 63–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33863-2_7/.

Berretti:2012:SSR

[411] Stefano Berretti, Alberto Del Bimbo, and Pietro Pala. Superfaces: a super-resolution model for 3D faces. Lecture Notes in Computer Science, 7583:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33863-2_8/.

Kovnatsky:2012:SSM

[412] Artiom Kovnatsky, Michael M. Bronstein, and Alexander M. Bronstein. Stable spectral mesh filtering. Lecture Notes in Computer Science, 7583: 83–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33863-2_9/.

Uchida:2012:ADP

[413] Seiichi Uchida, Satoshi Hokahori, and Yaokai Feng. Analytical dynamic programming matching. Lecture Notes in Computer Science, 7583: 92–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33863-2_10/.

Yang:2012:CPF

[414] Ying Yang, David Günther, Stefanie Wuhrer, Alan Brunton, and Ioannis Ivrissimtzis. Correspondences of persistent feature points on nearisometric surfaces. Lecture Notes in Computer Science, 7583:102–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33863-2_11/.

Agudo:2012:RNR

[415] Antonio Agudo, Begoña Calvo, and J. M. M. Montiel. 3D reconstruction of non-rigid surfaces in real-time using wedge elements. Lecture Notes in Computer Science, 7583:113–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33863-2_12/.

Iglesias:2012:SDS

[416] Jose A. Iglesias and Ron Kimmel. Schrödinger diffusion for shape analysis with texture. Lecture Notes in Computer Science, 7583:123–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33863-2_13/.

Cheng:2012:ADF

[417] Xin Cheng, Sridha Sridharan, Jason Saraghi, and Simon Lucey. Anchored deformable face ensemble alignment. Lecture Notes in Computer Science, 7583:133–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33863-2_14/.

Bardelli:2012:MOT

[418] Eleonora Bardelli, Maria Colombo, Andrea Mennucci, and Anthony Yezzi. Multiple object tracking via prediction and filtering with a Sobolevtype metric on curves. Lecture Notes in Computer Science, 7583:143– 152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33863-2_15/.

Kinoshita:2012:FMF

[419] Koichi Kinoshita, Yoshinori Konishi, Masato Kawade, and Hiroshi Murase. Facial model fitting based on perturbation learning and It’s evaluation on challenging real-world diversities images. Lecture Notes in Computer Science, 7583:153–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33863-2_16/.

Zhu:2012:ARL

[420] Jiejie Zhu, Mayank Bansal, Nick Vander Valk, and Hui Cheng. Adaptive rendering for large-scale skyline characterization and matching. Lecture Notes in Computer Science, 7583:163–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33863-2_17/.

Bansal:2012:UWB

[421] Mayank Bansal, Kostas Daniilidis, and Harpreet Sawhney. Ultra-wide baseline facade matching for geo-localization. Lecture Notes in Computer Science, 7583:175–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33863-2_18/.

Hedau:2012:MED

[422] Varsha Hedau, Sudipta N. Sinha, C. Lawrence Zitnick, and Richard Szeliski. A memory efficient discriminative approach for location aided recognition. Lecture Notes in Computer Science, 7583:187–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33863-2_19/.

Hartmann:2012:WSL

[423] Glenn Hartmann, Matthias Grundmann, Judy Hoffman, David Tsai, Vivek Kwatra, and Omid Madani. Weakly supervised learning of object segmentations from Web-scale video. Lecture Notes in Computer Science, 7583:198–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33863-2_20/.

Anonymous:2012:FMx

[424] Anonymous. Front matter. Lecture Notes in Computer Science, 7583: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33863-2/1.

Nair:2012:HAT

[425] Rahul Nair, Frank Lenzen, Stephan Meister, Henrik Schäfer, and Christoph Garbe. High accuracy TOF and stereo sensor fusion at interactive rates. Lecture Notes in Computer Science, 7584:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33868-7_1/.

Langmann:2012:MFM

[426] Benjamin Langmann, Klaus Hartmann, and Otmar Loffeld. A modular framework for 2D/3D and multi-modal segmentation with joint superresolution. Lecture Notes in Computer Science, 7584:12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33868-7_2/.

Wang:2012:RTP

[427] Zhe Wang, Hong Liu, Yueliang Qian, and Tao Xu. Real-time plane segmentation and obstacle detection of 3D point clouds for indoor scenes. Lecture Notes in Computer Science, 7584:22–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33868-7_3/.

Bayramoglu:2012:CTG

[428] Neslihan Bayramogğlu, Janne Heikkilä, and Matti Pietikäinen. Combining textural and geometrical descriptors for scene recognition. Lecture Notes in Computer Science, 7584:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33868-7_4/.

Lu:2012:HCI

[429] Jiwen Lu and Gang Wang. Human-centric indoor environment modeling from depth videos. Lecture Notes in Computer Science, 7584: 42–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33868-7_5/.

Cheng:2012:HDA

[430] Zhongwei Cheng, Lei Qin, Yituo Ye, Qingming Huang, and Qi Tian. Human daily action analysis with multi-view and color-depth data. Lecture Notes in Computer Science, 7584:52–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33868-7_6/.

Zeisl:2012:VIM

[431] Bernhard Zeisl, Kevin Köser, and Marc Pollefeys. Viewpoint invariant matching via developable surfaces. Lecture Notes in Computer Science, 7584:62–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33868-7_7/.

Ren:2012:UEM

[432] Carl Yuheng Ren and Ian Reid. A unified energy minimization framework for model fitting in depth. Lecture Notes in Computer Science, 7584:72–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33868-7_8/.

Fouhey:2012:ORR

[433] David F. Fouhey, Alvaro Collet, Martial Hebert, and Siddhartha Srinivasa. Object recognition robust to imperfect depth data. Lecture Notes in Computer Science, 7584:83–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33868-7_9/.

Susanto:2012:ODM

[434] Wandi Susanto, Marcus Rohrbach, and Bernt Schiele. 3D object detection with multiple kinects. Lecture Notes in Computer Science, 7584: 93–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33868-7_10/.

Kowdle:2012:CMG

[435] Adarsh Kowdle, Andrew Gallagher, and Tsuhan Chen. Combining monocular geometric cues with traditional stereo cues for consumer camera stereo. Lecture Notes in Computer Science, 7584:103–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33868-7_11/.

Sellent:2012:QAN

[436] Anita Sellent and Jochen Wingbermühle. Quality assessment of nondense image correspondences. Lecture Notes in Computer Science, 7584: 114–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33868-7_12/.

Marquez-Valle:2012:CCF

[437] Patricia Márquez-Valle, Debora Gil, and Aura Hernàndez-Sabaté. A complete confidence framework for optical flow. Lecture Notes in Computer Science, 7584:124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33868-7_13/.

Cigla:2012:ISM

[438] Cevahir Çığla and A. Aydın Alatan. An improved stereo matching algorithm with ground plane and temporal smoothness constraints. Lecture Notes in Computer Science, 7584:134–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33868-7_14/.

Mordohai:2012:ESF

[439] Philippos Mordohai. On the evaluation of scene flow estimation. Lecture Notes in Computer Science, 7584:148–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33868-7_15/.

Haeusler:2012:AKD

[440] Ralf Haeusler and Reinhard Klette. Analysis of KITTI data for stereo analysis with stereo confidence measures. Lecture Notes in Computer Science, 7584:158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33868-7_16/.

Wulff:2012:LIC

[441] Jonas Wulff, Daniel J. Butler, Garrett B. Stanley, and Michael J. Black. Lessons and insights from creating a synthetic optical flow benchmark. Lecture Notes in Computer Science, 7584:168–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33868-7_17/.

Batool:2012:MDW

[442] Nazre Batool and Rama Chellappa. Modeling and detection of wrinkles in aging human faces using marked point processes. Lecture Notes in Computer Science, 7584:178–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33868-7_18/.

Otto:2012:HDA

[443] Charles Otto, Hu Han, and Anil Jain. How does aging affect facial components? Lecture Notes in Computer Science, 7584:189–198, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33868-7_19/.

Riaz:2012:STM

[444] Zahid Riaz and Michael Beetz. Spatio-temporal multifeature for facial analysis. Lecture Notes in Computer Science, 7584:199–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33868-7_20/.

Anonymous:2012:FMy

[445] Anonymous. Front matter. Lecture Notes in Computer Science, 7584: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33868-7/1.

Pfeiffer:2012:PSO

[446] David Pfeiffer, Friedrich Erbs, and Uwe Franke. Pixels, stixels, and objects. Lecture Notes in Computer Science, 7585:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33885-4_1/.

Benenson:2012:FSC

[447] Rodrigo Benenson, Markus Mathias, Radu Timofte, and Luc Van Gool. Fast stixel computation for fast pedestrian detection. Lecture Notes in Computer Science, 7585:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33885-4_2/.

Maji:2012:DLP

[448] Subhransu Maji. Discovering a lexicon of parts and attributes. Lecture Notes in Computer Science, 7585:21–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33885-4_3/.

Divvala:2012:HID

[449] Santosh K. Divvala, Alexei A. Efros, and Martial Hebert. How important are “deformable parts” in the deformable parts model? Lecture Notes in Computer Science, 7585:31–40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33885-4_4/.

Kokkinos:2012:BPS

[450] Iasonas Kokkinos. Bounding part scores for rapid detection with deformable part models. Lecture Notes in Computer Science, 7585:41– 50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33885-4_5/.

Su:2012:LCV

[451] Yu Su and Frédéric Jurie. Learning compact visual attributes for largescale image classification. Lecture Notes in Computer Science, 7585: 51–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33885-4_6/.

Ma:2012:ULD

[452] Shugao Ma, Stan Sclaroff, and Nazli Ikizler-Cinbis. Unsupervised learning of discriminative relative visual attributes. Lecture Notes in Computer Science, 7585:61–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33885-4_7/.

Kosmopoulos:2012:MOA

[453] Dimitrios I. Kosmopoulos and Fillia Makedon. A method for online analysis of structured processes using Bayesian filters and echo state networks. Lecture Notes in Computer Science, 7585:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33885-4_8/.

Makantasis:2012:MCF

[454] Konstantinos Makantasis, Eftychios Protopapadakis, and Anastasios Doulamis. Monocular camera fall detection system exploiting 3D measures: a semi-supervised learning approach. Lecture Notes in Computer Science, 7585:81–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33885-4_9/.

Munsell:2012:PIU

[455] Brent C. Munsell, Andrew Temlyakov, Chengzheng Qu, and Song Wang. Person identification using full-body motion and anthropometric biometrics from Kinect Videos. Lecture Notes in Computer Science, 7585:

91–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33885-4_10/.

Ghamdi:2012:STV

[456] Manal Al Ghamdi, Nouf Al Harbi, and Yoshihiko Gotoh. Spatio-temporal video representation with locality-constrained linear coding. Lecture Notes in Computer Science, 7585:101–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33885-4_11/.

Rota:2012:RTD

[457] Paolo Rota, Nicola Conci, and Nicu Sebe. Real time detection of social interactions in surveillance video. Lecture Notes in Computer Science, 7585:111–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33885-4_12/.

Brkic:2012:TST

[458] Karla Brkić, Axel Pinz, Zoran Kalafatić, and Siniša Šegvić. Towards space-time semantics in two frames. Lecture Notes in Computer Science, 7585:121–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33885-4_13/.

Ravichandran:2012:SML

[459] Avinash Ravichandran, Chaohui Wang, Michalis Raptis, and Stefano Soatto. SuperFloxels: a mid-level representation for video sequences. Lecture Notes in Computer Science, 7585:131–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33885-4_14/.

John:2012:RCL

[460] Vijay John, Gwenn Englebienne, and Ben Krose. Relative camera localisation in non-overlapping camera networks using multiple trajectories. Lecture Notes in Computer Science, 7585:141–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33885-4_15/.

Ito:2012:DIE

[461] Yuichi Ito, Kris M. Kitani, James A. Bagnell, and Martial Hebert. Detecting interesting events using unsupervised density ratio estimation. Lecture Notes in Computer Science, 7585:151–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33885-4_16/.

Pellegrini:2012:DFC

[462] Stefano Pellegrini, Jürgen Gall, Leonid Sigal, and Luc Van Gool. Destination flow for crowd simulation. Lecture Notes in Computer Science, 7585: 162–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33885-4_17/.

Barthelemy:2012:RID

[463] Quentin Barthélemy, Anthony Larue, and Jérôme I. Mars. 3D rotation invariant decomposition of motion signals. Lecture Notes in Computer Science, 7585:172–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33885-4_18/.

Mauthner:2012:LMA

[464] Thomas Mauthner, Peter M. Roth, and Horst Bischof. Learn to move: Activity specific motion models for tracking by detection. Lecture Notes in Computer Science, 7585:183–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33885-4_19/.

Ardo:2012:FCU

[465] Håkan Ardö, Mikael Nilsson, and Rikard Berthilsson. Flow counting using realboosted multi-sized window detectors. Lecture Notes in Computer Science, 7585:193–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33885-4_20/.

Anonymous:2012:FMz

[466] Anonymous. Front matter. Lecture Notes in Computer Science, 7585: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33885-4/1.

Henzinger:2012:QRM

[467] Thomas A. Henzinger. Quantitative reactive models. Lecture Notes in Computer Science, 7590:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-33666-9_1.

Sanchez-Cuadrado:2012:BMM

[468] Jesús Sánchez-Cuadrado, Juan de Lara, and Esther Guerra. Bottom-up meta-modelling: An interactive approach. Lecture Notes in Computer Science, 7590:3–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33666-9_2/.

Noyrit:2012:FMU

[469] Florian Noyrit, Sébastien Gérard, and Bran Selic. FacadeMetamodel: Masking UML. Lecture Notes in Computer Science, 7590:20–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33666-9_3/.

Rabbi:2012:DSL

[470] Fazle Rabbi and Wendy MacCaull. T_{□} : a domain specific language for rapid workflow development. Lecture Notes in Computer Science, 7590: 36–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33666-9_4/.

Ramirez:2012:RCC

[471] Andres J. Ramirez, Betty H. C. Cheng, Nelly Bencomo, and Pete Sawyer. Relaxing claims: Coping with uncertainty while evaluating assumptions at run time. Lecture Notes in Computer Science, 7590:53–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33666-9_5/.

Alferez:2012:DEC

[472] Germán H. Alférez and Vicente Pelechano. Dynamic evolution of contextaware systems with models at runtime. Lecture Notes in Computer Science, 7590:70–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33666-9_6/.

Fouquet:2012:EMF

[473] François Fouquet, Grégory Nain, Brice Morin, Erwan Daubert, and Olivier Barais. An Eclipse modelling framework alternative to meet the models@runtime requirements. Lecture Notes in Computer Science, 7590: 87–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33666-9_7/.

Scheidgen:2012:ATM

[474] Markus Scheidgen, Anatolij Zubow, Joachim Fischer, and Thomas H. Kolbe. Automated and transparent model fragmentation for persisting large models. Lecture Notes in Computer Science, 7590:102–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33666-9_8/.

Combemale:2012:FDI

[475] Benoit Combemale, Xavier Thirioux, and Benoit Baudry. Formally defining and iterating infinite models. Lecture Notes in Computer Science, 7590:119–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33666-9_9/.

Hegedus:2012:QDS

[476] Ábel Hegedüs, Ákos Horváth, István Ráth, and Dániel Varró. Querydriven soft interconnection of EMF models. Lecture Notes in Computer Science, 7590:134–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33666-9_10/.

Favre:2012:MLA

[477] Jean-Marie Favre, Ralf Lämmel, and Andrei Varanovich. Modeling the linguistic architecture of software products. Lecture Notes in Computer Science, 7590:151–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33666-9_11/.

Pfeiffer:2012:CLS

[478] Rolf-Helge Pfeiffer and Andrzej Wasowski. Cross-language support mechanisms significantly aid software development. Lecture Notes in Computer Science, 7590:168–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33666-9_12/.

Gravino:2012:DPD

[479] Carmine Gravino, Michele Risi, Giuseppe Scanniello, and Genoveffa Tortora. Do professional developers benefit from design pattern documentation? A replication in the context of source code comprehension. Lecture Notes in Computer Science, 7590:185–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33666-9_13/.

Reder:2012:ICC

[480] Alexander Reder and Alexander Egyed. Incremental consistency checking for complex design rules and larger model changes. Lecture Notes in Computer Science, 7590:202–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33666-9_14/.

Farias:2012:EIA

[481] Kleinner Farias, Alessandro Garcia, and Carlos Lucena. Evaluating the impact of aspects on inconsistency detection effort: a controlled experiment. Lecture Notes in Computer Science, 7590:219–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33666-9_15/.

Hamann:2012:ISB

[482] Lars Hamann, Oliver Hofrichter, and Martin Gogolla. On integrating structure and behavior modeling with OCL. Lecture Notes in Computer Science, 7590:235–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33666-9_16/.

Schroeter:2012:MPF

[483] Julia Schroeter, Malte Lochau, and Tim Winkelmann. Multi-perspectives on feature models. Lecture Notes in Computer Science, 7590:252–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33666-9_17/.

Johansen:2012:GBP

[484] Martin Fagereng Johansen, Øystein Haugen, Franck Fleurey, and Anne Grete Eldegard. Generating better partial covering arrays by modeling weights on sub-product lines. Lecture Notes in Computer Science, 7590:269–284, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33666-9_18/.

Kulkarni:2012:TBA

[485] Vinay Kulkarni, Souvik Barat, and Suman Roychoudhury. Towards business application product lines. Lecture Notes in Computer Science, 7590: 285–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33666-9_19/.

Maraee:2012:IAC

[486] Azzam Maraee and Mira Balaban. Inter-association constraints in UML2: Comparative analysis, usage recommendations, and modeling guidelines. Lecture Notes in Computer Science, 7590:302–318, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33666-9_20/.

Anonymous:2012:FMba

[487] Anonymous. Front matter. Lecture Notes in Computer Science, 7590: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33666-9/1.

Jongmans:2012:ACG

[488] Sung-Shik T. Q. Jongmans, Francesco Santini, Mahdi Sargolzaei, and Farhad Arbab. Automatic code generation for the orchestration of Web services with reo. Lecture Notes in Computer Science, 7592:1– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33427-6_1/.

Barakat:2012:RSS

[489] Lina Barakat, Simon Miles, and Michael Luck. Reactive service selection in dynamic service environments. Lecture Notes in Computer Science, 7592:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33427-6_2/.

Demchenko:2012:GBS

[490] Yuri Demchenko, Canh Ngo, Pedro Martínez-Julia, Elena Torroglosa, and Mary Grammatikou. GEMBus based services composition platform for cloud PaaS. Lecture Notes in Computer Science, 7592:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33427-6_3/.

Preda:2012:IBS

[491] Mila Dalla Preda, Maurizio Gabbrielli, Claudio Guidi, Jacopo Mauro, and Fabrizio Montesi. Interface-based service composition with aggregation. Lecture Notes in Computer Science, 7592:48–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33427-6_4/.

Collinson:2012:FMS

[492] Matthew Collinson, David Pym, and Barry Taylor. A framework for modelling security architectures in services ecosystems. Lecture Notes in Computer Science, 7592:64–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33427-6_5/.

Wenge:2012:MAA

[493] Olga Wenge, Melanie Siebenhaar, Ulrich Lampe, Dieter Schuller, and Ralf Steinmetz. Much ado about security appeal: Cloud provider collaborations and their risks. Lecture Notes in Computer Science, 7592: 80–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33427-6_6/.

deBoer:2012:FMR

[494] Frank S. de Boer, Reiner Hähnle, Einar Broch Johnsen, Rudolf Schlatte, and Peter Y. H. Wong. Formal modeling of resource management for cloud architectures: An industrial case study. Lecture Notes in Computer Science, 7592:91–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33427-6_7/.

Nhan:2012:MDA

[495] Tam Le Nhan, Gerson Sunyé, and Jean-Marc Jézéquel. A model-driven approach for virtual machine image provisioning in cloud computing. Lecture Notes in Computer Science, 7592:107–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33427-6_8/.

Heike:2012:PCC

[496] Christian Heike, Wolf Zimmermann, and Andreas Both. Protocol conformance checking of services with exceptions. Lecture Notes in Computer Science, 7592:122–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33427-6_9/.

Pahl:2012:CSL

[497] Claus Pahl. Cloud service localisation. Lecture Notes in Computer Science, 7592:138–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33427-6_10/.

Sagbo:2012:QAR

[498] Kouessi Arafat Romaric Sagbo and Pélagie Houngue. Quality architecture for resource allocation in cloud computing. Lecture Notes in Computer Science, 7592:154–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33427-6_11/.

Zivkovic:2012:ARI

[499] Miroslav Živković and Hans van den Berg. Analysis of revenue improvements with runtime adaptation of service composition based on conditional request retries. Lecture Notes in Computer Science, 7592: 169–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33427-6_12/.

Dellkrantz:2012:PMA

[500] Manfred Dellkrantz, Maria Kihl, and Anders Robertsson. Performance modeling and analysis of a database server with write-heavy workload. Lecture Notes in Computer Science, 7592:184–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33427-6_13/.

Ferber:2012:MCC

[501] Marvin Ferber and Thomas Rauber. Mobile cloud computing in 3G cellular networks using pipelined tasks. Lecture Notes in Computer Science, 7592:192–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33427-6_14/.

Fittkau:2012:CUC

[502] Florian Fittkau, Sören Frey, and Wilhelm Hasselbring. Cloud usercentric enhancements of the simulator CloudSim to improve cloud deployment option analysis. Lecture Notes in Computer Science, 7592: 200–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33427-6_15/.

Guidi:2012:POP

[503] Claudio Guidi, Paolo Anedda, and Tullio Vardanega. PaaSSOA: An open PaaS architecture for service oriented applications. Lecture Notes in Computer Science, 7592:208–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-33427-6_16.

Vettor:2012:CML

[504] Pierre De Vettor, Michael Mrissa, and Carlos Pedrinaci. Context mediation as a linked service. Lecture Notes in Computer Science, 7592:210–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/accesspage/chapter/10.1007/978-
3-642-33427-6_17.

Wang:2012:UCP

[505] Ming-Xue Wang and Claus Pahl. User-customisable policy monitoring for multi-tenant cloud architectures. Lecture Notes in Computer Science, 7592:212–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33427-6_18.

Cheng:2012:ECB

[506] Bin Cheng, Sachin Agarwal, and Daniele Abbadessa. Enabling cobrowsing service across different browsers and devices. Lecture Notes in Computer Science, 7592:214–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33427-6_19/.

Trammel:2012:DTP

[507] John Trammel, Ümit Yalçinalp, Andrei Kalfas, James Boag, and Dan Brotsky. Device token protocol for persistent authentication shared across applications. Lecture Notes in Computer Science, 7592:230–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33427-6_20/.

Anonymous:2012:FMbb

[508] Anonymous. Front matter. Lecture Notes in Computer Science, 7592: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33427-6/1.

Ajmal:2012:VST

[509] Muhammad Ajmal, Muhammad Husnain Ashraf, Muhammad Shakir, Yasir Abbas, and Faiz Ali Shah. Video summarization: Techniques and classification. Lecture Notes in Computer Science, 7594:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33564-8_1/.

Bay:2012:DGM

[510] Thierry Bay, Romain Raffin, and Marc Daniel. Discrete geometric modeling of thick pelvic organs with a medial axis. Lecture Notes in Computer Science, 7594:14–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33564-8_2/.

Bielecka:2012:ENA

[511] Marzena Bielecka and Andrzej Bielecki. An evolutionary-neural algorithm for solving inverse IFS problem for images in two-dimensional space. Lecture Notes in Computer Science, 7594:22–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33564-8_3/.

Borkowski:2012:EAI

[512] Dariusz Borkowski. Euler’s approximations to image reconstruction. Lecture Notes in Computer Science, 7594:30–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33564-8_4/.

Borkowski:2012:ABS

[513] Dariusz Borkowski and Katarzyna Jańczak-Borkowska. Application of backward stochastic differential equations to reconstruction of vectorvalued images. Lecture Notes in Computer Science, 7594:38–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33564-8_5/.

Celebi:2012:BNG

[514] M. Emre Celebi, Quan Wen, Gerald Schaefer, and Huiyu Zhou. Batch neural gas with deterministic initialization for color quantization. Lecture Notes in Computer Science, 7594:48–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33564-8_6/.

Serrano:2012:CFD

[515] Eduardo E. Dávila Serrano, Laurent Guigues, Jean-Pierre Roux, and Frédéric Cervenansky. CreaTools: a framework to develop medical image processing software: Application to simulate pipeline stent deployment in intracranial vessels with aneurysms. Lecture Notes in Computer Science, 7594:55–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33564-8_7/.

Domanska:2012:VMD

[516] Diana Domańska, Marek Wojtylak, and Wiesław Kotarski. Visualization of multidimensional data in explorative forecast. Lecture Notes in Computer Science, 7594:63–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33564-8_8/.

Goinski:2012:ASG

[517] Adam Goiński, Tomasz Zawadzki, and Sławomir Nikiel. Automatic shape generation based on quadratic four-dimensional fractals. Lecture Notes in Computer Science, 7594:71–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33564-8_9/.

Grajek:2012:AAO

[518] Tomasz Grajek, Damian Karwowski, Adam Łuczak, Sławomir Maćkowiak, and Marek Domański. Architecture of algorithmically optimized MPEG-4 AVC/H.264 video encoder. Lecture Notes in Computer Science, 7594: 79–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33564-8_10/.

Iwaneczko:2012:PUA

[519] Paweł Iwaneczko, Karol Jedrasiak, Krzysztof Daniec, and Aleksander Nawrat. A prototype of unmanned aerial vehicle for image acquisition. Lecture Notes in Computer Science, 7594:87–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33564-8_11/.

Janik:2012:AHA

[520] Łukasz Janik, Karol Jedrasiak, Konrad Wojciechowski, and Andrzej Polański. Application of a hybrid algorithm for non–humanoid skeleton model estimation from motion capture data. Lecture Notes in Computer Science, 7594:95–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33564-8_12/.

Jopek:2012:NMS

[521] Łukasz Jopek, Laurent Babout, and Marcin Janaszewski. A new method to segment X-ray microtomography images of lamellar titanium alloy based on directional filter banks and gray level gradient. Lecture Notes in Computer Science, 7594:105–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33564-8_13/.

Kapec:2012:IGE

[522] Peter Kapec, Michal Paprčka, and Adam Pažitnaj. Intelligent 3D graph exploration with time-travel features. Lecture Notes in Computer Science, 7594:113–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33564-8_14/.

Karwowski:2012:IAA

[523] Damian Karwowski. Improved adaptive arithmetic coding for HEVC video compression technology. Lecture Notes in Computer Science, 7594: 121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33564-8_15/.

Komorowski:2012:ECC

[524] Jacek Komorowski and Przemysław Rokita. Extrinsic camera calibration method and its performance evaluation. Lecture Notes in Computer Science, 7594:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33564-8_16/.

Kozera:2012:LEA

[525] Ryszard Kozera, Lyle Noakes, and Mariusz Rasiński. Length estimation for the adjusted exponential parameterization. Lecture Notes in Computer Science, 7594:139–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33564-8_17/.

Kozera:2012:STE

[526] Ryszard Kozera and Mateusz Śmietanka. Sharpness in trajectory estimation by piecewise-quadratics(-cubics) and cumulative chords. Lecture Notes in Computer Science, 7594:148–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33564-8_18/.

Krupinski:2012:EEB

[527] Robert Krupiński and Przemysław Mazurek. Estimation of electrooculography and blinking signals based on filter banks. Lecture Notes in Computer Science, 7594:156–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33564-8_19/.

Luczak:2012:LBD

[528] Adam Łuczak, Sławomir Maákowiak, Damian Karwowski, and Tomasz Grajek. A large barrel distortion in an acquisition system for multifocal images extraction. Lecture Notes in Computer Science, 7594: 164–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33564-8_20/.

Anonymous:2012:FMbc

[529] Anonymous. Front matter. Lecture Notes in Computer Science, 7594: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33564-8/1.

Basten:2012:MDD

[530] Twan Basten, Martijn Hendriks, Lou Somers, and Nikola Trčka. Modeldriven design-space exploration for software-intensive embedded systems. Lecture Notes in Computer Science, 7595:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33365-1_1/.

Larsen:2012:SMC

[531] Kim G. Larsen. Statistical model checking, refinement checking, optimization, … for stochastic hybrid systems. Lecture Notes in Computer Science, 7595:7–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33365-1_2/.

Akshay:2012:RTPb

[532] S. Akshay, Loïc Hélouët, Claude Jard, Didier Lime, and Olivier H. Roux. Robustness of time Petri nets under architectural constraints. Lecture Notes in Computer Science, 7595:11–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33365-1_3/.

Asarin:2012:TTT

[533] Eugene Asarin, Nicolas Basset, Marie-Pierre Béal, Aldric Degorre, and Dominique Perrin. Toward a timed theory of channel coding. Lecture Notes in Computer Science, 7595:27–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33365-1_4/.

Bertrand:2012:POT

[534] Nathalie Bertrand and Sven Schewe. Playing optimally on timed automata with random delays. Lecture Notes in Computer Science, 7595: 43–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33365-1_5/.

Carter:2012:DDT

[535] Rebekah Carter and Eva M. Navarro-López. Dynamically-driven timed automaton abstractions for proving liveness of continuous systems. Lecture Notes in Computer Science, 7595:59–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33365-1_6/.

Chilton:2012:RTS

[536] Chris Chilton, Marta Kwiatkowska, and Xu Wang. Revisiting timed specification theories: a linear-time perspective. Lecture Notes in Computer Science, 7595:75–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33365-1_7/.

Dalsgaard:2012:MCR

[537] Andreas E. Dalsgaard, Alfons Laarman, Kim G. Larsen, Mads Chr. Olesen, and Jaco van de Pol. Multi-core reachability for timed automata. Lecture Notes in Computer Science, 7595:91–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33365-1_8/.

Dimitrova:2012:CGS

[538] Rayna Dimitrova and Bernd Finkbeiner. Counterexample-guided synthesis of observation predicates. Lecture Notes in Computer Science, 7595: 107–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33365-1_9/.

Ellen:2012:CBS

[539] Christian Ellen, Sebastian Gerwinn, and Martin Fränzle. Confidence bounds for statistical model checking of probabilistic hybrid systems. Lecture Notes in Computer Science, 7595:123–138, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33365-1_10/.

Ghasemieh:2012:RBA

[540] Hamed Ghasemieh, Anne Remke, Boudewijn Haverkort, and Marco Gribaudo. Region-based analysis of hybrid Petri nets with a single general one-shot transition. Lecture Notes in Computer Science, 7595: 139–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33365-1_11/.

Herrera:2012:RQE

[541] Christian Herrera, Bernd Westphal, Sergio Feo-Arenis, Marco Muñiz, and Andreas Podelski. Reducing quasi-equal clocks in networks of timed automata. Lecture Notes in Computer Science, 7595:155–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33365-1_12/.

Kindermann:2012:SBI

[542] Roland Kindermann, Tommi Junttila, and Ilkka Niemelä. SMT-based induction methods for timed systems. Lecture Notes in Computer Science, 7595:171–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33365-1_13/.

Muniz:2012:TAD

[543] Marco Muñiz, Bernd Westphal, and Andreas Podelski. Timed automata with disjoint activity. Lecture Notes in Computer Science, 7595: 188–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33365-1_14/.

Peter:2012:CBS

[544] Hans-Jörg Peter and Bernd Finkbeiner. The complexity of bounded synthesis for timed control with partial observability. Lecture Notes in Computer Science, 7595:204–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33365-1_15/.

Rinast:2012:SDZ

[545] Jonas Rinast and Sibylle Schupp. Static detection of Zeno runs in UPPAAL networks based on synchronization matrices and two datavariable heuristics. Lecture Notes in Computer Science, 7595:220–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33365-1_16/.

Stainer:2012:FFT

[546] Amélie Stainer. Frequencies in forgetful timed automata. Lecture Notes in Computer Science, 7595:236–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33365-1_17/.

Wehrle:2012:MHS

[547] Martin Wehrle and Sebastian Kupferschmid. Mcta: Heuristics and search for timed systems. Lecture Notes in Computer Science, 7595: 252–266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33365-1_18/.

Anonymous:2012:BMg

[548] Anonymous. Back matter. Lecture Notes in Computer Science, 7595: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33365-1/1.

Anonymous:2012:FMbd

[549] Anonymous. Front matter. Lecture Notes in Computer Science, 7595: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33365-1/1.

Takimoto:2012:CES

[550] Tomoya Takimoto and Fukuhito Ooshita. Communication-efficient selfstabilization in wireless networks. Lecture Notes in Computer Science, 7596:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_1/.

Kohler:2012:SSL

[551] Sven Köhler, Volker Turau, and Gerhard Mentges. Self-stabilizing local k-placement of replicas with minimal variance. Lecture Notes in Computer Science, 7596:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33536-5_2/.

Neggazi:2012:SSA

[552] Brahim Neggazi and Mohammed Haddad. Self-stabilizing algorithm for maximal graph partitioning into triangles. Lecture Notes in Computer Science, 7596:31–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33536-5_3/.

Malekpour:2012:BAS

[553] Mahyar R. Malekpour. Brief announcement: Self-stabilizing synchronization of arbitrary digraphs in presence of faults. Lecture Notes in Computer Science, 7596:43–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33536-5_4/.

Datta:2012:BAS

[554] Ajoy K. Datta and Stéphane Devismes. Brief announcement: Selfstabilizing silent disjunction in an anonymous network. Lecture Notes in Computer Science, 7596:46–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33536-5_5/.

Ooshita:2012:SSM

[555] Fukuhito Ooshita and Sébastien Tixeuil. On the self-stabilization of mobile oblivious robots in uniform rings. Lecture Notes in Computer Science, 7596:49–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_6/.

Devismes:2012:OGE

[556] Stéphane Devismes and Anissa Lamani. Optimal grid exploration by asynchronous oblivious robots. Lecture Notes in Computer Science, 7596: 64–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_7/.

Michail:2012:TPP

[557] Othon Michail and Ioannis Chatzigiannakis. Terminating population protocols via some minimal global knowledge assumptions. Lecture Notes in Computer Science, 7596:77–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33536-5_8/.

Fugger:2012:BAD

[558] Matthias Függer and Alexander Kößler. Brief announcement: The degrading effect of forgetting on a synchronizer. Lecture Notes in Computer Science, 7596:90–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-33536-5_9.

Izumi:2012:BAM

[559] Tomoko Izumi, Yukiko Yamauchi, and Sayaka Kamei. Brief announcement: Mobile agent rendezvous on edge evolving rings. Lecture Notes in Computer Science, 7596:92–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33536-5_10/.

Liu:2012:HLE

[560] Yanhong A. Liu, Scott D. Stoller, and Bo Lin. High-level executable specifications of distributed algorithms. Lecture Notes in Computer Science, 7596:95–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_11/.

Jarraya:2012:FVS

[561] Yosr Jarraya and Arash Eghtesadi. Formal verification of security preservation for migrating virtual machines in the cloud. Lecture Notes in Computer Science, 7596:111–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33536-5_12/.

Adamek:2012:EPT

[562] Jordan Adamek and Mikhail Nesterenko. Evaluating practical tolerance properties of stabilizing programs through simulation: The case of propagation of information with feedback. Lecture Notes in Computer Science, 7596:126–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33536-5_13/.

Dolev:2012:SSE

[563] Shlomi Dolev and Ariel Hanemann. Self-stabilizing end-to-end communication in (bounded capacity, omitting, duplicating and non-FIFO) dynamic networks. Lecture Notes in Computer Science, 7596:133–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33536-5_14/.

Ducourthial:2012:SSD

[564] Bertrand Ducourthial and Véronique Cherfaoui. Self-stabilizing distributed data fusion. Lecture Notes in Computer Science, 7596:148– 162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_15/.

Johnen:2012:SSS

[565] Colette Johnen. From self- to self-stabilizing with service guarantee 1hop weight-based clustering. Lecture Notes in Computer Science, 7596: 163–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_16/.

Chen:2012:BAV

[566] Jingshu Chen. Brief announcement: Verification of stabilizing programs with SMT solvers. Lecture Notes in Computer Science, 7596: 179–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33536-5_17/.

Muftuoglu:2012:BAM

[567] Can Arda Muftuoglu, Péter Bokor, and Neeraj Suri. Brief announcement: MP-state: State-aware software model checking of message-passing systems. Lecture Notes in Computer Science, 7596:183–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33536-5_18/.

Ateniese:2012:OAM

[568] Giuseppe Ateniese and Roberto Baldoni. Oblivious assignment with m slots. Lecture Notes in Computer Science, 7596:187–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33536-5_19/.

Alaggan:2012:BNI

[569] Mohammad Alaggan and Sébastien Gambs. BLIP: Non-interactive differentially-private similarity computation on Bloom filters. Lecture Notes in Computer Science, 7596:202–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33536-5_20/.

Anonymous:2012:FMbe

[570] Anonymous. Front matter. Lecture Notes in Computer Science, 7596: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33536-5/1.

Gong:2012:FRS

[571] Tao Gong. Face recognition by searching most similar sample with immune learning. Lecture Notes in Computer Science, 7597:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33757-4_1/.

Pierrard:2012:MOA

[572] Thomas Pierrard and Carlos A. Coello Coello. A multi-objective artificial immune system based on hypervolume. Lecture Notes in Computer Science, 7597:14–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33757-4_2/.

Textor:2012:CSN

[573] Johannes Textor. A comparative study of negative selection based anomaly detection in sequence data. Lecture Notes in Computer Science, 7597:28–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_3/.

Lim:2012:IIS

[574] TiongHoo Lim, HuiKeng Lau, Jon Timmis, and Iain Bate. Immuneinspired self healing in wireless sensor networks. Lecture Notes in Computer Science, 7597:42–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33757-4_4/.

Figueredo:2012:BGS

[575] Grazziela P. Figueredo, Peer-Olaf Siebers, Uwe Aickelin, and Stephanie Foan. A beginner’s guide to systems simulation in immunology. Lecture Notes in Computer Science, 7597:57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33757-4_5/.

Tsang:2012:CBM

[576] Wilburn W. P. Tsang and Henry Y. K. Lau. Clustering-based multiobjective immune optimization evolutionary algorithm. Lecture Notes in Computer Science, 7597:72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33757-4_6/.

Gogolinska:2012:PNA

[577] Anna Gogolinska and Wieslaw Nowak. Petri nets approach to modeling of immune system and autism. Lecture Notes in Computer Science, 7597:86–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_7/.

deOliveira:2012:MII

[578] Cicero Hildenberg Lima de Oliveira, Thayna Baptista Moroso, and Fabio Hugo Souza Matos. Mathematical implementation of interaction between malaria and immune system. Lecture Notes in Computer Science, 7597: 100–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_8/.

Jansen:2012:CLC

[579] Thomas Jansen and Christine Zarges. Computing longest common subsequences with the B-cell algorithm. Lecture Notes in Computer Science, 7597:111–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/chapter/10.1007/
978-3-642-33757-4_9/.

Fu:2012:BTI

[580] Jun Fu, Huan Yang, Yiwen Liang, and Chengyu Tan. Bait a trap: Introducing natural killer cells to artificial immune system for spyware detection. Lecture Notes in Computer Science, 7597:125–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33757-4_10/.

Elsayed:2012:ACI

[581] Samir A. Mohamed Elsayed, Sanguthevar Rajasekaran, and Reda A. Ammar. AC-CS: An immune-inspired associative classification algorithm. Lecture Notes in Computer Science, 7597:139–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33757-4_11/.

Chelly:2012:RDN

[582] Zeineb Chelly and Zied Elouedi. RC-DCA: a new feature selection and signal categorization technique for the Dendritic Cell Algorithm based on rough set theory. Lecture Notes in Computer Science, 7597: 152–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_12/.

Mariani:2012:AIN

[583] Viviana Cocco Mariani, Leandro dos Santos Coelho, and Anderson Duck. Artificial immune network approach with Beta differential operator applied to optimization of heat exchangers. Lecture Notes in Computer Science, 7597:166–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33757-4_13/.

Mostardinha:2012:NSA

[584] Patricia Mostardinha, Bruno Filipe Faria, André Zúquete, and Fernão Vistulo de Abreu. A negative selection approach to intrusion detection. Lecture Notes in Computer Science, 7597:178–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33757-4_14/.

Tarapore:2012:CNS

[585] Danesh Tarapore, Anders Lyhne Christensen, Pedro U. Lima, and Jorge Carneiro. Clonal Expansion without Self-replicating Entities. Lecture Notes in Computer Science, 7597:191–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33757-4_15/.

Mohammadi:2012:RTA

[586] Mahdi Mohammadi, Ahmad Akbari, Bijan Raahemi, and Babak Nassersharif. A real time anomaly detection system based on probabilistic artificial immune based algorithm. Lecture Notes in Computer Science, 7597: 205–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_16/.

Cai:2012:CIN

[587] Zixing Cai, Xingbao Liu, and Xiaoping Ren. CSA/IE: Novel clonal selection algorithm with information exchange for high dimensional global optimization problems. Lecture Notes in Computer Science, 7597:218– 231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_17/.

Pinacho:2012:EAA

[588] Pedro Pinacho, Iván Pau, Max Chacón, and Sergio Sánchez. An ecological approach to anomaly detection: The EIA model. Lecture Notes in Computer Science, 7597:232–245, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33757-4_18/.

Musselle:2012:RCD

[589] Chris Musselle. Rethinking concepts of the Dendritic Cell Algorithm for multiple data stream analysis. Lecture Notes in Computer Science, 7597: 246–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33757-4_19/.

Giancarlo:2012:SBM

[590] Raffaele Giancarlo and Filippo Utro. Stability-based model selection for high throughput genomic data: An algorithmic paradigm. Lecture Notes in Computer Science, 7597:260–270, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33757-4_20/.

Anonymous:2012:FMbf

[591] Anonymous. Front matter. Lecture Notes in Computer Science, 7597: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33757-4/1.

Faraj:2012:RSI

[592] Noura Faraj, Jean-Marc Thiery, Isabelle Bloch, Nadège Varsier, and Joe Wiart. Robust and scalable interactive freeform modeling of high definition medical images. Lecture Notes in Computer Science, 7599: 1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_1/.

Cui:2012:ASA

[593] Xinyi Cui, Shaoting Zhang, Yiqiang Zhan, Mingchen Gao, and Junzhou Huang. 3D anatomical shape Atlas construction using mesh quality preserved deformable models. Lecture Notes in Computer Science, 7599: 12–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_2/.

Albrecht:2012:AFR

[594] Thomas Albrecht and Thomas Vetter. Automatic fracture reduction. Lecture Notes in Computer Science, 7599:22–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33463-4_3/.

Whitmarsh:2012:ARC

[595] Tristan Whitmarsh, Luis M. Del Río Barquero, Silvana Di Gregorio, and Jorge Malouf Sierra. Age-related changes in vertebral morphometry by statistical shape analysis. Lecture Notes in Computer Science, 7599:30–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_4/.

Zhang:2012:AMF

[596] Ju Zhang, Duane Malcolm, Jacqui Hislop-Jambrich, C. David L. Thomas, and Poul Nielsen. Automatic meshing of femur cortical surfaces from clinical CT images. Lecture Notes in Computer Science, 7599:40–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33463-4_5/.

Jensen:2012:SSR

[597] Rasmus R. Jensen, Oline V. Olesen, Rasmus R. Paulsen, Mike van der Poel, and Rasmus Larsen. Statistical surface recovery: a study on ear canals. Lecture Notes in Computer Science, 7599:49–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/chapter/10.1007/978-3-642-33463-4_6/.

Zhang:2012:ASC

[598] Hong Zhang, Yuanfeng Jiao, Yongjie Zhang, and Kenji Shimada. Automated segmentation of cerebral aneurysms based on conditional random field and gentle adaboost. Lecture Notes in Computer Science, 7599: 59–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_7/.

Kim:2012:CMD

[599] Youngjun Kim, Seungbin Lee, Frédérick Roy, Deukhee Lee, Laehyun Kim, and Sehyung Park. Carving mesh with deformation for soft tissue removal simulation. Lecture Notes in Computer Science, 7599:70– 79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-33463-4_8/.

Gao:2012:SRS

[600] Yi Gao and Sylvain Bouix. Synthesis of realistic subcortical anatomy with known surface deformations. Lecture Notes in Computer Science, 7599:80–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_9/.

Gutman:2012:RSC

[601] Boris A. Gutman, Ryan McComb, Jay Sung, Won Moon, and Paul M. Thompson. Robust shape correspondence via spherical patch matching for atlases of partial skull models. Lecture Notes in Computer Science, 7599:89–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_10/.

Kahnt:2012:FPM

[602] Max Kahnt, Heiko Ramm, Hans Lamecker, and Stefan Zachow. Featurepreserving, multi-material mesh generation using hierarchical oracles. Lecture Notes in Computer Science, 7599:101–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-33463-4_11/.

Li:2012:PCS

[603] Gang Li, Jingxin Nie, and Dinggang Shen. Partition cortical surfaces into supervertices: Method and application. Lecture Notes in Computer Science, 7599:112–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33463-4_12/.

Park:2012:ABE

[604] Jeonghyung Park, Suzanne M. Shontz, and Corina S. Drapaca. Automatic boundary evolution tracking via a combined level set method and mesh warping technique: Application to hydrocephalus. Lecture Notes in Computer Science, 7599:122–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-33463-4_13/.

Pepe:2012:SRG

[605] Antonietta Pepe, Laura Brandolini, Marco Piastra, Juha Koikkalainen, and Jarmo Hietala. Simplified Reeb graph as effective shape descriptor for the striatum. Lecture Notes in Computer Science, 7599:134– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/chapter/10.1007/978-
3-642-33463-4_14/.

Sibbing:2012:TAQ

[606] Dominik Sibbing, Hans-Christian Ebke, Kai Ingo Esser, and Leif Kobbelt. Topology aware quad dominant meshing for vascular structures. Lecture Notes in Computer Science, 7599:147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-33463-4_15/.

Voglreiter:2012:VRT

[607] Philip Voglreiter, Markus Steinberger, Dieter Schmalstieg, and Bernhard Kainz. Volumetric real-time particle-based representation of large unstructured tetrahedral polygon meshes. Lecture Notes in Computer Science, 7599:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-33463-4_16/.

Anonymous:2012:BMh

[608] Anonymous. Back matter. Lecture Notes in Computer Science, 7599: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-33463-4/1.

Anonymous:2012:FMbg

[609] Anonymous. Front matter. Lecture Notes in Computer Science, 7599: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-33463-4/1.

Finkel:2012:RPI

[610] Alain Finkel, Jérôme Leroux, and Igor Potapov, editors. Reachability Problems: 6th International Workshop, RP 2012, Bordeaux, France, September 17–19, 2012. Proceedings, volume 7550 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33511-X (print), 3-642-33512-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33512-9.

Golumbic:2012:GTC

[611] Martin Charles Golumbic, Michal Stern, Avivit Levy, and Gila Morgenstern, editors. Graph-Theoretic Concepts in Computer Science: 38th International Workshop, WG 2012, Jerusalem, Israel, June 26–28, 2012, Revised Selcted Papers, volume 7551 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34610-3 (print), 3-642-34611-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-34611-8.

Villa:2012:ANNa

[612] Alessandro E. P. Villa, Włodzisław Duch, Péter Érdi, Francesco Masulli, and Günther Palm, editors. Artificial Neural Networks and Machine Learning — ICANN 2012: 22nd International Conference on Artificial Neural Networks, Lausanne, Switzerland, September 11–14, 2012, Proceedings, Part I, volume 7552 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33268-4 (print), 3-642-33269-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-33269-2.

Villa:2012:ANNb

[613] Alessandro E. Villa, Włodzisław Duch, Péter Érdi, Francesco Masulli, and Günther Palm, editors. Artificial Neural Networks and Machine Learning — ICANN 2012: 22nd International Conference on Artificial Neural Networks, Lausanne, Switzerland, September 11–14, 2012, Proceedings, Part II, volume 7553 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33265-X (print), 3-642-33266-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33266-1.

Junior:2012:PLB

[614] Francisco Heron de Carvalho Junior and Luis Soares Barbosa, editors. Programming Languages: 16th Brazilian Symposium, SBLP 2012, Natal, Brazil, September 23–28, 2012. Proceedings, volume 7554 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33181-5 (print), 3-642-33182-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33182-4.

Hu:2012:CLT

[615] Hao Hu, Xiaoning Shi, Robert Stahlbock, and Stefan Voß, editors. Computational Logistics: Third International Conference, ICCL 2012, Shanghai, China, September 24–26, 2012. Proceedings, volume 7555 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33586-1 (print), 3-642-33587-X (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-33587-7.

Domingo-Ferrer:2012:PSD

[616] Josep Domingo-Ferrer and Ilenia Tinnirello, editors. Privacy in Statistical Databases: UNESCO Chair in Data Privacy, International Conference, PSD 2012, Palermo, Italy, September 26–28, 2012. Proceedings, volume 7556 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33626-4 (print), 3-642-33627-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33627-0.

Ramsay:2012:AIM

[617] Allan Ramsay and Gennady Agre, editors. Artificial Intelligence: Methodology, Systems, and Applications: 15th International Conference, AIMSA 2012, Varna, Bulgaria, September 12–15, 2012. Proceedings, volume 7557 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33184-X (print), 3-642-33185-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33185-5.

Popescu:2012:AWB

[618] Elvira Popescu, Qing Li, Ralf Klamma, Howard Leung, and Marcus Specht, editors. Advances in Web-Based Learning — ICWL 2012: 11th International Conference, Sinaia, Romania, September 2–4, 2012. Proceedings, volume 7558 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33641-8 (print), 3-642-33642-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33642-3.

Salah:2012:HBUb

[619] Albert Ali Salah, Javier Ruiz del Solar, Çetin Meriçli, and Pierre-Yves Oudeyer, editors. Human Behavior Understanding: Third International Workshop, HBU 2012, Vilamoura, Portugal, October 7, 2012. Proceedings, volume 7559 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34013-X (print), 3-642-34014-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-34014-7.

Kahl:2012:RAM

[620] Wolfram Kahl and Timothy G. Griffin, editors. Relational and Algebraic Methods in Computer Science: 13th International Conference, RAMiCS 2012, Cambridge, UK, September 17–20, 2012. Proceedings, volume 7560 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33313-3 (print), 3-642-33314-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33314-9.

Ehrig:2012:GTI

[621] Hartmut Ehrig, Gregor Engels, Hans-Jörg Kreowski, and Grzegorz Rozenberg, editors. Graph Transformations: 6th International Conference, ICGT 2012, Bremen, Germany, September 24–29, 2012. Proceedings, volume 7562 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33653-1 (print), 3-642-33654-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33654-6.

Ravenscroft:2012:CLC

[622] Andrew Ravenscroft, Stefanie Lindstaedt, Carlos Delgado Kloos, and Davinia Hernández-Leo, editors. 21st Century Learning for 21st Century Skills: 7th European Conference of Technology Enhanced Learning, EC-TEL 2012, Saarbrücken, Germany, September 18–21, 2012. Proceedings, volume 7563 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33262-5 (print), 3-642-33263-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33263-0.

Cortesi:2012:CIS

[623] Agostino Cortesi, Nabendu Chaki, Khalid Saeed, and Sławomir Wierzchoń, editors. Computer Information Systems and Industrial Management: 11th IFIP TC 8 International Conference, CISIM 2012, Venice, Italy, September 26–28, 2012. Proceedings, volume 7564 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33259-5 (print), 3-642-33260-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33260-9.

Ganascia:2012:DSI

[624] Jean-Gabriel Ganascia, Philippe Lenca, and Jean-Marc Petit, editors. Discovery Science: 15th International Conference, DS 2012, Lyon, France, October 29–31, 2012. Proceedings, volume 7569 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33491-1 (print), 3-642-33492-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33492-4.

Durrleman:2012:STI

[625] Stanley Durrleman, Tom Fletcher, Guido Gerig, and Marc Niethammer, editors. Spatio-temporal Image Analysis for Longitudinal and Time-Series Image Data: Second International Workshop, STIA 2012, Held in Conjunction with MICCAI 2012, Nice, France, October 1, 2012. Proceedings, volume 7570 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33554-3 (print), 3-642-33555-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33555-6.

Fitzgibbon:2012:CVEa

[626] Andrew Fitzgibbon, Svetlana Lazebnik, Pietro Perona, Yoichi Sato, and Cordelia Schmid, editors. Computer Vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings, Part I, volume 7572 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33717-1 (print), 3-642-33718-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33718-5.

Fitzgibbon:2012:CVEb

[627] Andrew Fitzgibbon, Svetlana Lazebnik, Pietro Perona, Yoichi Sato, and Cordelia Schmid, editors. Computer Vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings, Part III, volume 7574 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33711-2 (print), 3-642-33712-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-33712-3.

Fitzgibbon:2012:CVEc

[628] Andrew Fitzgibbon, Svetlana Lazebnik, Pietro Perona, Yoichi Sato, and Cordelia Schmid, editors. Computer Vision — ECCV 2012: 12th European Conference on Computer Vision, Florence, Italy, October 7–13, 2012, Proceedings, Part VII, volume 7578 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-33785-6 (print), 3-642-33786-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-33786-4.

Marquez:2012:CGX

[629] Alberto Márquez, Pedro Ramos, and Jorge Urrutia, editors. Computational Geometry: XIV Spanish Meeting on Computational Geometry, EGC 2011, Dedicated to Ferran Hurtado on the Occasion of His 60th Birthday, Alcalá de Henares, Spain, June 27–30, 2011, Revised Selected Papers, volume 7579 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-34190-X (print), 3-642-34191-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-34191-5.

Cranefield:2012:ABS

[630] Stephen Cranefield and Insu Song, editors. Agent Based Simulation for a Sustainable Society and Multi-agent Smart Computing: International Workshops, PRIMA 2011, Wollongong, Australia, November 14, 2011 Revised Selected Papers, volume 7580 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-35611-7 (print), 3-642-35612-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-35612-4.

Fusiello:2012:CVEa

[631] Andrea Fusiello, Vittorio Murino, and Rita Cucchiara, editors. Computer Vision — ECCV 2012. Workshops and Demonstrations: Florence, Italy, October 7–13, 2012, Proceedings, Part I, volume 7583 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33862-3 (print), 3-642-33863-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33863-2.

Fusiello:2012:CVEb

[632] Andrea Fusiello, Vittorio Murino, and Rita Cucchiara, editors. Computer Vision — ECCV 2012. Workshops and Demonstrations: Florence, Italy, October 7–13, 2012, Proceedings, Part II, volume 7584 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33867-4 (print), 3-642-33868-2 (e-book).

ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33868-7.

Fusiello:2012:CVEc

[633] Andrea Fusiello, Vittorio Murino, and Rita Cucchiara, editors. Computer Vision — ECCV 2012. Workshops and Demonstrations: Florence, Italy, October 7–13, 2012, Proceedings, Part III, volume 7585 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33884-4 (print), 3-642-33885-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-33885-4.

France:2012:MDE

[634] Robert B. France, Jürgen Kazmeier, Ruth Breu, and Colin Atkinson, editors. Model Driven Engineering Languages and Systems: 15th International Conference, MODELS 2012, Innsbruck, Austria, September 30–October 5, 2012. Proceedings, volume 7590 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33665-5 (print), 3-642-33666-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-33666-9.

Paoli:2012:SOC

[635] Flavio De Paoli, Ernesto Pimentel, and Gianluigi Zavattaro, editors. Service-Oriented and Cloud Computing: First European Conference, ES-OCC 2012, Bertinoro, Italy, September 19–21, 2012. Proceedings, volume 7592 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33426-1 (print), 3-642-33427-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-642-
33427-6.

Bolc:2012:CVG

[636] Leonard Bolc, Ryszard Tadeusiewicz, Leszek J. Chmielewski, and Konrad Wojciechowski, editors. Computer Vision and Graphics: International Conference, ICCVG 2012, Warsaw, Poland, September 24–26, 2012. Proceedings, volume 7594 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33563-2 (print), 3-642-33564-0 (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-33564-8.

Jurdzinski:2012:FMA

[637] Marcin Jurdziński and Dejan Ničković, editors. Formal Modeling and Analysis of Timed Systems: 10th International Conference, FORMATS 2012, London, UK, September 18–20, 2012. Proceedings, volume 7595 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33364-8 (print), 3-642-33365-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33365-1.

Richa:2012:SSS

[638] Andréa W. Richa and Christian Scheideler, editors. Stabilization, Safety, and Security of Distributed Systems: 14th International Symposium, SSS 2012, Toronto, Canada, October 1–4, 2012. Proceedings, volume 7596 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33535-7 (print), 3-642-33536-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-33536-5.

Coello:2012:AIS

[639] Carlos A. Coello Coello, Julie Greensmith, Natalio Krasnogor, Pietro Liò, Giuseppe Nicosia, and Mario Pavone, editors. Artificial Immune Systems: 11th International Conference, ICARIS 2012, Taormina, Italy, August 28–31, 2012. Proceedings, volume 7597 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33756-2 (print), 3-642-33757-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-33757-4.

Levine:2012:MPM

[640] Joshua A. Levine, Rasmus R. Paulsen, and Yongjie Zhang, editors. Mesh Processing in Medical Image Analysis 2012: MICCAI 2012 International Workshop, MeshMed 2012, Nice, France, October 1, 2012. Proceedings, volume 7599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-33462-8 (print), 3-642-33463-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-33463-4.